### v1

# Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations

Abstract
Large language models like GPT, LLAMA, and Claude have become incredibly powerful at generating text, but they are still black boxes, so it is hard to understand how they decide what to say. That lack of transparency can be problematic, especially in fields where trust and accountability matter. To help with this, we introduce SMILE, a new method that explains how these models respond to different parts of a prompt. SMILE is model-agnostic and works by slightly changing the input, measuring how the output changes, and then highlighting which words had the most impact. Create simple visual heat maps showing which parts of a prompt matter the most. We tested SMILE on several leading LLMs and used metrics such as accuracy, consistency, stability, and fidelity to show that it gives clear and reliable explanations. By making these models easier to understand, SMILE brings us one step closer to making AI more transparent and trustworthy.
Figure 1: Highlighting the contribution of each word in the input prompt of a Commercial LLM
1 Introduction
Large Language Models (LLMs) have become powerful tools in recent years. They can perform tasks ranging from answering questions with remarkable accuracy [1] to drafting essays and other written content with coherence and style [2], creating art that spans various mediums and styles [3], coding across multiple programming languages [4], debugging software [4], writing technical documentation [5], and even generating algorithms [1].
LLMs function by leveraging deep learning architectures, particularly the Transformer model. This model uses self-attention mechanisms to process and generate text by analysing relationships between words in large data sets [6]. These models are trained in various corpora, including books, articles, code, and web content, allowing them to learn language patterns, context, and syntax on an unheard scale [7].
Their capabilities continue to grow, turning them into essential tools for a wide range of tasks [2]. These models, trained in various texts, have transformed how we interact with technology and how machines understand and generate human language [6]. LLMs such as GPT, BERT, and Claude leverage the Transformer architecture to process and generate language based on patterns learned from extensive datasets [7]. Although they excel in many NLP tasks, understanding their strengths and limitations is the key to using them effectively and responsibly [8].
These models continue to evolve, pushing the boundaries of AI applications in language understanding and generation [2]. However, the nature of these models is known as black boxes and lacks transparency [9]. One of their important problems is how these models decide which words or parts of input to focus on, which remains a complex and fascinating process [10].
At their core, LLMs rely on patterns and relationships learned during training to decide which input elements are the most important [6]. This process is more than just identifying individual keywords; it is about picking up sophisticated connections and context that shape the model’s responses [10]. Understanding how this works is crucial: it explains why these models excel in some scenarios but struggle in others. It also highlights why they can occasionally produce responses that are biased, irrelevant, or even harmful [11].
By studying how LLMs emphasise words and phrases, we can better understand how they work [10]. This can help us make them more transparent and reliable, ensuring they focus on what truly matters in a given context. It also allows us to address potential problems, such as biases or alignment problems, making these tools safer and more ethical [11].
As LLMs continue to integrate into our daily lives and decision-making processes, understanding their inner workings is more important than ever [9]. Research on how they prioritise language helps us refine their performance, ensure ethical use, and unlock their full potential to benefit society [11] responsibly.
To solve this problem, we propose a novel approach of SMILE (Statistical Model-agnostic Interpretability with Local Explanations) to extract a visual heat map that shows the influence of each word in responding to these models [12]. This method extends the foundation laid by LIME [13] by incorporating Empirical Cumulative Distribution Function (ECDF) statistical distances, significantly improving robustness and making interpretability more effective.
SMILE produces visual heatmaps that reveal the impact of individual words or phrases in text input on the model’s output as illustrated in Fig 1. For example, when a language model processes an instruction or query, the heatmap highlights the importance of specific terms in influencing the outcome. This offers users a clear visual representation of how the model interprets and prioritises different parts of the input text [10].
To ensure a thorough evaluation of our approach, we employ a suite of metrics, including accuracy, fidelity, stability, and consistency [14]. These metrics demonstrate that SMILE provides reliable insights into model behaviour [15].
The contributions of our work include the following.
  • Introducing a New Interpretability Framework: SMILE is a cutting-edge method that visualises the role of individual words in shaping model output, offering a deeper understanding of how LLMs process language [12].
  • Promoting Transparency and Trust in LLMs: By visually linking text input to model decisions, SMILE enhances transparency, making LLMs more trustworthy and enabling their practical use in high-stakes domains such as healthcare, education, and legal systems.
  • Comprehensive Evaluation for Robustness: Our evaluation leverages metrics such as stability, fidelity, accuracy, and consistency, establishing a new standard for evaluating the interpretability and reliability of language models [16].
2 Literature review
This section summarises the most recent advances in large language models (LLMs) and their growth in various applications. We also discuss Explainable Artificial Intelligence (XAI), a field dedicated to making AI systems more transparent and easier to understand. Finally, we explore how these two areas come together, focusing on the ongoing efforts to create explainable large language models. We discuss the key challenges involved and the techniques being developed to improve how we interpret and trust these powerful models.
2.1 Large Language Models
Large language models (LLMs) have become increasingly common thanks to significant strides in deep learning, better access to data, and more powerful computing resources. These models are trained on vast amounts of text and can generate responses that sound natural, answer questions, and support a wide range of language-based tasks. Over time, several notable LLMs have helped push the field forward, each bringing new ideas and improvements.
GPT (OpenAI) [1] is one of the most well-known autoregressive transformer-based models, capable of generating coherent and contextually rich text. Google’s T5 [17] introduced the text-to-text framework, where all NLP tasks are framed as text generation problems, improving transfer learning and efficiency. Similarly, PaLM (Google) [18] leverages the Pathways architecture for large-scale language modelling, optimising multitask learning across different NLP tasks.
Anthropic’s Claude [19] focuses on ethical AI and safe interactions designed to minimise harmful outputs while maintaining interpretability. Google’s Gemma 2 [20] is optimised for logical reasoning and structured data interpretation, making it a strong candidate for academic and enterprise use cases. Meta’s LLaMA [5] is an efficient, open-source model tailored for various NLP applications, offering researchers an accessible yet powerful alternative to proprietary models.
EleutherAI’s GPT-NeoX [21] is an open-source alternative to GPT, providing robust text generation capabilities while allowing customisation. AI21 Labs’ Jurassic-2 [22] is designed for large-scale enterprise applications, emphasising contextual understanding and user intent. Cohere’s Command R+ [23] integrates retrieval-augmented generation, improving fact-based responses and minimising hallucinations.
Furthermore, models like BERT (Google) [24] have NLP through bidirectional context understanding, significantly improving sentiment analysis and question-answering tasks. Microsoft’s Phi-4 focuses on efficiency and cost-effectiveness, delivering competitive performance with fewer parameters. Mistral and Alibaba’s Qwen 2.5 push the boundaries of model adaptability and fine-tuning capabilities, making them suitable for multilingual and domain-specific applications.
DeepSeek-LLM [25] is another promising open-source model, competing with top-tier models regarding benchmark performance. Models such as Mixtral and StableLM have been developed with efficiency and specific NLP task performance in mind, ensuring a balance between accuracy and computational demands.
Recent developments have introduced models such as DeepSeek-R1, which focused on advanced reasoning capabilities, and Qwen2.5-Max, which explores large-scale Mixture-of-Experts (MoE) models to enhance efficiency and performance. OpenAI’s o3-mini pushes the frontier of cost-effective reasoning, balancing performance with computational efficiency. DeepSeek-V3, the first open-sourced GPT-4o level model, represents a significant step forward in open-source LLM development.
The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications:
Category Models Description General-Purpose GPT-4 [26], GPT-4o [27], Claude 3 [28], Gemini 1.5 [29], Command R+ [30], Mistral [31] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [32], Mistral [31], Mixtral [33], Falcon 180B [34], DeepSeek-V3 [35], OpenChat [36] Customizable models for research applications Retrieval-Augmented Command R+ [30], GPT-4 Bing [26], Claude 3 RAG [28], Gemini Search [29] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [28], Gemini 1.5 [29], GPT-4o [27], Grok-1.5 [37] Specialized in processing long-form text Advanced Reasoning GPT-4 [26], Claude 3 Opus [28], Gemini 1.5 [29], Mixtral [33], Grok-1.5 [37] Logical reasoning, multi-task learning, problem-solving Ethically-Aligned Claude 3 [28], Gemini [29], GPT-4 [26], Grok [37] Emphasize safety, reducing bias, ethical AI Mixture-of-Experts (MoE) Mixtral [33], Qwen2.5-Max [38], GShard [39], Switch Transformer [40] Improved efficiency and performance via MoE Multimodal GPT-4o [27], Gemini 1.5 [29], Claude 3 [28], Kosmos-2 [41], Grok-1.5V [37] Capable of processing both text and visual inputs Multilingual BLOOM [42], XGLM [43], mBERT [44], XLM-R [45], ByT5 [46] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [47], Galactica [48], Legal-BERT [49], StarCoder2 [50], Code LLaMA [51] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [52], Vicuna [53], ChatGLM [54], Baize [55], OpenChat [36] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [56], TinyLLaMA [57], DistilGPT2 [58], GPT-2 [59], LLaMA 2 7B [60] Designed for fast inference and deployment on edge devices
Figure 2: Extended LLM Categories, Models, and Descriptions
2.2 Explainable AI (XAI)
The increasing complexity of machine learning models has made understanding how they make decisions more important than ever. Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [61, 14].
Explainability methods can be broadly categorised as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [9].
In contrast, post-hoc methods, such as Local Interpretable Model-agnostic Explanations (LIME) [13] and SHAP (Shapley Additive Explanations) [134], are applied after model training to explain the decisions of complex, often “black-box” models like neural networks and ensemble methods. These approaches provide valuable insights without altering the underlying model structure [63]. Post-hoc explainability methods can be classified into global and local explanations.
Global explanations provide an overarching view of the model’s decision-making patterns across an entire dataset, helping to identify trends and biases [62, 64]. In contrast, local explanations focus on specific predictions, providing a detailed breakdown of how individual inputs influence model outputs. LIME and its numerous enhancements fall under local explainability methods and are designed to provide granular, case-specific interpretations [13].
LIME has become one of the most widely used explainability techniques, but over time, researchers have introduced various improvements to address its limitations and expand its usefulness across different applications. As shown in Figure 3, enhancements to LIME can be grouped into four main categories:
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [65], S-LIME [66], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behaviour, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimisation strategies to make LIME more efficient and stable. Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [67], TS-MULE for time series data [68], Graph LIME for graph-structured data [69], Sound LIME for audio data [70], and B-LIME for ECG signal data [71] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to improve LIME’s approach, including using alternative surrogate models, fine-tuning distance parameters, optimising sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
Figure 3: Key strategies for enhancing LIME, including improvements in sampling, distance parameters, surrogate models, and optimisation.
2.3 Explainable AI (XAI) for Large Language Models
Large Language Models (LLMs) have revolutionised natural language processing, enabling text generation, translation, and question answering tasks. However, their opaque “black-box” nature raises concerns regarding interpretability, fairness, and trustworthiness [72, 73]. Explainability in LLMs is crucial for ensuring accountability, enhancing user trust, and mitigating biases. This literature review explores existing research on explainability methods, applications, and challenges in LLMs.
The conceptual framework for explainability in LLMs includes four key dimensions: Post-hoc Explainability, Mechanistic Interpretability, Human-centric Explainability, and Applications and Challenges. These categories reflect the main approaches and practical considerations in enhancing the transparency and trustworthiness of LLMs.
After training, post-hoc explainability methods aim to interpret LLMs, providing insights into their decision-making processes [72, 74]. Zhao et al. [72] present a taxonomy of explainability techniques for LLMs, categorising approaches into local and global explanations, and discussing limitations and future research directions. Similarly, Bender et al. [73] highlight the risks of opaque LLMs, including biases and misinformation, and call for increased transparency and accountability. Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [75, 76]. Amara et al. [77] recently introduced ConceptX, a concept-level attribution method highlighting semantically meaningful input phrases through a coalition-based Shapley framework, offering robust and human-aligned explanations.
Mechanistic interpretability seeks to uncover how LLMs process information internally [78, 79, 80]. Olah et al. [78] propose reverse-engineering techniques to dissect internal representations. Nanda et al. [79] examine how LLMs develop abstract representations and generalisation, introducing progress measures for interpretability research. Recent work by Anthropic [80] demonstrates the role of mechanistic interpretability in understanding and controlling model behaviour. Additionally, the SEER framework [81] introduces self-explainability mechanisms to enhance the interpretability of internal representations in LLMs. The CELL framework [82] further advances mechanistic interpretability by integrating concept-based explanations directly into the training and representation learning of language models.
Recent advancements from the Transformer Circuits framework have further enriched mechanistic interpretability. Using causal scrubbing, Meng et al. [83] introduced Attribution Graphs to uncover emergent structures within neural networks. In contrast, Nanda et al. [84] investigated the scaling of monosemanticity, extracting interpretable features from state-of-the-art models. Additional recent contributions explore specific tasks such as relevance estimation [85], emotion inference [86], and geospatial reasoning [87]. These works leverage techniques like activation patching and sparse autoencoders to reveal LLM internals. Sparse interpretability methods are further reviewed comprehensively in [88]. Neuroscientific perspectives have also been applied, using dynamical systems theory to model token-level trajectories within transformers [89].
Human-centric explainability focuses on making LLM outputs understandable to non-experts [90, 91, 92]. Ji et al. [90] investigate hallucination in LLM outputs and propose techniques for detecting and mitigating incorrect responses. Krause et al. [91] assess reasoning abilities compared to human performance, and Martens et al. [92] explore counterfactual explanations to improve human understanding. Recent works extend this perspective by enabling interactive explanations of black-box models through fine-tuned LLMs [93] and by benchmarking explanation quality in clinical domains [94]. These approaches improve the accessibility and transparency of LLM outputs in real-world decision-making.
Explainability enhances debugging, bias detection, regulatory compliance, and user trust in LLMs [95]. Challenges remain, including defining “meaning” in LLM-generated outputs [95], managing the ethical and societal implications of black-box models, and balancing performance with interpretability. Recent work has demonstrated SMILE’s applicability beyond language, such as in instruction-based image editing [96], highlighting its modality-agnostic nature. New efforts address rare concept discovery via tailored sparse autoencoders [97] and propose hybrid architectures like concept bottleneck models [98] and symbolic program compositions [99] to enable more transparent reasoning. Future research should emphasise standardised benchmarks, hybrid mechanistic and human-centric methods, and integrating explainability into AI governance frameworks.
3 Problem Definition
This section introduces the fundamental concepts and mathematical framework necessary to understand our approach. We begin by defining key notation and problem settings, followed by a discussion of measuring the impact of perturbations, the surrogate model used for approximation, and a theoretical justification for its fidelity [13, 100, 101]. Finally, we outline a structured workflow for implementing a Wasserstein-based LIME surrogate, ensuring interpretability in the context of large language models [102].
3.1 Notation and Problem Setting
Setting the Stage.
In this study, we aim to develop a local surrogate model that explains the behaviour of a black-box model $\pi^{(n)}$ when provided with input text prompts. The basic notations are as follows:
  • Input: $x$, representing the original prompt.
  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, small modifications of $x$ used to explore the neighbourhood around the original input.
  • Model outputs: $\pi^{(n)}(y\mid x)$, the model’s probability distribution over possible outputs $y$ for the given input $x$.
3.2 Input-Level Distance
To measure how different each perturbed input $\hat{x}_{j}$ is from the original prompt $x$, we compute the semantic distance using Word Mover’s Distance (WMD) [103]:
$\delta_{x_{j}}:=\mathrm{WMD}(x,\hat{x}_{j}).$ (1)
Equation 1 quantifies the semantic dissimilarity between $x$ and $\hat{x}_{j}$, ensuring that closer perturbations are treated as more relevant.
3.3 Output-Level Distribution Shift
To assess how the model’s outputs change when the input is perturbed, we compute the Wasserstein distance between the output distributions for $x$ and $\hat{x}_{j}$ [101]:
$\Delta(x,\hat{x}_{j}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x}_{j})\right).$ (2)
Equation 2 captures the magnitude of the change in the model’s behaviour due to input perturbations.
3.4 Weighting via Input Similarity
Since not all perturbations are equally meaningful, we prioritise those that are semantically closer to the original input. We define a Gaussian-based weighting scheme [13]:
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right).$ (3)
As shown in Equation 3, perturbations with smaller semantic distances $\delta_{x_{j}}$ receive higher weights, reflecting their greater relevance.
3.5 Fitting a Local Surrogate
Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. The surrogate model $h_{\theta}$ adopts a linear form:
$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j}.$ (4)
The parameters $\theta$ are learned by minimising the weighted squared error between the surrogate’s predictions and the observed distribution shifts:
$\min_{\theta}\sum_{j}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2}.$ (5)
Equation 5 ensures that the surrogate closely approximates the black-box model’s behaviour within the local neighbourhood around $x$.
3.6 Theoretical Justification (Lipschitz Smoothness)
Shift Function Definition.
We formally define the shift function:
$f(x,\hat{x}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x})\right).$ (6)
Assuming that $\pi^{(n)}$ is Lipschitz continuous with respect to its input, the shift function $f$ is also Lipschitz [104]:
$|f(x,\hat{x})-f(x^{\prime},\hat{x}^{\prime})|\leq L\cdot\left(\|x-x^{\prime}\|+\|\hat{x}-\hat{x}^{\prime}\|\right),$ (7)
where $L>0$ is a Lipschitz constant (Equation 7). This smoothness justifies the use of a local linear surrogate model to approximate $f$.
Relation to Total Variation.
The Kantorovich-Rubinstein duality [105] relates the Wasserstein distance and Total Variation (TV) distance:
$W(p,q)\leq\text{diam}(\mathcal{Y})\cdot\mathrm{TV}(p,q),$ (8)
where:
$\mathrm{TV}(p,q)=\frac{1}{2}\sum_{y\in\mathcal{Y}}|p(y)-q(y)|.$ (9)
Equations 8 and 9 indicate that small changes in the input resulting in small TV distances also imply bounded Wasserstein shifts.
3.7 Linear Surrogates Approximate Lipschitz Functions
Due to the Lipschitz property, any smooth function $f$ can be locally approximated by a linear model in a small neighbourhood $U$:
$f(x^{\prime})\approx f(x)+\nabla f(x)\cdot(x^{\prime}-x).$ (10)
This principle, expressed in Equation 10, underlies the surrogate modelling approach adopted by LIME and related methods [13].
3.8 Putting It All Together
The complete Wasserstein-based LIME surrogate process consists of the following steps:
  1. 1.
    Generate perturbations $\{\hat{x}_{j}\}$ near $x$.
  2. 2.
    Compute semantic distances $\delta_{x_{j}}$ (Equation 1) and output-level shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
  3. 3.
    Assign weights $w_{j}$ using Equation 3.
  4. 4.
    Extract feature vectors $z_{j}$.
  5. 5.
    Fit a local surrogate model $h_{\theta}$ by minimising the loss (Equation 5).
  6. 6.
    Use $h_{\theta}$ to interpret how input perturbations affect the model’s output distribution.
This framework provides a rigorous, interpretable method for understanding how small changes to an input affect the outputs of complex models, particularly in applications involving large language models [102, 13, 101, 105].
4 PROPOSED METHOD
The proposed approach enhances the interpretability of large language models for text generation by analysing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behaviour [8, 102].
In the context of classification tasks, Fig. 4 illustrates SMILE, a tool designed to explain model predictions by isolating critical features that drive decision-making [12]. SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or logical sections for text, and creates perturbations by selectively modifying these components. These perturbations reveal how different input elements contribute to the final prediction.
Figure 4: SMILE flowchart for explaining image classification [12]
SMILE records the model’s predictions for each perturbed input and calculates similarity scores between the original and modified inputs using metrics such as the Wasserstein distance [100, 101]. These scores are transformed into weights using a kernel function, enabling a balanced weighting scheme. The results are used to train a weighted linear regression model that identifies the most influential input components.
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig. 5. Instead of perturbing the generated image, as in SMILE, we modify the instruction by including or excluding certain words. We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between these outputs and the one produced by the original text prompt [103]. This distance is a similarity measure, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
These similarity scores are then used as the outcome variable in a weighted linear regression model, with weights derived from the text distances and perturbations [106]. This regression model helps determine how each word affects the generated text. The coefficients derived from this model quantify the influence of each word, which we use to create a visual heatmap highlighting the most influential words in the text prompt. This heatmap provides an intuitive visual representation of the elements of the prompt that contribute the most to the generated output.
Using this enhanced three-step process, which builds on SMILE’s methodology [12], we aim to help users understand how particular elements of text prompts impact text generation. This technique improves user control and predictability in text generation models by providing insightful information about which elements of textual input, such as stylistic instructions or descriptive terms, most impact the final output.
The first step employs a text generation model based on input text. The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output. The text generation model is expected to produce responses similar to the original response that vary only based on critical components of the input texts [107].
Different text permutations are utilised from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into individual words. The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [108].
While the text generation model faces challenges in tasks such as counting objects and spatial reasoning, these issues can be mitigated by carefully selecting perturbation texts and restricting the input domain for perturbations [109]. The text generation models enable the proposed method to create responses that differ from the original text solely in terms of keywords related to the perturbation texts.
We use the Wasserstein distance to quantify these differences, comparing the distributions in both text embeddings [110, 100]. This metric is employed because it provides a more comprehensive measure of the differences between text embeddings compared to other ECDF-based distances, such as Kolmogorov–Smirnov or Cramér–von Mises. Although these methods focus on specific aspects, the Wasserstein distance captures both broad and subtle changes in the data [110, 101].
Finally, by applying linear regression, the space of input prompts is linked to the output space of the text generation model. As a result, the model’s behaviour can be interpreted by examining the regression coefficients [111, 106].
4.1 Text Generating and perturbed prompts
To begin, we generate variations of the original text prompt, known as “perturbed prompts”, to examine how these subtle changes affect the output of a text generation model [112, 113]. The original text is first broken down into individual words. Multiple text versions are then created by selectively including or excluding specific words. Each perturbed prompt is then provided to the text generation model, which produces a unique output for each variation. In this process, the output generated from the original prompt is the reference, or baseline output, against which all other outputs are compared.
For the original text prompt $p_{\text{org}}$, the text generation function $\phi_{\text{gen}}$ generates a corresponding output text $t_{\text{org}}^{\prime}$, as shown in Eq. 11:
$t_{\text{org}}^{\prime}=\phi_{\text{gen}}(p_{\text{org}})$ (11)
Similarly, for each perturbed prompt $p_{\text{pert}}^{i}$ within the set $S_{\text{pert}}$, the text generation model produces a corresponding output text $t_{\text{pert}}^{\prime i}$, according to Eq. 12:
$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in S_{\text{pert}}.$ (12)
Here, $p_{\text{org}}$ represents the original prompt, $p_{\text{pert}}^{i}$ describes a perturbation generated from the original text, and $\phi_{\text{gen}}$ is the text generation model function used to produce outputs $t^{\prime}$ that reveal the influence of prompt variations. The set of all perturbed prompts is denoted by $S_{\text{pert}}$ [13, 108].
4.2 Creating the interpretable space
In order to map the output space of the text generation model to a one-dimensional space, the Wasserstein distance between each text generated from the perturbed prompts and the original prompt (baseline output) is computed as a measure of similarity [100, 101]. In all cases, the input prompt remains consistent. Word Mover’s Distance (WMD) is employed to extract meaningful embeddings from the texts, allowing the focus to remain on the primary semantic components of the generated text [103]. This reduces noise in the distance estimation, enabling the accurately calculated distance to serve as the output for the linear regression model.
The distance between the outputs generated by the perturbation texts and the original prompt is given by Eq. 13:
$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{p}\right)^{\frac{1}{p}},\quad\forall i\in S_{\text{pert}}.$ (13)
In the above relation, the $\delta_{\text{WMD}}$ function represents the embedding extraction model applied to the texts before calculating the distance [103]. In Eq. 13, $n$ represents the number of features, $p$ denotes the norm order, and $W(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ demonstrates the Wasserstein distance between the outputs produced by the original and perturbed prompts.
Using p-values is crucial for validating the significance of the observed Wasserstein distances (WD) between texts generated under different prompt perturbations. By calculating a p-value, we test whether the observed WD could have occurred by chance, determining if the perturbation-induced changes are meaningful [114]. To compute this p-value, we use the Bootstrap algorithm [115, 116], which estimates the probability of observing a WD as large as the observed one. This involves generating distributions of WDs through repeated sampling and comparing each sampled WD with the observed value. If a small proportion of sampled WDs exceed the observed WD, the resulting p-value confirms that the differences are statistically significant, supporting the validity of the perturbation effects.
The Wasserstein-$p$ metric allows us to adjust WD’s sensitivity to different text variations by fine-tuning the norm order $p$. Smaller $p$ values emphasise local differences, while larger $p$ values capture broader distributional shifts [101]. By testing various $p$-values, we identify the norm that best captures meaningful variations caused by prompt changes, enhancing the reliability and interpretability of the results [101].
When analysing Wasserstein distances, measures with high p-values, indicating non-significant differences, are excluded from the procedure [114]. As described in Algorithm 1, the Bootstrap algorithm calculates these p-values for all relevant ECDF-based distance measures, including WD. The algorithm computes the WD and its associated p-value.
For a univariate example, let X and Y be the inputs. The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y. In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed. If boostWD exceeds the observed WD, a counter (bigger) is incremented. After completing all iterations, the p-value (pVal) is calculated as the proportion of iterations where boostWD is greater than the observed WD, as shown in Algorithm 1.
Input: Two sets of text-based distances: $X$ and $Y$
Output: Observed Wasserstein Distance (WD) and corresponding p-value (pVal)
1
[…]
6 $n\leftarrow\texttt{LX}+\texttt{LY}$;
7 $\texttt{bigger}\leftarrow 0$;
8
9for $i\leftarrow 1$ to MaxItr do
10       Draw a random sample $e$ of size LX from $XY$;
11       Draw a random sample $f$ of size LY from $XY$;
[…]
14             $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
15
16       end if
17
18 end for
[…]
Algorithm 1 Bootstrap Algorithm for Wasserstein Distance (WD) P-Value Calculation
Figure 5: SMILE flowchart for Large Language Models
When generating perturbation texts derived from the original text, the keywords of the original text play a crucial role in maintaining certain levels of similarity and fidelity between the perturbation texts and the source content. This is important when the interpretable model is trained on perturbed texts as input. Therefore, a similarity criterion is used to measure the degree of similarity between each perturbation text and the original, and this value is employed as a sample weight in the interpretable model. Eq. 15 calculates this similarity, and a Gaussian kernel is applied to normalise it into the range $[0,1]$ [103].
$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\,d(w_{k},w_{l})$ (14)
subject to $\sum_{l=1}^{m}T_{kl}=p_{k},\quad\forall k\in S_{\text{org}}$
$\sum_{k=1}^{n}T_{kl}=q_{l},\quad\forall l\in S_{\text{pert}}$
$T_{kl}\geq 0$
$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\exp\left(-\left(\frac{C_{i}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma^{2}}\right)^{2}\right),\quad\forall i\in S_{\text{pert}}$ (15)
The function $\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ quantifies the semantic distance between the perturbation text and the original. The resulting weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is obtained by applying a Gaussian kernel to this distance [106], allowing the surrogate model to assign higher importance to more semantically similar inputs.
4.3 Developing the Interpretable Surrogate Model
We are innovative in mapping the output space of the text generation model to a one-dimensional distance scale, accompanied by corresponding perturbation texts for the calculated distances [117]. This method enables us to train an interpretable model on the mapped output space of the text generation model, effectively demonstrating how input words influence text generation.
In our linear regression model, the vectors of the perturbation texts are critical, as they are treated as independent variables. Following the extraction of text embeddings, the distance between the outputs of these perturbation texts and the corresponding original text output is regarded as the response variable. Additionally [118], the degree of similarity between each perturbation text and the original text is assumed to represent the weight of the samples [102]. Eq. 16 outlines the square loss objective function for interpretable weighted linear regression:
$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=$ (16)
$\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\times\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$
Here, $\phi_{\text{gen}}$ represents the text generation model, and $\delta_{\text{WMD}}$ is the Word Mover’s Distance function applied to compute similarity between generated outputs. The weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is derived using a Gaussian kernel function and reflects the degree of similarity between the original prompt and each perturbed prompt. The term $\hat{f}_{t}(p_{\text{pert}}^{i})$ is a predicted value from the linear regression model.
After feature extraction, the $\hat{f}_{i}(p_{\text{pert}}^{i,01})$ function is a linear regression model trained over the one-hot vector space of perturbation texts $(p_{\text{pert}}^{i,01})$ and the calculated distances [102].
4.4 Evaluation Metrics
In our investigation, we adopt a suite of evaluation metrics inspired by foundational work provided by Google [119], emphasising the multifaceted nature of assessing explainable models. This work highlights the significance of metrics such as accuracy, stability, fidelity and consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models. The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [119].
Figure 6:
4.4.1 Accuracy
Accuracy measures how well the model’s output aligns with the expected results (ground truth). Specifically, we compare the model’s attribution scores (attention to text elements) against ground truth labels that identify the most relevant text elements for specific generated text segments [120].
To quantify this, we use the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) [121]. AUC reflects the model’s ability to rank relevant elements (from the ground truth) higher than irrelevant ones:
  • AUC $\sim$1: The model effectively distinguishes relevant elements, closely matching human-identified ground truth.
  • AUC $\sim$0.5: The model’s ranking is random, showing poor alignment with the ground truth.
For example, Fig. 6 shows that the ground truth identifies the words “meaning” and “life” as the most relevant text elements for generating the corresponding text output. The metric AttAUC evaluates the model’s alignment with the ground truth, where:
  • AttAUC = 1.0: Perfect alignment.
  • Lower AttAUC (0.8): Less accurate [122].
The heatmaps produced by the model visually represent these attention scores, as shown in Fig. 6, where darker red shades indicate higher attribution to specific text elements. An accurate model should correctly assign higher attention to the text elements that most influence the generated text output, improving the interpretability and reliability of the model [120].
4.4.2 Stability
Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [123]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. Stability is a key property for fostering trust in explainable artificial intelligence (XAI) systems [14]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [15].
In the example text, the attention scores are visualised for two slightly different text prompts. The heatmaps show how the model assigns attention to each word. A stable model, as demonstrated in Fig. 6, assigns similar importance to the words make and rainy in both prompts, even though the token  was added in one case. This consistency is critical for ensuring the model behaves predictably across minor input variations [16]. Without such stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [13].
We use the Jaccard index to quantify stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 17:
$\text{Jaccard}(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (17)
This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater stability and similarity in explanations [123]. A higher Jaccard index confirms the model’s ability to yield stable outputs, thus providing consistent and reliable interpretations [6]. By ensuring stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [14].
4.4.3 Consistency
In evaluating the consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times. Consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [13].
As shown in Fig. 6, when the input phrase “what is the meaning of life?” was fed into the model multiple times, the output remained stable across iterations. This consistency is crucial for applications where repeatable and reliable results are required. In our case, the model consistently assigns higher weights to the words “meaning” and “life”, demonstrating that it understands the importance of these words in generating the desired output [6]. Such behaviour underscores the model’s ability to focus on the semantic core of the input, a trait essential for maintaining coherence in practical use cases [15].
In practical applications—such as automated content generation, interactive systems, or scenario simulations—users depend on models to produce results that do not fluctuate unexpectedly. Any inconsistencies could undermine user trust and compromise the system’s overall effectiveness. Our model enhances user confidence by demonstrating consistent outputs and establishes a foundation for further optimisation and integration into more complex systems. This ability to maintain consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [14].
4.4.4 Fidelity
Fidelity is a metric to evaluate how well an explainable model aligns with a black-box model. The predictions of the black-box model, denoted as $f(x_{i})$, and the explainable model, $g(x_{i})$, for the $i$-th data perturbation are compared. Various loss functions and coefficient metrics quantify the similarity between the models’ predictions, measuring how effectively the explainable model approximates the black-box model.
The coefficient of determination $R^{2}$ is an accuracy-based metric that measures how well the explainable model correlates with the black-box model [124], and is defined in Eq. 18:
$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}(X_{i}))^{2}}$ (18)
where $\bar{f}(X_{i})$ represents the average of $f(X_{i})$. Values of $R^{2}$ close to 1 indicate a strong alignment between the two models.
For scenarios where weighting is necessary, the weighted coefficient of determination $R_{w}^{2}$ provides an alternative version [16, 125], defined as Eq. 19:
$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f_{w}}(X_{i}))^{2}}$ (19)
where $\bar{f_{w}}(X_{i})$ represents the weighted average of $f(X_{i})$. These accuracy-based metrics, $R^{2}$ and $R_{w}^{2}$, offer insight into how well the explainable model approximates the black-box model.
To account for both sample size and the number of features used in the local explanation, the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ [16] is used, as shown in Eq. 20:
$\hat{R}_{w}^{2}=1-(1-R_{w}^{2})\left[\frac{N_{p}-1}{N_{p}-N_{s}-1}\right]$ (20)
To measure the error or risk between the models’ predictions, error-based metrics, including the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE), are used [126]. These metrics quantify the weighted differences between the predictions of the explainable model and those of the black-box model. Lower values indicate better alignment. WMSE and WMAE are defined in Eqs. 21 and 22:
$\text{WMSE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}w_{i}}$ (21)
$\text{WMAE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot|f(X_{i})-g(X_{i})|}{\sum_{i=1}^{N_{p}}w_{i}}$ (22)
Here, $w_{i}$ denotes the weight for each data point $i$, $f(X_{i})$ represents the predictions from the black-box model, $g(X_{i})$ represents the predictions from the explainable model, and $N_{p}$ is the total number of data points.
Additionally, the mean $L_{1}$ and mean $L_{2}$ losses, defined below, are commonly used to assess prediction discrepancies [127, 16], as shown in Eqs. 23 and 24:
$L_{1}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (23)
$L_{2}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (24)
These error-based metrics, including MSE, MAE, $L_{1}$, and $L_{2}$, quantify the differences between the black-box and explainable models’ predicted scores, reflecting the risk of prediction discrepancies.
We compute fidelity across various scenarios. Fidelity is measured by comparing the predictions of the explainable model to those of the black-box text generation model. We analyse how different perturbations in the input text affect fidelity scores and how various distance metrics for text-to-text comparisons perform when using linear regression and Bayesian ridge as surrogate models.
5 Experimental Results
This section evaluates the proposed method’s ability to enhance explainability in instruction-based text generation. We analyse how the structure of textual prompts affects interpretability, how variations in input text influence the clarity of explanations, and how the proposed solution performs across diverse scenarios.
Experiments were conducted using a range of datasets and four state-of-the-art text generation models, OpenAI GPT [1], LLaMA [5], and Claude-AI [128] to ensure a comprehensive assessment.
5.1 Qualitative Results
The qualitative evaluation demonstrates the proposed framework’s interpretability using scenario-based testing and visualisations. We designed diverse testing scenarios by varying textual prompts, including changes in sentence complexity, rephrasing, and the inclusion of domain-specific terminology. These scenarios examine how the model interprets and generates outputs based on nuanced variations in textual inputs.
5.1.1 Input Prompt Explainability
The proposed framework applies scenario-based testing across multiple prompt variations to evaluate the interpretability of the model’s behaviour regarding different inputs. These variations include rephrased questions, descriptive modifiers, and changes in word choice. The system generates a text heatmap highlighting individual words’ contribution to the model’s output for each input, with colours indicating the importance scores.
Figure 7 shows how varying the input prompt’s phrasing or focus leads to different word importance patterns, demonstrating the model’s sensitivity to linguistic changes.
Figure 7: Text heatmaps showing word-level importance for various input prompts. The colour intensity represents the influence of each word on the model’s output.
5.1.2 Reasoning Path and Token Attribution
Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response. Figure 8 presents an example where the SMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram. This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models. As shown in Figure 8, SMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [129].
Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the SMILE framework. The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates SMILE’s token-level importance heatmap for the same prompt. [129]
5.1.3 Word-Level Attribution Analysis on MMLU Using SMILE
To evaluate the interpretability of ChatGPT using the SMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [130]. This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities. For this study, we accessed a pre-formatted version of the dataset available on Kaggle [131] to facilitate prompt construction and experimentation.
Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision. Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter. These visualisations demonstrate SMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
Figure 9: Attribution heatmaps generated by SMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
5.2 Quantitative Results
In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to asses our explainability method across different text generation models and different scenarios.
5.2.1 Accuracy
To evaluate the performance of instruction-based text generation models, we tested ten different prompts and their variations through 64 perturbations. Heatmaps were generated to display the weights of each keyword in the prompts across various models, including OpenAI GPT, LLAMA, and Claude-AI. A ground truth was defined, assigning weights of 1 to critical keywords in the prompts and 0 to less relevant or auxiliary words. These ground truth values were compared with the extracted heatmaps to assess the interpretability and accuracy of each model.
To quantify accuracy, we employed multiple metrics, including Attention Accuracy (ATT ACC), F1-Score for Attention (ATT F1), and Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC). The results were averaged across the ten different prompts, and each model’s performance is summarised in Table 1, showcasing their effectiveness based on these metrics.
Table 1: Performance metrics (ATT ACC, ATT F1, and ATT AUROC) for various
models, evaluated under two conditions: different prompts and different images.
Model name ATT ACC ATT F1 ATT AUROC
OpenAI-GPT 0.70 0.59 0.84
LLaMA 0.76 0.40 0.84
Claude-AI 0.82 0.67 0.88
5.2.2 Stability
To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap. For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word. To evaluate stability, we added the token ‘‘ at the end of each prompt as input text, extracted the coefficients again, and computed the Jaccard index to compare the sets of coefficients before and after the perturbation. Finally, we calculated the average Jaccard index for each model, and the results are presented in Table 2.
Table 2: Stability across different models for ten prompts and 30 perturbations
Model name Jaccard Index
OpenAI-GPT 0.62
LLaMA 0.45
Claude-AI 0.44
5.2.3 Consistency
To compute consistency, we ran the same code with 64 perturbations for each prompt over 10 iterations, ensuring that the outputs remained consistent across these repetitions for text generation models such as OpenAI GPT, LLAMA, and Claude-AI, as shown in Table 3. We computed variance and standard deviation metrics for each word coefficient. This step is essential to confirm that the model’s predictions are not random or heavily dependent on initialisation factors but are instead deterministic and robust.
Table 3: Consistency metrics for different models on the prompt
“What is the meaning of life?”
Model name Variance Standard Deviation
OpenAI-GPT 0.0000 0.0046
LLaMA 0.0048 0.0663
Claude-AI 0.0011 0.0331
5.2.4 Fidelity for Different Instruct image editing diffusion based Models (IED)
[…]
Table 4: Fidelity metrics for different models on the prompt
“What is the meaning of life?”
[…]
OpenAI-GPT 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
LLaMA 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
Claude-AI 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
5.2.5 Fidelity Across Different number of Text Perturbations
Tables 56, and 7 present the results of fidelity computations for varying numbers of perturbations across the OpenAI GPT, LLAMA, and Claude-AI models, respectively. These results are derived using a weighted linear regression as the surrogate model and the Wasserstein distance for computing distances. The evaluation metrics include weighted mean squared error (WMSE), weighted mean absolute error (WMAE), coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$), and average loss metrics such as mean-L1 and mean-L2.
Table 5: Performance metrics for different numbers of perturbations in OpenAI GPT
[…]
32 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
64 0.0617 0.3406 0.1695 0.1756 0.0526 0.2712
128 0.0329 0.4468 0.1385 0.1519 0.0378 0.4193
256 0.0116 0.4276 0.0675 0.0872 0.0163 0.4138
Table 6: Performance metrics for different numbers of perturbations in LLAMA
[…]
32 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
64 0.0240 0.4844 0.1284 0.1433 0.0382 0.4301
128 0.0085 0.2430 0.0638 0.0612 0.0096 0.2055
256 0.0058 0.2387 0.0605 0.0667 0.0118 0.2203
Table 7: Performance metrics for different numbers of perturbations in Claude-AI
[…]
32 0.0368 0.7209 0.1495 0.2262 0.0733 0.6539
64 0.0115 0.3708 0.0564 0.1036 0.0370 0.3046
128 0.0089 0.1506 0.0535 0.0920 0.0192 0.1085
256 0.0191 0.2670 0.0902 0.0873 0.0163 0.2494
5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
[…]
Table 8: Fidelity results for different distance measures with 30 perturbations
WLR Fidelity Metrics
[…]
Cosine Cosine 0.0172 0.3151 0.0659 0.1277 0.0412 0.1508
Cosine WD 0.0216 0.4197 0.0899 0.1332 0.0385 0.2805
WD WD 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
WD Cosine 0.0048 0.4026 0.0329 0.0871 0.0296 0.2593
WD+C WD+C 0.0349 0.5349 0.1589 0.3050 0.1468 0.4233
BayLIME
Cosine Cosine 0.0200 0.2285 0.0499 0.1067 0.0464 0.0434
Cosine WD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
WD WD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
WD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
WD+C WD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
5.2.7 Computation complexity
We apply our model to three instruction-based text generation frameworks: OpenAI GPT, LLaMA, and Claude-AI. We also evaluate the execution time of different explainability methods: LIME, SMILE, and Bay-LIME. Each method is tested with 60 perturbations on the same hardware to ensure consistency in computational conditions.
The models exhibit notable differences in execution time, primarily due to how each explainability method computes similarity and assigns weights. SMILE, in particular, employs the Wasserstein distance to quantify text dissimilarities, making it substantially more computationally expensive than both LIME and Bay-LIME. As reported in Table 9, SMILE consistently incurs higher execution times across all three model frameworks. This pattern also holds across other frameworks, indicating that SMILE’s computational overhead is a systemic characteristic of its underlying methodology.
Wasserstein distance is known for capturing subtle differences between text embeddings, making it especially effective in high-dimensional semantic spaces [101]. However, this precision comes at a cost. Computing the Wasserstein-2 distance involves operations on covariance matrices with time complexity proportional to $O(d^{3})$, where $d$ is the embedding dimensionality. When applied across $N$ samples, the total complexity becomes $O(Nd^{3})$. In contrast, LIME relies on cosine similarity, which has a much lower computational cost of $O(Nd)$.
This disparity arises because Wasserstein-2 distance requires solving matrix optimisation problems over $d$-dimensional Gaussian distributions [132]. While this added complexity leads to more stable and expressive explanations, as demonstrated by SMILE, it significantly increases runtime. Conversely, LIME avoids this overhead by utilising simpler vector-based similarity measures.
In summary, SMILE offers greater fidelity and stability but at a higher computational cost. Its use of the Wasserstein distance makes it ideal for tasks where precision and interpretability are essential, even if that entails longer execution times. LIME and Bay-LIME, while faster, may trade off some robustness in favour of efficiency.
Table 9: Execution times (in seconds) for each framework and explainability method with 60 perturbations
Framework LIME SMILE Bay-LIME
OpenAI-GPT 156.82 170.70 161.43
LLaMA 113.00 118.99 116.40
Claude-AI 355.09 372.95 349.88
5.2.8 Optimised Instructions from Google Research
To further explore how instruction phrasing influences interpretability, we applied SMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [133]. These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
We first applied SMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem. ” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, SMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using SMILE with GPT-3.5.
Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using SMILE with GPT-4.
5.2.9 Token-Level Attention for Different Sentence Structures
To investigate how sentence structure affects token-level attention in large language models (LLMs), we analyzed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts. These prompts ranged from questions (e.g., “What is the meaning of life?”), commands (e.g., “Please give me the meaning of life”), to declarative formulations (e.g., “You must explain the meaning of life”).
For each sentence, we extracted the attention values directly from the final layer of the LLM. We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualized using a box plot (see Figure 12), where each point represents a unique sentence. Full-sentence prompts are viewable through hover text, allowing qualitative comparison of sentence structure with attention distribution.
The analysis reveals noticeable variability in how these tokens are weighted depending on syntactic form. For instance, in imperative constructions, attention to “life” tends to dominate, whereas in interrogative or abstract prompts, “meaning” often receives comparable or higher attention. This suggests that sentence formulation can subtly steer the internal focus of the model, even when the semantic content remains consistent.
These findings underscore the sensitivity of LLMs to surface-level structure in shaping token-level interpretability, offering insight into both model behaviour and prompt engineering strategies.
Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures. Hovering over each point reveals the full sentence associated with that score.
6 Limitations
While SMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, even when the semantic meaning remains largely intact. This sensitivity challenges the assumption of local smoothness required for regression-based attribution and can reduce explanation stability.
2) Some LLMs exhibit hallucination, producing fluent but factually incorrect or irrelevant outputs, especially when presented with ambiguous or syntactically degraded perturbations. This introduces noise into the attribution signal and affects explanation reliability.
3) The computational cost of generating and processing many perturbations (e.g., 1024 per input) is significant. Commercial APIs can lead to high inference costs and latency. For local models, substantial GPU resources are required, limiting practical scalability.
4) Some models, particularly research-stage or open-source LLMs, do not provide API access and require manual local deployment.
7 Conclusion
Interpretability remains a cornerstone of responsible AI, particularly in the context of large language models (LLMs) and instruction-based text generation. As AI systems become increasingly embedded in domains such as business, education, healthcare, and public services, it becomes essential for users to understand how these models generate outputs from their prompts. Without such insight, trust in AI systems diminishes, limiting adoption in real-world workflows. This underscores the need for transparent and interpretable solutions to demystify the behaviour of these black-box models.
Interpretability is not only a research priority but is likely to become a legal and operational requirement in alignment with emerging AI regulations and ethical standards. Organisations that fail to address interpretability may encounter regulatory hurdles and reduced user acceptance.
This work introduced a novel interpretability framework tailored for instruction-based text generation models. Leveraging SMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output. Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
Whereas most existing interpretability methods highlight important tokens in the model’s output [13, 134], SMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that SMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma. The resulting heatmaps and attribution scores intuitively understand model behaviour, empowering users to control better and refine prompt design.
This work lays the groundwork for more transparent and trustworthy natural language systems. Future directions include extending SMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
8 Future Works
  1. 1.
    Extension to Mathematical and Equation-Based Text Generation: Future work will explore the application of the proposed interpretability framework to models focused on symbolic reasoning and math-related tasks, such as equation solving, theorem generation, and formal derivation. Instruction-based mathematical models like Minerva and MathGLM require high precision in prompt design, and understanding how prompt components affect symbolic output will be essential in educational and scientific settings.
  2. 2.
    Application to Paraphrase Generation Models: The methodology can be extended to paraphrasing models, where semantic preservation and variation are central. Analysing how individual words influence paraphrased outputs’ style, tone, or fidelity can support better user control in writing assistance, content rephrasing, and translation tools.
  3. 3.
    Adaptation to Multi-Turn and Context-Aware Generation: Another direction is to adapt the framework for multi-turn and context-aware systems, including dialogue agents. This would involve tracking how prompt changes impact responses’ consistency, coherence, and relevance across conversational sequences.
  4. 4.
    Cross-Task Generalisation: While this work focuses on instruction-based text generation, the interpretability strategy can be generalised to broader NLP tasks, including summarisation, classification, and question answering. Evaluating how different model families respond to prompt perturbations can uncover model-specific behaviours and improve transparency.
  5. 5.
    Benchmarking Interpretability and Prompt Sensitivity: We plan to contribute to standardised evaluation protocols for prompt sensitivity, attribution robustness, and local fidelity. These benchmarks will support consistent assessment across models and help identify best practices for interpretable generation.
Data Availability
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the SMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.

### v2

@@
-Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations
+Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations
 Abstract
-Large language models like GPT, LLAMA, and Claude have become incredibly powerful at generating text, but they are still black boxes, so it is hard to understand how they decide what to say.
-That lack of transparency can be problematic, especially in fields where trust and accountability matter.
-To help with this, we introduce SMILE, a new method that explains how these models respond to different parts of a prompt.
-SMILE is model-agnostic and works by slightly changing the input, measuring how the output changes, and then highlighting which words had the most impact.
-Create simple visual heat maps showing which parts of a prompt matter the most.
-We tested SMILE on several leading LLMs and used metrics such as accuracy, consistency, stability, and fidelity to show that it gives clear and reliable explanations.
-By making these models easier to understand, SMILE brings us one step closer to making AI more transparent and trustworthy.
+Large Language Models (LLMs) such as GPT, LLaMA, and Claude have achieved remarkable performance in text generation, yet their decision-making processes remain opaque.
+This lack of transparency poses challenges in domains where trust and accountability are critical.
+In this work, we introduce gSMILE (Generative SMILE), a generalized and LLM-oriented extension of the previously proposed SMILE framework, for interpreting how these models respond to specific parts of an input prompt.
+gSMILE is model-agnostic and operates by introducing controlled perturbations to the input, observing the resulting output changes, and identifying which input words have the most significant impact on the output.
+This process enables the generation of intuitive heatmaps that visually highlight influential words in the prompt.
+We evaluate gSMILE across several leading large language models (LLMs) using metrics such as accuracy, consistency, stability, and fidelity, demonstrating its ability to provide precise, robust, and reliable explanations.
+By making LLMs more interpretable, gSMILE takes an important step toward building more transparent and trustworthy AI systems.
 Figure 1: Highlighting the contribution of each word in the input prompt of a Commercial LLM
 1 Introduction
-Large Language Models (LLMs) have become powerful tools in recent years.
-They can perform tasks ranging from answering questions with remarkable accuracy [1] to drafting essays and other written content with coherence and style [2], creating art that spans various mediums and styles [3], coding across multiple programming languages [4], debugging software [4], writing technical documentation [5], and even generating algorithms [1].
-LLMs function by leveraging deep learning architectures, particularly the Transformer model.
-This model uses self-attention mechanisms to process and generate text by analysing relationships between words in large data sets [6].
-These models are trained in various corpora, including books, articles, code, and web content, allowing them to learn language patterns, context, and syntax on an unheard scale [7].
-Their capabilities continue to grow, turning them into essential tools for a wide range of tasks [2].
-These models, trained in various texts, have transformed how we interact with technology and how machines understand and generate human language [6].
-LLMs such as GPT, BERT, and Claude leverage the Transformer architecture to process and generate language based on patterns learned from extensive datasets [7].
-Although they excel in many NLP tasks, understanding their strengths and limitations is the key to using them effectively and responsibly [8].
-These models continue to evolve, pushing the boundaries of AI applications in language understanding and generation [2].
-However, the nature of these models is known as black boxes and lacks transparency [9].
-One of their important problems is how these models decide which words or parts of input to focus on, which remains a complex and fascinating process [10].
-At their core, LLMs rely on patterns and relationships learned during training to decide which input elements are the most important [6].
-This process is more than just identifying individual keywords; it is about picking up sophisticated connections and context that shape the model’s responses [10].
-Understanding how this works is crucial: it explains why these models excel in some scenarios but struggle in others.
-It also highlights why they can occasionally produce responses that are biased, irrelevant, or even harmful [11].
-By studying how LLMs emphasise words and phrases, we can better understand how they work [10].
-This can help us make them more transparent and reliable, ensuring they focus on what truly matters in a given context.
-It also allows us to address potential problems, such as biases or alignment problems, making these tools safer and more ethical [11].
-As LLMs continue to integrate into our daily lives and decision-making processes, understanding their inner workings is more important than ever [9].
-Research on how they prioritise language helps us refine their performance, ensure ethical use, and unlock their full potential to benefit society [11] responsibly.
-To solve this problem, we propose a novel approach of SMILE (Statistical Model-agnostic Interpretability with Local Explanations) to extract a visual heat map that shows the influence of each word in responding to these models [12].
-This method extends the foundation laid by LIME [13] by incorporating Empirical Cumulative Distribution Function (ECDF) statistical distances, significantly improving robustness and making interpretability more effective.
-SMILE produces visual heatmaps that reveal the impact of individual words or phrases in text input on the model’s output as illustrated in Fig 1.
-For example, when a language model processes an instruction or query, the heatmap highlights the importance of specific terms in influencing the outcome.
-This offers users a clear visual representation of how the model interprets and prioritises different parts of the input text [10].
-To ensure a thorough evaluation of our approach, we employ a suite of metrics, including accuracy, fidelity, stability, and consistency [14].
-These metrics demonstrate that SMILE provides reliable insights into model behaviour [15].
-The contributions of our work include the following.
-  • Introducing a New Interpretability Framework: SMILE is a cutting-edge method that visualises the role of individual words in shaping model output, offering a deeper understanding of how LLMs process language [12].
-  • Promoting Transparency and Trust in LLMs: By visually linking text input to model decisions, SMILE enhances transparency, making LLMs more trustworthy and enabling their practical use in high-stakes domains such as healthcare, education, and legal systems.
-  • Comprehensive Evaluation for Robustness: Our evaluation leverages metrics such as stability, fidelity, accuracy, and consistency, establishing a new standard for evaluating the interpretability and reliability of language models [16].
+Large Language Models (LLMs), such as GPT, Claude, and LLaMA, have significantly advanced the capabilities of natural language processing systems [1, 2, 3].
+These models, based on the Transformer architecture [4], are trained on extensive corpora and exhibit remarkable performance across a wide range of tasks, including text generation, summarisation, code completion, and dialogue reasoning [5].
+However, despite their widespread adoption, LLMs remain largely opaque in how they make predictions, leading to concerns about transparency, trust, and safety, especially in high-stakes domains such as healthcare and law [6, 7].
+The black-box nature of LLMs limits our ability to understand which parts of an input prompt have the most significant influence on the model’s output.
+This lack of interpretability hinders debugging, ethical assessment, and responsible deployment [8, 9].
+Interpretability methods, especially those that operate post hoc and in a model-agnostic fashion, offer a way to open these black boxes by explaining the model’s behavior without requiring access to internal parameters [10].
+In this work, we propose gSMILE, a generalized and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [11].
+gSMILE interprets model outputs by perturbing input prompts, measuring output shifts using Wasserstein distance, and fitting a local linear surrogate model.
+The result is a weighted heatmap that visualizes how individual words in a prompt contribute to the model’s response.
+We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including fidelity, consistency, stability, and accuracy.
+Our findings demonstrate that gSMILE yields stable, interpretable, and reliable attributions, rendering it suitable for practical applications in real-world scenarios involving instruction-based generation.
+The key contributions of this paper are:
+  • Token-level interpretability for LLMs: We present gSMILE, a model-agnostic method that quantifies the influence of each input token on the output of large language models using perturbation-based analysis and Wasserstein distance.
+  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behavior.
+  • Robust cross-model evaluation: We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, accuracy, fidelity, stability, and consistency that demonstrate its reliability and adaptability.
 2 Literature review
-This section summarises the most recent advances in large language models (LLMs) and their growth in various applications.
-We also discuss Explainable Artificial Intelligence (XAI), a field dedicated to making AI systems more transparent and easier to understand.
-Finally, we explore how these two areas come together, focusing on the ongoing efforts to create explainable large language models.
-We discuss the key challenges involved and the techniques being developed to improve how we interpret and trust these powerful models.
+This section provides an overview of recent advancements in large language models (LLMs) and their increasing impact across various applications.
+We also introduce Explainable Artificial Intelligence (XAI), which aims to make AI models more transparent, interpretable, and trustworthy.
+As LLMs grow in complexity and usage, the need to understand their decision-making processes becomes more critical.
+Therefore, we highlight the emerging efforts to develop explainable large language models (LLMs), outlining key challenges and recent techniques that aim to improve their interpretability and foster greater trust in their outputs.
 2.1 Large Language Models
-Large language models (LLMs) have become increasingly common thanks to significant strides in deep learning, better access to data, and more powerful computing resources.
-These models are trained on vast amounts of text and can generate responses that sound natural, answer questions, and support a wide range of language-based tasks.
-Over time, several notable LLMs have helped push the field forward, each bringing new ideas and improvements.
-GPT (OpenAI) [1] is one of the most well-known autoregressive transformer-based models, capable of generating coherent and contextually rich text.
-Google’s T5 [17] introduced the text-to-text framework, where all NLP tasks are framed as text generation problems, improving transfer learning and efficiency.
-Similarly, PaLM (Google) [18] leverages the Pathways architecture for large-scale language modelling, optimising multitask learning across different NLP tasks.
-Anthropic’s Claude [19] focuses on ethical AI and safe interactions designed to minimise harmful outputs while maintaining interpretability.
-Google’s Gemma 2 [20] is optimised for logical reasoning and structured data interpretation, making it a strong candidate for academic and enterprise use cases.
-Meta’s LLaMA [5] is an efficient, open-source model tailored for various NLP applications, offering researchers an accessible yet powerful alternative to proprietary models.
-EleutherAI’s GPT-NeoX [21] is an open-source alternative to GPT, providing robust text generation capabilities while allowing customisation.
-AI21 Labs’ Jurassic-2 [22] is designed for large-scale enterprise applications, emphasising contextual understanding and user intent.
-Cohere’s Command R+ [23] integrates retrieval-augmented generation, improving fact-based responses and minimising hallucinations.
-Furthermore, models like BERT (Google) [24] have NLP through bidirectional context understanding, significantly improving sentiment analysis and question-answering tasks.
-Microsoft’s Phi-4 focuses on efficiency and cost-effectiveness, delivering competitive performance with fewer parameters.
-Mistral and Alibaba’s Qwen 2.5 push the boundaries of model adaptability and fine-tuning capabilities, making them suitable for multilingual and domain-specific applications.
-DeepSeek-LLM [25] is another promising open-source model, competing with top-tier models regarding benchmark performance.
-Models such as Mixtral and StableLM have been developed with efficiency and specific NLP task performance in mind, ensuring a balance between accuracy and computational demands.
-Recent developments have introduced models such as DeepSeek-R1, which focused on advanced reasoning capabilities, and Qwen2.5-Max, which explores large-scale Mixture-of-Experts (MoE) models to enhance efficiency and performance.
-OpenAI’s o3-mini pushes the frontier of cost-effective reasoning, balancing performance with computational efficiency.
-DeepSeek-V3, the first open-sourced GPT-4o level model, represents a significant step forward in open-source LLM development.
-The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications:
-Category Models Description General-Purpose GPT-4 [26], GPT-4o [27], Claude 3 [28], Gemini 1.5 [29], Command R+ [30], Mistral [31] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [32], Mistral [31], Mixtral [33], Falcon 180B [34], DeepSeek-V3 [35], OpenChat [36] Customizable models for research applications Retrieval-Augmented Command R+ [30], GPT-4 Bing [26], Claude 3 RAG [28], Gemini Search [29] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [28], Gemini 1.5 [29], GPT-4o [27], Grok-1.5 [37] Specialized in processing long-form text Advanced Reasoning GPT-4 [26], Claude 3 Opus [28], Gemini 1.5 [29], Mixtral [33], Grok-1.5 [37] Logical reasoning, multi-task learning, problem-solving Ethically-Aligned Claude 3 [28], Gemini [29], GPT-4 [26], Grok [37] Emphasize safety, reducing bias, ethical AI Mixture-of-Experts (MoE) Mixtral [33], Qwen2.5-Max [38], GShard [39], Switch Transformer [40] Improved efficiency and performance via MoE Multimodal GPT-4o [27], Gemini 1.5 [29], Claude 3 [28], Kosmos-2 [41], Grok-1.5V [37] Capable of processing both text and visual inputs Multilingual BLOOM [42], XGLM [43], mBERT [44], XLM-R [45], ByT5 [46] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [47], Galactica [48], Legal-BERT [49], StarCoder2 [50], Code LLaMA [51] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [52], Vicuna [53], ChatGLM [54], Baize [55], OpenChat [36] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [56], TinyLLaMA [57], DistilGPT2 [58], GPT-2 [59], LLaMA 2 7B [60] Designed for fast inference and deployment on edge devices
+Large language models (LLMs) have become increasingly prevalent due to significant advancements in deep learning [12], improved access to large-scale datasets, and the growing availability of powerful computing resources [13].
+These models are typically trained on massive corpora of text.
+They are capable of generating fluent, contextually appropriate responses across a wide variety of language-based tasks, including question answering, summarisation, translation, and content generation.
+Over time, several influential large language models (LLMs) have helped push the field forward, each introducing novel architectural innovations [4], training strategies [14], or alignment techniques [15] that have contributed to their growing capabilities and broader adoption.
+Among the most influential models, GPT (OpenAI) [1] stands out as a leading autoregressive transformer-based architecture capable of generating coherent and contextually rich text.
+It set a new standard for generative language models and introduced the broader public to powerful applications of AI through tools like ChatGPT.
+Expanding on this paradigm, Google’s T5 [14] reframed all NLP tasks as text-to-text problems, streamlining the transfer of learning across diverse tasks and contributing to more unified and efficient architectures.
+Palm (Google) [16], with its Pathways design, extended the multitask learning capability at scale, allowing a single model to generalize effectively across different NLP applications.
+In parallel, Anthropic’s Claude [17] introduced a new approach centered on safe and aligned AI behavior.
+Claude utilizes constitutional AI principles to ensure its outputs are consistent with predefined ethical standards, thereby promoting greater transparency and reducing the risk of harm.
+More recently, Google’s Gemma 2 [18] has focused on logical reasoning and structured data interpretation, offering a competitive option for research-driven and enterprise-level deployment.
+Meta’s LLaMA [19] series continues to draw attention as a scalable, open-source alternative, providing the research community with accessible tools that deliver high performance across a range of benchmarks.
+Community-driven efforts have also enriched the landscape.
+EleutherAI’s GPT-NeoX [20] was developed to provide an open-source alternative to proprietary GPT models, offering researchers greater flexibility in experimentation and deployment.
+AI21 Labs’ Jurassic-2 [21] targets commercial-scale applications with a focus on contextual awareness and enterprise integration.
+Meanwhile, Cohere’s Command R+ [22] integrates retrieval-augmented generation (RAG), enhancing factual consistency and minimizing hallucinations in the generative output — a key concern for the practical adoption of large language models (LLMs).
+Additionally, transformer-based models, such as BERT [23], have revolutionized NLP through bidirectional encoding, providing substantial improvements in tasks like question answering, named entity recognition, and sentiment analysis.
+Efficiency-focused models, such as Microsoft’s Phi-4, highlight a growing trend toward parameter-efficient design without sacrificing core performance, making large language models (LLMs) more accessible for low-resource environments.
+Mistral and Qwen 2.5 (by Alibaba) continue this direction by offering fine-tuning-friendly models with strong multilingual and domain adaptation capabilities.
+Several recent open-source contenders further enrich the ecosystem.
+DeepSeek-LLM [24] delivers performance comparable to proprietary models while remaining open and extensible.
+Mixtral and StableLM focus on task efficiency and targeted performance, ensuring their utility in production environments.
+DeepSeek-R1 has been noted for its emphasis on reasoning, reflecting a broader shift from general language generation to models optimized for critical thinking, chain-of-thought prompting, and interpretability.
+Cutting-edge advancements have introduced even more specialized architectures.
+Qwen2.5-Max utilizes large-scale Mixture-of-Experts (MoE) models to enhance throughput and scalability.
+OpenAI’s o3-mini emphasizes low-latency, cost-effective deployment while retaining high-quality output — ideal for mobile and edge computing.
+Perhaps most notably, DeepSeek-V3 represents a significant milestone by achieving GPT-4o-level performance in an open-source setting, suggesting that open-source models may soon rival or surpass closed models in both capability and adoption.
+The following models in Fig.
+2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications.
+These models continue to define the state of the art in machine-generated language, with a growing focus on transparency, efficiency, and alignment with human values.
+Category Models Description General-Purpose GPT-4 [25], GPT-4o [26], Claude 3 [27], Gemini 1.5 [28], Command R+ [29], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [32], Falcon 180B [33], DeepSeek-V3 [34], OpenChat [35] Customizable models for research applications Retrieval-Augmented Command R+ [29], GPT-4 Bing [25], Claude 3 RAG [27], Gemini Search [28] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [27], Gemini 1.5 [28], GPT-4o [26], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [25], Claude 3 Opus [27], Gemini 1.5 [28], Mixtral [32], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [27], Gemini [28], GPT-4 [25], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [32], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [26], Gemini 1.5 [28], Claude 3 [27], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [43], XLM-R [44], ByT5 [45] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [46], Galactica [47], Legal-BERT [48], StarCoder2 [49], Code LLaMA [50] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [51], Vicuna [52], ChatGLM [53], Baize [54], OpenChat [35] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [55], TinyLLaMA [56], DistilGPT2 [57], GPT-2 [58], LLaMA 2 7B [59] Designed for fast inference and deployment on edge devices
 Figure 2: Extended LLM Categories, Models, and Descriptions
 2.2 Explainable AI (XAI)
-The increasing complexity of machine learning models has made understanding how they make decisions more important than ever.
+The increasing complexity of machine learning models has made understanding how they make decisions more critical than ever.
 Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [60, 61].
-Explainability methods can be broadly categorised as intrinsic or post-hoc.
+Explainability methods can be broadly categorized as intrinsic or post-hoc.
 Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62].
@@
 Post-hoc explainability methods can be classified into global and local explanations.
-Global explanations provide an overarching view of the model’s decision-making patterns across an entire dataset, helping to identify trends and biases [62, 64].
+Global explanations offer an overarching view of the model’s decision-making patterns across an entire dataset, facilitating the identification of trends and biases [62, 66].
 In contrast, local explanations focus on specific predictions, providing a detailed breakdown of how individual inputs influence model outputs.
 LIME and its numerous enhancements fall under local explainability methods and are designed to provide granular, case-specific interpretations [63].
-LIME has become one of the most widely used explainability techniques, but over time, researchers have introduced various improvements to address its limitations and expand its usefulness across different applications.
-As shown in Figure 3, enhancements to LIME can be grouped into four main categories:
+LIME has become one of the most widely used explainability techniques; however, over time, researchers have introduced various improvements to address its limitations and expand its applicability across different applications.
+As shown in Figure 3, enhancements to LIME can be grouped into four main categories.
 A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points.
@@
 Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives.
-Methods such as Q-LIME [65], S-LIME [66], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behaviour, capturing non-linear relationships that the original method might miss.
-Furthermore, researchers have explored optimisation strategies to make LIME more efficient and stable.
+Methods such as Q-LIME [67], S-LIME [68], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behavior, capturing non-linear relationships that the original method might miss.
+Furthermore, researchers have explored optimization strategies to enhance the efficiency and stability of LIME.
 Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
@@
 Adaptations such as PointNet LIME for 3D point cloud data [69], TS-MULE for time series data [70], Graph LIME for graph-structured data [71], Sound LIME for audio data [72], and B-LIME for ECG signal data [73] demonstrate how LIME can be tailored to diverse applications.
-In addition, several modifications have been proposed to improve LIME’s approach, including using alternative surrogate models, fine-tuning distance parameters, optimising sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
-Figure 3: Key strategies for enhancing LIME, including improvements in sampling, distance parameters, surrogate models, and optimisation.
+In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimizing sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
+Other extensions include DSEG-LIME [74], which tailors LIME for segmentation models by generating perturbations at the object level rather than the pixel level, improving interpretability in image segmentation tasks.
+SLICE [75] introduces structured sampling and clustering to ensure diverse and representative local explanations, reducing redundancy in LIME samples.
+Stratified LIME [76] improves fidelity by stratifying the input space and generating more targeted local samples for different subpopulations.
+US-LIME [77] employs uncertainty sampling to guide instance selection, increasing the reliability and faithfulness of local explanations.
+SS-LIME [78] uses self-supervised learning to identify important regions in the feature space and provide richer explanations.
+K-LIME [79] incorporates kernel-based methods into LIME’s regression step to model more complex local decision boundaries.
+Attention-LIME [80] leverages attention mechanisms to prioritize features during the explanation, aligning more closely with deep learning models.
+BMB-LIME [81] employs a boosted multi-branch approach to combine multiple local models, thereby enhancing consistency and robustness.
+These methods aim to enhance interpretability across various domains, including segmentation, survival analysis, multimodal learning, and attention-based contexts.
+Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [82], LIME-SUP, LORE, ILIME, Aleph LIME, bLIMEy, NormLIME, DLIME, LIMEtree, MPS-LIME, Explain Explore, survLIME, survLIME-inf, CBR-LIME, GuidedLIME, audioLIME, MeLIME, C-LIME, gLIME, DIME, s-LIME, LIMESegment, UnRAvEL-LIME, Kernel-LIME, Specific-Input LIME, ExpLIMEable, GLIME, GMM-LIME, SEGAL, and Sig-LIME, reflecting the diverse and rapidly evolving landscape of post-hoc explainability [83].
+Figure 3: Taxonomy of LIME enhancement strategies: categorized into sampling techniques, surrogate model improvements, distance parameter adjustments, and optimization approaches.
 2.3 Explainable AI (XAI) for Large Language Models
@@
 3 Problem Definition
-This section introduces the fundamental concepts and mathematical framework necessary to understand our approach.
-We begin by defining key notation and problem settings, followed by a discussion of measuring the impact of perturbations, the surrogate model used for approximation, and a theoretical justification for its fidelity [13, 100, 101].
-Finally, we outline a structured workflow for implementing a Wasserstein-based LIME surrogate, ensuring interpretability in the context of large language models [102].
+This section introduces the core concepts and mathematical framework underlying our approach.
+We begin by defining the key notation and problem setup, followed by a formal description of how we quantify the impact of input perturbations.
+We then describe the use of a surrogate model to approximate model behavior locally and provide theoretical justification based on the Wasserstein distance [63, 112, 113].
+Lastly, we outline a structured methodology for implementing a Wasserstein-based surrogate framework tailored to large language models [114].
 3.1 Notation and Problem Setting
 Setting the Stage.
-In this study, we aim to develop a local surrogate model that explains the behaviour of a black-box model $\pi^{(n)}$ when provided with input text prompts. The basic notations are as follows:
-  • Input: $x$, representing the original prompt.
-  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, small modifications of $x$ used to explore the neighbourhood around the original input.
-  • Model outputs: $\pi^{(n)}(y\mid x)$, the model’s probability distribution over possible outputs $y$ for the given input $x$.
+We aim to construct a local surrogate model that explains the behavior of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt.
+The goal is to understand how specific parts of the input affect the model’s output distribution.
+The following notation is used:
+  • Input: $x\in\mathcal{X}$, representing the original input prompt.
+  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, a set of $J$ perturbed versions of $x$, generated via minor token-level edits.
+  • Model Outputs: $\pi^{(n)}(y\mid\hat{x}_{j})$, the model’s probability distribution over output space $\mathcal{Y}$, conditioned on the perturbed input $\hat{x}_{j}$.
+For the unperturbed input $x$, we write $\pi^{(n)}(y\mid x)$.
 3.2 Input-Level Distance
-To measure how different each perturbed input $\hat{x}_{j}$ is from the original prompt $x$, we compute the semantic distance using Word Mover’s Distance (WMD) [103]:
-$\delta_{x_{j}}:=\mathrm{WMD}(x,\hat{x}_{j}).$ (1)
-Equation 1 quantifies the semantic dissimilarity between $x$ and $\hat{x}_{j}$, ensuring that closer perturbations are treated as more relevant.
+To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [115]:
+$\delta_{x_{j}}=\mathrm{WMD}(x,\hat{x}_{j}),$ (1)
+Where $\delta_{x_{j}}\in\mathbb{R}_{\geq 0}$ denotes the semantic distance between the original and perturbed prompts.
 3.3 Output-Level Distribution Shift
-To assess how the model’s outputs change when the input is perturbed, we compute the Wasserstein distance between the output distributions for $x$ and $\hat{x}_{j}$ [101]:
-$\Delta(x,\hat{x}_{j}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x}_{j})\right).$ (2)
-Equation 2 captures the magnitude of the change in the model’s behaviour due to input perturbations.
+To assess the effect of input perturbation on the model’s behavior, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs:
+$\Delta(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
+where $y\in\mathcal{Y}$, and $W(p,q)$ denotes the Wasserstein distance between two probability distributions $p$ and $q$.
 3.4 Weighting via Input Similarity
-Since not all perturbations are equally meaningful, we prioritise those that are semantically closer to the original input. We define a Gaussian-based weighting scheme [13]:
-$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right).$ (3)
-As shown in Equation 3, perturbations with smaller semantic distances $\delta_{x_{j}}$ receive higher weights, reflecting their greater relevance.
+Perturbations that are semantically closer to the original input are weighted more heavily. We use a Gaussian-based weighting scheme [63]:
+$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right),$ (3)
+Where $w_{j}\in(0,1]$ denotes the relevance weight of $\hat{x}_{j}$, and $\sigma>0$ is a kernel width parameter that controls the rate of exponential decay.
 3.5 Fitting a Local Surrogate
-Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. The surrogate model $h_{\theta}$ adopts a linear form:
-$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j}.$ (4)
-The parameters $\theta$ are learned by minimising the weighted squared error between the surrogate’s predictions and the observed distribution shifts:
-$\min_{\theta}\sum_{j}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2}.$ (5)
-Equation 5 ensures that the surrogate closely approximates the black-box model’s behaviour within the local neighbourhood around $x$.
+Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. We fit a linear surrogate model of the form:
+$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j},$ (4)
+where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimizing the weighted squared error.
+$\min_{\theta}\sum_{j=1}^{J}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2},$ (5)
+Where the loss ensures local fidelity between the surrogate and the observed output shifts.
 3.6 Theoretical Justification (Lipschitz Smoothness)
 Shift Function Definition.
-We formally define the shift function:
-$f(x,\hat{x}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x})\right).$ (6)
-Assuming that $\pi^{(n)}$ is Lipschitz continuous with respect to its input, the shift function $f$ is also Lipschitz [104]:
-$|f(x,\hat{x})-f(x^{\prime},\hat{x}^{\prime})|\leq L\cdot\left(\|x-x^{\prime}\|+\|\hat{x}-\hat{x}^{\prime}\|\right),$ (7)
-where $L>0$ is a Lipschitz constant (Equation 7). This smoothness justifies the use of a local linear surrogate model to approximate $f$.
+We define a shift function that quantifies the model’s response to perturbation:
+$f(x_{a},\hat{x}_{a})=W\left(\pi^{(n)}(y\mid x_{a}),\pi^{(n)}(y\mid\hat{x}_{a})\right),$ (6)
+If $\pi^{(n)}$ is Lipschitz continuous concerning its input, then the function $f$ is also Lipschitz:
+$\left|f(x_{a},\hat{x}_{a})-f(x_{b},\hat{x}_{b})\right|\leq L\times\left(\|x_{a}-x_{b}\|+\|\hat{x}_{a}-\hat{x}_{b}\|\right),\forall a,b\in\{1,\dots,N_{pert}\}$ (7)
+The notation $x_{a}$ and $\hat{x}_{a}$ represents a specific input and its corresponding perturbation.
+The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analyzing the Lipschitz continuity or stability across perturbations.
+$N_{pert}$ denotes the total number of perturbations.
+[…]
 Relation to Total Variation.
-The Kantorovich-Rubinstein duality [105] relates the Wasserstein distance and Total Variation (TV) distance:
+For two distributions $p$ and $q$ over output space $\mathcal{Y}$, the Kantorovich–Rubinstein duality yields:
 $W(p,q)\leq\text{diam}(\mathcal{Y})\cdot\mathrm{TV}(p,q),$ (8)
-where:
+Where the total variation distance ($TV$) is given by:
 $\mathrm{TV}(p,q)=\frac{1}{2}\sum_{y\in\mathcal{Y}}|p(y)-q(y)|.$ (9)
-Equations 8 and 9 indicate that small changes in the input resulting in small TV distances also imply bounded Wasserstein shifts.
+This bound highlights that small changes in distribution (in the TV sense) imply bounded Wasserstein distance.
 3.7 Linear Surrogates Approximate Lipschitz Functions
-Due to the Lipschitz property, any smooth function $f$ can be locally approximated by a linear model in a small neighbourhood $U$:
-$f(x^{\prime})\approx f(x)+\nabla f(x)\cdot(x^{\prime}-x).$ (10)
-This principle, expressed in Equation 10, underlies the surrogate modelling approach adopted by LIME and related methods [13].
+Due to the differentiability of Lipschitz functions, the shift function $f$ can be locally approximated by a first-order Taylor expansion:
+$f(x^{\prime})\approx f(x)+\nabla f(x)^{\top}(x^{\prime}-x),$ (10)
+where $\nabla f(x)^{\top}(x^{\prime}-x)$ denotes the dot product between the gradient of $f$ and the local displacement in input space.
 3.8 Putting It All Together
-The complete Wasserstein-based LIME surrogate process consists of the following steps:
+The complete Wasserstein-based LIME surrogate procedure consists of:
   1. 1.
-    Generate perturbations $\{\hat{x}_{j}\}$ near $x$.
+    Generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$.
   2. 2.
-    Compute semantic distances $\delta_{x_{j}}$ (Equation 1) and output-level shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
+    Computing semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
   3. 3.
-    Assign weights $w_{j}$ using Equation 3.
+    Assigning similarity weights $w_{j}$ via Equation 3.
   4. 4.
-    Extract feature vectors $z_{j}$.
+    Extracting feature representations $z_{j}$ for each perturbation.
   5. 5.
-    Fit a local surrogate model $h_{\theta}$ by minimising the loss (Equation 5).
+    Fitting a local surrogate model $h_{\theta}$ by minimizing the loss (Equation 5).
   6. 6.
-    Use $h_{\theta}$ to interpret how input perturbations affect the model’s output distribution.
-This framework provides a rigorous, interpretable method for understanding how small changes to an input affect the outputs of complex models, particularly in applications involving large language models [102, 13, 101, 105].
+    Using $h_{\theta}$ to interpret the influence of each input feature on model behavior.
+This framework offers a principled and interpretable approach to analyzing large language models through localized perturbation and response analysis.
 4 PROPOSED METHOD
-The proposed approach enhances the interpretability of large language models for text generation by analysing how specific textual inputs influence the generated outputs.
-Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behaviour [8, 102].
+The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analyzing how specific textual inputs influence the generated outputs.
+Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behavior [9, 114].
 In the context of classification tasks, Fig.
 4 illustrates SMILE, a tool designed to explain model predictions by isolating critical features that drive decision-making [11].
-SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or logical sections for text, and creates perturbations by selectively modifying these components.
+SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or “logical sections” for text, and creates perturbations by selectively modifying these components.
 These perturbations reveal how different input elements contribute to the final prediction.
 Figure 4: SMILE flowchart for explaining image classification [11]
-SMILE records the model’s predictions for each perturbed input and calculates similarity scores between the original and modified inputs using metrics such as the Wasserstein distance [100, 101].
-These scores are transformed into weights using a kernel function, enabling a balanced weighting scheme.
+SMILE records the model’s predictions for each perturbed input.
+It calculates distance scores between the original and modified outputs using metrics such as the Wasserstein distance [112, 113].
+These distances are transformed into similarity-based weights using a kernel function, enabling a balanced weighting scheme.
 The results are used to train a weighted linear regression model that identifies the most influential input components.
@@
 5.
-Instead of perturbing the generated image, as in SMILE, we modify the instruction by including or excluding certain words.
-We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between these outputs and the one produced by the original text prompt [103].
-This distance is a similarity measure, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
+Instead of perturbing the generated image, as in SMILE, we perturb the input prompt by including or excluding certain words.
+We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between each generated output and the original output [115].
+This distance quantifies the difference in the production, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
 These similarity scores are then used as the outcome variable in a weighted linear regression model, with weights derived from the text distances and perturbations [116].
@@
 Using this enhanced three-step process, which builds on SMILE’s methodology [11], we aim to help users understand how particular elements of text prompts impact text generation.
-This technique improves user control and predictability in text generation models by providing insightful information about which elements of textual input, such as stylistic instructions or descriptive terms, most impact the final output.
-The first step employs a text generation model based on input text.
+This technique enhances user control and predictability in text generation models by providing insight into which aspects of textual input, such as stylistic instructions or descriptive terms, have the most significant impact on the final output.
+The first step utilizes a text generation model that is trained on input text.
 The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output.
-The text generation model is expected to produce responses similar to the original response that vary only based on critical components of the input texts [107].
-Different text permutations are utilised from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into individual words.
+The text generation model is expected to produce responses similar to the original response, varying only based on the critical components of the input texts [117].
+Different text permutations are utilized from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into separate words.
 The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [118].
@@
 4.1 Text Generating and Perturbed Prompts
-To begin, we generate variations of the original text prompt, known as “perturbed prompts”, to examine how these subtle changes affect the output of a text generation model [112, 113].
-The original text is first broken down into individual words.
-Multiple text versions are then created by selectively including or excluding specific words.
-Each perturbed prompt is then provided to the text generation model, which produces a unique output for each variation.
-In this process, the output generated from the original prompt is the reference, or baseline output, against which all other outputs are compared.
-For the original text prompt $p_{\text{org}}$, the text generation function $\phi_{\text{gen}}$ generates a corresponding output text $t_{\text{org}}^{\prime}$, as shown in Eq. 11:
+To begin, we generate variations of the original text prompt, referred to as “perturbed prompts,” to analyze how subtle changes in the input affect the output of a text generation model [122, 123].
+The original prompt $p_{\text{org}}$ is first tokenized into individual words.
+From these tokens, multiple perturbed prompts $\{p_{\text{pert}}^{i}\}_{i=1}^{N}$ are constructed by selectively including or excluding specific words.
+Each perturbed prompt is then passed to the text generation model, which produces a corresponding output.
+The output generated from the original prompt, $t_{\text{org}}^{\prime}$, serves as the reference (or baseline) for comparison.
+For the original text prompt $p_{\text{org}}$, the text generation model $\phi_{\text{gen}}$ produces the output:
 $t_{\text{org}}^{\prime}=\phi_{\text{gen}}(p_{\text{org}})$ (11)
-Similarly, for each perturbed prompt $p_{\text{pert}}^{i}$ within the set $S_{\text{pert}}$, the text generation model produces a corresponding output text $t_{\text{pert}}^{\prime i}$, according to Eq.
-12:
-$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in S_{\text{pert}}.$ (12)
-Here, $p_{\text{org}}$ represents the original prompt, $p_{\text{pert}}^{i}$ describes a perturbation generated from the original text, and $\phi_{\text{gen}}$ is the text generation model function used to produce outputs $t^{\prime}$ that reveal the influence of prompt variations.
-The set of all perturbed prompts is denoted by $S_{\text{pert}}$ [13, 108].
+Similarly, for each perturbed prompt $p_{\text{pert}}^{i}\in S_{\text{pert}}$, the model generates:
+$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in\{1,\dots,N\}$ (12)
+Here, $\phi_{\text{gen}}$ is the text generation function, and $t_{\text{pert}}^{\prime i}$ denotes the generated output for the $i$-th perturbation.
+These outputs are used to measure how individual word-level changes in the prompt influence the model’s behavior [63, 118].
 4.2 Creating the Interpretable Space
-In order to map the output space of the text generation model to a one-dimensional space, the Wasserstein distance between each text generated from the perturbed prompts and the original prompt (baseline output) is computed as a measure of similarity [100, 101].
-In all cases, the input prompt remains consistent.
-Word Mover’s Distance (WMD) is employed to extract meaningful embeddings from the texts, allowing the focus to remain on the primary semantic components of the generated text [103].
-This reduces noise in the distance estimation, enabling the accurately calculated distance to serve as the output for the linear regression model.
-The distance between the outputs generated by the perturbation texts and the original prompt is given by Eq. 13:
-$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{p}\right)^{\frac{1}{p}},\quad\forall i\in S_{\text{pert}}.$ (13)
-In the above relation, the $\delta_{\text{WMD}}$ function represents the embedding extraction model applied to the texts before calculating the distance [103].
-In Eq.
-13, $n$ represents the number of features, $p$ denotes the norm order, and $W(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ demonstrates the Wasserstein distance between the outputs produced by the original and perturbed prompts.
+To represent the output space of the text generation model in a one-dimensional similarity space, we compute the Wasserstein distance between each generated text from a perturbed prompt and the output generated from the original prompt [112, 113].
+Throughout this process, the input structure remains unchanged, and only the perturbed words differ across inputs.
+To capture the semantic similarity between generated outputs, we use Word Mover’s Distance (WMD) to embed the texts into a feature space [115].
+These embeddings enable the focus to remain on high-level semantic content, thereby mitigating lexical noise and enhancing the reliability of the computed distances.
+The resulting scalar distances are used as the dependent variable in the subsequent regression model.
+The Wasserstein distance between each perturbed output $t_{\text{pert}}^{\prime i}$ and the original output $t_{\text{org}}^{\prime}$ is computed as:
+$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{P}\right)^{\frac{1}{P}},\quad\forall i\in\{1,\dots,N\}$ (13)
+In this equation:
+  • $\delta_{\text{WMD}}(\cdot)$ denotes the semantic embedding of a generated text,
+  • $n$ is the dimensionality of the embedding space,
+  • $P$ specifies the norm order used in the Wasserstein metric,
+  • and $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance computed for the $i$-th perturbation.
 Using p-values is crucial for validating the significance of the observed Wasserstein distances (WD) between texts generated under different prompt perturbations.
@@
 If a small proportion of sampled WDs exceed the observed WD, the resulting p-value confirms that the differences are statistically significant, supporting the validity of the perturbation effects.
-The Wasserstein-$p$ metric allows us to adjust WD’s sensitivity to different text variations by fine-tuning the norm order $p$.
-Smaller $p$ values emphasise local differences, while larger $p$ values capture broader distributional shifts [101].
-By testing various $p$-values, we identify the norm that best captures meaningful variations caused by prompt changes, enhancing the reliability and interpretability of the results [101].
-When analysing Wasserstein distances, measures with high p-values, indicating non-significant differences, are excluded from the procedure [114].
-As described in Algorithm 1, the Bootstrap algorithm calculates these p-values for all relevant ECDF-based distance measures, including WD.
-The algorithm computes the WD and its associated p-value.
-For a univariate example, let X and Y be the inputs.
+The Wasserstein-$P$ metric enables adjustment of the sensitivity of the Wasserstein Distance (WD) to different types of text variations by tuning the norm order $P$.
+Smaller values of $P$ emphasize fine-grained, local differences between distributions, while larger values capture broader, more global shifts [113].
+By experimenting with various values of $P$, we can select the one that best captures the semantic differences introduced by prompt perturbations, thereby improving both the reliability and interpretability of the analysis [113].
+To assess the statistical significance of the computed Wasserstein distances, we apply hypothesis testing using p-values.
+Distances associated with high p-values, indicating non-significant variation, are excluded from further consideration [124].
+As detailed in Algorithm 1, the Bootstrap algorithm [127, 108] is used to estimate these p-values for each observed WD.
+For example, let X and Y denote two sets of text-based embeddings.
 The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y.
 In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed.
-If boostWD exceeds the observed WD, a counter (bigger) is incremented.
-After completing all iterations, the p-value (pVal) is calculated as the proportion of iterations where boostWD is greater than the observed WD, as shown in Algorithm 1.
+If boostWD exceeds the original observed WD, a counter (bigger) is incremented.
+After all iterations, the p-value (pVal) is estimated as the proportion of iterations where boostWD exceeds the observed WD, as outlined in Algorithm 1.
 Input: Two sets of text-based distances: $X$ and $Y$
@@
 Algorithm 1 Bootstrap Algorithm for Wasserstein Distance (WD) P-Value Calculation
-Figure 5: SMILE flowchart for Large Language Models
-When generating perturbation texts derived from the original text, the keywords of the original text play a crucial role in maintaining certain levels of similarity and fidelity between the perturbation texts and the source content.
-This is important when the interpretable model is trained on perturbed texts as input.
-Therefore, a similarity criterion is used to measure the degree of similarity between each perturbation text and the original, and this value is employed as a sample weight in the interpretable model.
-Eq.
-15 calculates this similarity, and a Gaussian kernel is applied to normalise it into the range $[0,1]$ [103].
-$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\,d(w_{k},w_{l})$ (14)
-subject to $\sum_{l=1}^{m}T_{kl}=p_{k},\quad\forall k\in S_{\text{org}}$
-$\sum_{k=1}^{n}T_{kl}=q_{l},\quad\forall l\in S_{\text{pert}}$
+Figure 5: gSMILE flowchart for Large Language Models
+When generating perturbation texts from the original prompt, the key terms in the prompt are essential for maintaining semantic similarity between perturbed prompts and the source.
+To account for this, we define a similarity metric between each perturbed prompt and the original, which is then used as a sample weight in the interpretable model.
+This similarity is quantified using Word Mover’s Distance (WMD), which calculates the minimum cost to transport the embedded words of one document to match another [115].
+The WMD between the original prompt $p_{\text{org}}$ and a perturbed prompt $p_{\text{pert}}^{i}$ is defined as:
+$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\cdot d(w_{k},w_{l})$ (14)
+subject to $\sum_{l=1}^{m}T_{kl}=d_{k},\quad\forall k=1,\dots,n$
+$\sum_{k=1}^{n}T_{kl}=d^{\prime}_{l},\quad\forall l=1,\dots,m$
 $T_{kl}\geq 0$
-$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\exp\left(-\left(\frac{C_{i}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma^{2}}\right)^{2}\right),\quad\forall i\in S_{\text{pert}}$ (15)
-The function $\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ quantifies the semantic distance between the perturbation text and the original.
-The resulting weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is obtained by applying a Gaussian kernel to this distance [106], allowing the surrogate model to assign higher importance to more semantically similar inputs.
+Here:
+  • $w_{k}$ and $w_{l}$ denote the $k$-th and $l$-th words from the original and perturbed prompts respectively,
+  • $d(w_{k},w_{l})$ is the distance (typically cosine or Euclidean) between the embeddings of $w_{k}$ and $w_{l}$,
+  • $T_{kl}$ is the flow variable indicating how much of $w_{k}$’s weight is transported to $w_{l}$,
+  • $d_{k}$ and $d^{\prime}_{l}$ represent the normalized bag-of-words weights for the original and perturbed texts.
+To use this distance as a weighting factor in the regression model, we normalize it using a Gaussian kernel:
+$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})=\exp\left(-\left(\frac{\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma}\right)^{2}\right),\quad\forall i\in\{1,\dots,N\}$ (15)
+Here, $\sigma$ is a kernel width parameter that controls sensitivity to dissimilar inputs.
+The resulting similarity score $\pi_{i}$ falls in the range $[0,1]$, assigning higher weights to perturbed prompts that are semantically closer to the original.
 4.3 Developing the Interpretable Surrogate Model
-We are innovative in mapping the output space of the text generation model to a one-dimensional distance scale, accompanied by corresponding perturbation texts for the calculated distances [117].
-This method enables us to train an interpretable model on the mapped output space of the text generation model, effectively demonstrating how input words influence text generation.
-In our linear regression model, the vectors of the perturbation texts are critical, as they are treated as independent variables.
-Following the extraction of text embeddings, the distance between the outputs of these perturbation texts and the corresponding original text output is regarded as the response variable.
-Additionally [118], the degree of similarity between each perturbation text and the original text is assumed to represent the weight of the samples [102].
-Eq.
-16 outlines the square loss objective function for interpretable weighted linear regression:
-$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=$ (16)
-$\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\times\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$
-Here, $\phi_{\text{gen}}$ represents the text generation model, and $\delta_{\text{WMD}}$ is the Word Mover’s Distance function applied to compute similarity between generated outputs.
-The weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is derived using a Gaussian kernel function and reflects the degree of similarity between the original prompt and each perturbed prompt.
-The term $\hat{f}_{t}(p_{\text{pert}}^{i})$ is a predicted value from the linear regression model.
-After feature extraction, the $\hat{f}_{i}(p_{\text{pert}}^{i,01})$ function is a linear regression model trained over the one-hot vector space of perturbation texts $(p_{\text{pert}}^{i,01})$ and the calculated distances [102].
+To interpret the influence of individual words on text generation, we map the output space of the generation model to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [128].
+This enables the training of a local surrogate model that relates changes in input structure to variations in output behavior.
+In this weighted linear regression framework, the perturbation vectors (i.e., binary or one-hot encodings of included or excluded tokens) serve as independent variables.
+The response variable is the Wasserstein distance between the output of a perturbed prompt and that of the original prompt.
+Each perturbation is assigned a weight based on its semantic similarity to the original prompt [114, 129].
+The loss function minimized during training is a weighted squared error:
+$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\cdot\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$ (16)
+Here:
+  • $\phi_{\text{gen}}$ is the text generation model,
+  • $\delta_{\text{WMD}}$ denotes the function used to embed generated texts,
+  • $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance between the generated outputs,
+  • $\pi_{i}$ is the similarity-based kernel weight for the $i$-th perturbation,
+  • and $\hat{f}_{t}(p_{\text{pert}}^{i})$ is the prediction from the linear surrogate model.
+The surrogate model $\hat{f}_{t}$ is trained on the one-hot vector representation of the perturbation prompts $p_{\text{pert}}^{i,01}$, using the computed Wasserstein distances as targets [114].
 4.4 Evaluation Metrics
@@
 In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations.
-Stability is a key property for fostering trust in explainable artificial intelligence (XAI) systems [14].
+Stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61].
 It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [10].
@@
 4.4.4 Fidelity
-Fidelity is a metric to evaluate how well an explainable model aligns with a black-box model.
-The predictions of the black-box model, denoted as $f(x_{i})$, and the explainable model, $g(x_{i})$, for the $i$-th data perturbation are compared.
-Various loss functions and coefficient metrics quantify the similarity between the models’ predictions, measuring how effectively the explainable model approximates the black-box model.
-The coefficient of determination $R^{2}$ is an accuracy-based metric that measures how well the explainable model correlates with the black-box model [124], and is defined in Eq. 18:
-$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}(X_{i}))^{2}}$ (18)
-where $\bar{f}(X_{i})$ represents the average of $f(X_{i})$. Values of $R^{2}$ close to 1 indicate a strong alignment between the two models.
-For scenarios where weighting is necessary, the weighted coefficient of determination $R_{w}^{2}$ provides an alternative version [16, 125], defined as Eq. 19:
-$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f_{w}}(X_{i}))^{2}}$ (19)
-where $\bar{f_{w}}(X_{i})$ represents the weighted average of $f(X_{i})$.
-These accuracy-based metrics, $R^{2}$ and $R_{w}^{2}$, offer insight into how well the explainable model approximates the black-box model.
-To account for both sample size and the number of features used in the local explanation, the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ [16] is used, as shown in Eq. 20:
+Fidelity is a metric used to evaluate how closely an explainable model replicates the behavior of a black-box model.
+Specifically, the predictions of the black-box model denoted as $f(X_{i})$, are compared to those of the explainable model, $g(X_{i})$, for each input perturbation $i$.
+A variety of accuracy-based and error-based metrics are employed to assess this alignment, thereby quantifying how well the explainable model approximates the responses of the black-box model.
+Let $N_{p}$ denote the total number of perturbation samples (i.e., data points).
+One of the most commonly used accuracy metrics is the coefficient of determination $R^{2}$, which measures how well the explainable model’s predictions correlate with those of the black-box model [135].
+It is defined as:
+$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f})^{2}}$ (18)
+where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{i})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
+In cases where data samples have different levels of importance, a weighted variant is used. The weighted coefficient of determination $R_{w}^{2}$ is given by [136, 137]:
+$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}_{w})^{2}}$ (19)
+Here, $\bar{f}_{w}$ is the weighted mean of the black-box predictions.
+These accuracy-based metrics ($R^{2}$ and $R_{w}^{2}$) provide insight into how closely the explainable model aligns with the black-box model.
+To account for both sample size and model complexity (i.e., number of features used), the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ is used [136]:
 $\hat{R}_{w}^{2}=1-(1-R_{w}^{2})\left[\frac{N_{p}-1}{N_{p}-N_{s}-1}\right]$ (20)
-To measure the error or risk between the models’ predictions, error-based metrics, including the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE), are used [126].
-These metrics quantify the weighted differences between the predictions of the explainable model and those of the black-box model.
-Lower values indicate better alignment.
-WMSE and WMAE are defined in Eqs.
-21 and 22:
+Where $N_{s}$ is the number of features in the interpretable model.
+In addition to correlation-based metrics, error-based metrics are widely used to quantify differences in prediction values.
+The classical unweighted forms Mean Squared Error (MSE) and Mean Absolute Error (MAE) are:
+$\text{MSE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (21)
+$\text{MAE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (22)
+For situations requiring sample-level importance, the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE) are defined as [138]:
 $\text{WMSE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}w_{i}}$ (23)
 $\text{WMAE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot|f(X_{i})-g(X_{i})|}{\sum_{i=1}^{N_{p}}w_{i}}$ (24)
-Here, $w_{i}$ denotes the weight for each data point $i$, $f(X_{i})$ represents the predictions from the black-box model, $g(X_{i})$ represents the predictions from the explainable model, and $N_{p}$ is the total number of data points.
-Additionally, the mean $L_{1}$ and mean $L_{2}$ losses, defined below, are commonly used to assess prediction discrepancies [127, 16], as shown in Eqs. 23 and 24:
+Here, $w_{i}$ denotes the sample weight associated with the $i$-th perturbation.
+Moreover, the mean $L_{1}$ and mean $L_{2}$ norms are also applied to evaluate absolute and squared differences respectively [139, 136]:
 $L_{1}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (25)
 $L_{2}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (26)
-These error-based metrics, including MSE, MAE, $L_{1}$, and $L_{2}$, quantify the differences between the black-box and explainable models’ predicted scores, reflecting the risk of prediction discrepancies.
-We compute fidelity across various scenarios.
-Fidelity is measured by comparing the predictions of the explainable model to those of the black-box text generation model.
-We analyse how different perturbations in the input text affect fidelity scores and how various distance metrics for text-to-text comparisons perform when using linear regression and Bayesian ridge as surrogate models.
+These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$—together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
+We compute fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model.
+This includes analyzing how different input perturbations affect fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behavior.
 5 Experimental Results
@@
 We analyse how the structure of textual prompts affects interpretability, how variations in input text influence the clarity of explanations, and how the proposed solution performs across diverse scenarios.
-Experiments were conducted using a range of datasets and four state-of-the-art text generation models, OpenAI GPT [1], LLaMA [5], and Claude-AI [128] to ensure a comprehensive assessment.
+Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI GPT [1], LLaMA [19], and Claude-AI [140] to ensure a comprehensive assessment.
 5.1 Qualitative Results
@@
 Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response.
-Figure 8 presents an example where the SMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram.
+Figure 8 presents an example where the gSMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram.
 This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
 Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models.
-As shown in Figure 8, SMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [129].
-Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the SMILE framework.
-The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates SMILE’s token-level importance heatmap for the same prompt.
+As shown in Figure 8, gSMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [141].
+Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the gSMILE framework.
+The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates gSMILE’s token-level importance heatmap for the same prompt.
 [141]
-5.1.3 Word-Level Attribution Analysis on MMLU Using SMILE
-To evaluate the interpretability of ChatGPT using the SMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [130].
+5.1.3 Word-Level Attribution Analysis on MMLU Using gSMILE
+To evaluate the interpretability of ChatGPT using the gSMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [142].
 This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities.
@@
 Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter.
-These visualisations demonstrate SMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
-Figure 9: Attribution heatmaps generated by SMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
+These visualisations demonstrate gSMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
+Figure 9: Attribution heatmaps generated by gSMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
+5.1.4 Optimised Instructions from Google Research
+To further explore how instruction phrasing influences interpretability, we applied gSMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [144].
+These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
+We first applied gSMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem.
+” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
+We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, gSMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
+Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using gSMILE with GPT-3.5.
+Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using gSMILE with GPT-4.
 5.2 Quantitative Results
-In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to asses our explainability method across different text generation models and different scenarios.
+In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to assess our explainability method across different text generation models and different scenarios.
 5.2.1 Accuracy
@@
 The results were averaged across the ten different prompts, and each model’s performance is summarised in Table 1, showcasing their effectiveness based on these metrics.
+Among the evaluated models, Claude-AI demonstrated the strongest performance, achieving the highest scores in all three metrics: 0.82 in ATT ACC, 0.67 in ATT F1, and 0.88 in ATT AUROC.
+These results indicate Claude-AI’s superior ability to consistently focus on essential input tokens across varying prompts.
+In contrast, LLAMA achieved the second-highest accuracy but showed a notably lower F1-score (0.40), suggesting potential inconsistency in identifying key tokens.
+OpenAI GPT yielded balanced results but lagged slightly behind Claude-AI in all metrics, highlighting room for improvement in its attention fidelity.
+Ll, Claude-AI provides more reliable and semantically focused attention compared to its counterparts.
 Table 1: Performance metrics (ATT ACC, ATT F1, and ATT AUROC) for various
@@
 For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word.
-To evaluate stability, we added the token ‘‘ at the end of each prompt as input text, extracted the coefficients again, and computed the Jaccard index to compare the sets of coefficients before and after the perturbation.
-Finally, we calculated the average Jaccard index for each model, and the results are presented in Table 2.
-Table 2: Stability across different models for ten prompts and 30 perturbations
+To evaluate stability, we added only the token  at the end of the text, extracted the coefficients again, and computed the Jaccard index between the coefficients before and after the perturbation.
+The averaged Jaccard indices across all ten prompts for each model are reported in Table 2.
+Higher scores indicate greater stability under minor input changes.
+Table 2: Stability scores across models based on average Jaccard index over ten prompts
 Model name Jaccard Index
@@
 This step is essential to confirm that the model’s predictions are not random or heavily dependent on initialisation factors but are instead deterministic and robust.
+Among all models, OpenAI-GPT exhibited the highest consistency, achieving a variance of 0.0000 and a standard deviation of 0.0046.
+This indicates that its predictions are the most stable across repeated runs, confirming its robustness and reliability in generating interpretable outputs.
 Table 3: Consistency metrics for different models on the prompt
@@
 Claude-AI 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
+From the results, we observe that Claude-AI achieves the lowest WMSE, WMAE, and mean-L1, indicating superior performance in minimizing error.
+On the other hand, OpenAI-GPT performs best in terms of $R^{2}_{\omega}$, mean-L2, and adjusted $R^{2}_{\hat{\omega}}$, demonstrating stronger alignment and generalization of the surrogate model.
+LLaMA ranks in the middle across most metrics but does not lead in any.
+This suggests that different models excel in different fidelity dimensions, highlighting the importance of multi-metric evaluation when interpreting black-box model behavior.
 5.2.5 Fidelity Across Different number of Text Perturbations
@@
 256 0.0191 0.2670 0.0902 0.0873 0.0163 0.2494
+For OpenAI GPT and LLaMA, increasing the number of perturbations generally improves fidelity across most metrics, with 256 perturbations achieving the best WMSE, WMAE, and loss values.
+Interestingly, for Claude-AI, 64 perturbations yield the lowest WMSE and WMAE, while 256 achieves the best loss (mean-L2 and mean-L1).
+However, the highest $R^{2}$ scores for all three models occur at 32 perturbations, suggesting that beyond a certain point, more perturbations may dilute interpretability despite improving local fit.
+Therefore, there is a trade-off between the number of perturbations and the reliability of the surrogate model, depending on the metric of interest.
 5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
-[…]
+Fidelity is also computed by comparing various distance metrics for text generation models such as OpenAI GPT using text-to-text comparisons.
+The surrogate models used are Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME), as shown in Table 8.
+We explore different combinations of distance metrics, including Cosine similarity and Wasserstein Distance (WD).
+Fidelity is measured using WMSE, WMAE, mean-$L_{1}$, mean-$L_{2}$, and coefficient of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$).
 Table 8: Fidelity results for different distance measures with 30 perturbations
@@
 WD+C WD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
+Our Proposed Method: gSMILE.
+Based on the results in Table 8, the combination where Wasserstein Distance (WD) is used for both input similarity and output shift consistently yields the highest values of the weighted coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$).
+Specifically, underweighted linear regression, the WD–WD configuration achieves $R^{2}_{\omega}=\textbf{0.7104}$ and $R^{2}_{\hat{\omega}}=\textbf{0.6409}$, which significantly outperform other configurations.
+Motivated by these findings, we introduce our method, called gSMILE (generalized SMILE), which utilizes the Wasserstein distance for both input and output comparisons.
+This choice enhances the surrogate model’s fidelity and stability in capturing local shifts in model behavior.
+We use gSMILE throughout subsequent sections as the foundation for explainability evaluations on large language models (LLMs).
+These include assessments of fidelity, consistency, and stability, ensuring the proposed approach is both theoretically grounded and empirically adequate.
 5.2.7 Computation Complexity
-We apply our model to three instruction-based text generation frameworks: OpenAI GPT, LLaMA, and Claude-AI.
-We also evaluate the execution time of different explainability methods: LIME, SMILE, and Bay-LIME.
-Each method is tested with 60 perturbations on the same hardware to ensure consistency in computational conditions.
-The models exhibit notable differences in execution time, primarily due to how each explainability method computes similarity and assigns weights.
-SMILE, in particular, employs the Wasserstein distance to quantify text dissimilarities, making it substantially more computationally expensive than both LIME and Bay-LIME.
-As reported in Table 9, SMILE consistently incurs higher execution times across all three model frameworks.
-This pattern also holds across other frameworks, indicating that SMILE’s computational overhead is a systemic characteristic of its underlying methodology.
-Wasserstein distance is known for capturing subtle differences between text embeddings, making it especially effective in high-dimensional semantic spaces [101].
-However, this precision comes at a cost.
-Computing the Wasserstein-2 distance involves operations on covariance matrices with time complexity proportional to $O(d^{3})$, where $d$ is the embedding dimensionality.
-When applied across $N$ samples, the total complexity becomes $O(Nd^{3})$.
-In contrast, LIME relies on cosine similarity, which has a much lower computational cost of $O(Nd)$.
-This disparity arises because Wasserstein-2 distance requires solving matrix optimisation problems over $d$-dimensional Gaussian distributions [132].
-While this added complexity leads to more stable and expressive explanations, as demonstrated by SMILE, it significantly increases runtime.
-Conversely, LIME avoids this overhead by utilising simpler vector-based similarity measures.
-In summary, SMILE offers greater fidelity and stability but at a higher computational cost.
-Its use of the Wasserstein distance makes it ideal for tasks where precision and interpretability are essential, even if that entails longer execution times.
-LIME and Bay-LIME, while faster, may trade off some robustness in favour of efficiency.
+We evaluate the computational efficiency of our proposed method, gSMILE, using three instruction-based text generation models: OpenAI GPT, LLaMA, and Claude-AI.
+Although these models differ in scale and architecture, the objective here is not to compare the models themselves but rather to assess the performance of different explainability methods applied to each model under identical hardware and runtime conditions.
+All evaluations are conducted with 60 perturbations per input prompt.
+We compare three explainability approaches—LIME, gSMILE, and Bay-LIME—based on their execution time.
+As reported in Table 9, gSMILE consistently requires the highest runtime across all model frameworks.
+This is mainly due to its reliance on the Wasserstein distance to compute similarity between output distributions, which involves solving high-dimensional optimization problems with time complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [145].
+By contrast, LIME and Bay-LIME utilize cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
+Despite the higher computational cost, gSMILE delivers significantly better fidelity, stability, and robustness to adversarial perturbations—essential properties for trustworthy explanations.
+These advantages are supported by results in Tables 8 and 2 and align with findings from prior work on SMILE [11], which demonstrated superior robustness against adversarial input variations.
+These findings suggest that gSMILE’s computational overhead is justified by the quality of explanations it provides, making it highly suitable for applications where interpretability, consistency, and resistance to manipulation are paramount.
 Table 9: Execution times (in seconds) for each framework and explainability method with 60 perturbations
-Framework LIME SMILE Bay-LIME
+Framework LIME gSMILE Bay-LIME
 OpenAI-GPT 156.82 170.70 161.43
@@
 Claude-AI 355.09 372.95 349.88
-5.2.8 Optimised Instructions from Google Research
-To further explore how instruction phrasing influences interpretability, we applied SMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [133].
-These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
-We first applied SMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem.
-” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
-We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, SMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
-Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using SMILE with GPT-3.5.
-Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using SMILE with GPT-4.
 5.2.8 Token-Level Attention for Different Sentence Structures
@@
 Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures.
-Hovering over each point reveals the full sentence associated with that score.
+Each point represents the attention score assigned to the token in a specific sentence instance.
 6 Limitations
-While SMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
-1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, even when the semantic meaning remains largely intact.
+While gSMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
+1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, particularly in English, even when the semantic meaning remains largely intact.
 This sensitivity challenges the assumption of local smoothness required for regression-based attribution and can reduce explanation stability.
@@
 This work introduced a novel interpretability framework tailored for instruction-based text generation models.
-Leveraging SMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output.
+Leveraging gSMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output.
 Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
-Whereas most existing interpretability methods highlight important tokens in the model’s output [13, 134], SMILE shifts attention to the input space.
+Whereas most existing interpretability methods highlight important tokens in the model’s output [63, 64], gSMILE shifts attention to the input space.
 By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
-Our experiments show that SMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma.
-The resulting heatmaps and attribution scores intuitively understand model behaviour, empowering users to control better and refine prompt design.
+Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma.
+The resulting heatmaps and attribution scores provide an intuitive understanding of model behavior, empowering users to control and refine prompt design better.
+gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.
 This work lays the groundwork for more transparent and trustworthy natural language systems.
-Future directions include extending SMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
+Future directions include extending gSMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
 8 Future Works
@@
 These benchmarks will support consistent assessment across models and help identify best practices for interpretable generation.
+  6. 6.
+    Optimisation of Computational Efficiency: To address the computational demands of gSMILE, future work will investigate dimensionality reduction techniques and approximation strategies.
+These approaches aim to reduce execution time while preserving the integrity and interpretability of the model’s outputs.
 Data Availability
 The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub.
-For the SMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy.
+For the gSMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy.
 For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.

### v4

@@
-Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations
+Explaining Large Language Models with gSMILE
 Abstract
-Large Language Models (LLMs) such as GPT, LLaMA, and Claude have achieved remarkable performance in text generation, yet their decision-making processes remain opaque.
-This lack of transparency poses challenges in domains where trust and accountability are critical.
-In this work, we introduce gSMILE (Generative SMILE), a generalized and LLM-oriented extension of the previously proposed SMILE framework, for interpreting how these models respond to specific parts of an input prompt.
-gSMILE is model-agnostic and operates by introducing controlled perturbations to the input, observing the resulting output changes, and identifying which input words have the most significant impact on the output.
-This process enables the generation of intuitive heatmaps that visually highlight influential words in the prompt.
-We evaluate gSMILE across several leading large language models (LLMs) using metrics such as accuracy, consistency, stability, and fidelity, demonstrating its ability to provide precise, robust, and reliable explanations.
-By making LLMs more interpretable, gSMILE takes an important step toward building more transparent and trustworthy AI systems.
-Figure 1: Highlighting the contribution of each word in the input prompt of a Commercial LLM
+Large Language Models (LLMs) such as GPT, LLaMA, and Claude achieve remarkable performance in text generation but remain opaque in their decision-making processes, limiting trust and accountability in high-stakes applications.
+We present gSMILE (generative SMILE), a model-agnostic, perturbation-based framework for token-level interpretability in LLMs.
+Extending the SMILE methodology, gSMILE uses controlled prompt perturbations, Wasserstein distance metrics, and weighted linear surrogates to identify input tokens with the most significant impact on the output.
+This process enables the generation of intuitive heatmaps that visually highlight influential tokens and reasoning paths.
+We evaluate gSMILE across leading LLMs (OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo, and Anthropic’s Claude 2.1) using attribution fidelity, attribution consistency, attribution stability, attribution faithfulness, and attribution accuracy as metrics.
+Results show that gSMILE delivers reliable human-aligned attributions, with Claude 2.1 excelling in attention fidelity and GPT-3.5 achieving the highest output consistency.
+These findings demonstrate gSMILE’s ability to balance model performance and interpretability, enabling more transparent and trustworthy AI systems.
+Figure 1: Overview of gSMILE interpretability process: An input prompt is processed by a transformer-based LLM, and gSMILE generates a token-level heatmap highlighting influential words contributing to the output.
 1 Introduction
-Large Language Models (LLMs), such as GPT, Claude, and LLaMA, have significantly advanced the capabilities of natural language processing systems [1, 2, 3].
-These models, based on the Transformer architecture [4], are trained on extensive corpora and exhibit remarkable performance across a wide range of tasks, including text generation, summarisation, code completion, and dialogue reasoning [5].
-However, despite their widespread adoption, LLMs remain largely opaque in how they make predictions, leading to concerns about transparency, trust, and safety, especially in high-stakes domains such as healthcare and law [6, 7].
-The black-box nature of LLMs limits our ability to understand which parts of an input prompt have the most significant influence on the model’s output.
-This lack of interpretability hinders debugging, ethical assessment, and responsible deployment [8, 9].
-Interpretability methods, especially those that operate post hoc and in a model-agnostic fashion, offer a way to open these black boxes by explaining the model’s behavior without requiring access to internal parameters [10].
-In this work, we propose gSMILE, a generalized and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [11].
+Large Language Models (LLMs) such as GPT-4 [1], Claude [2], and LLaMA [3] have rapidly become central to artificial intelligence, powering applications in healthcare, education, law, and scientific research [4, 5, 6, 7].
+Built on the Transformer architecture [8] and trained on massive corpora, they generate fluent, context-aware responses for tasks ranging from dialogue and summarisation to reasoning and information extraction [6].
+Today, LLMs are embedded in widely used technologies, assisting clinicians in healthcare [4], supporting personalised education [5], streamlining customer service [6], and aiding legal research [7].
+Their growing influence underscores both their transformative potential and the urgent need to address concerns about reliability, transparency, and trust.
+Despite their remarkable capabilities, a fundamental question remains: how much control do we truly have over LLM behaviour, and to what extent do we understand their decision-making processes?
+While they can generate compelling outputs, studies show that these responses are not always reliable, often containing hallucinations or logical inconsistencies that damage trust [4, 5].
+In practice, users are left uncertain about why a model provides one answer instead of another, raising concerns about whether we genuinely comprehend the mechanisms driving their outputs [9].
+Recent work has also highlighted the critical role of prompt design in shaping LLM performance.
+Simple changes to instructions, for example, adding phrases such as “Let’s think step by step” or “Take a deep breath and work on this problem step-by-step, have been shown to improve accuracy on reasoning benchmarks significantly.
+OPRO reports that the latter phrase boosts GSM8K accuracy to 80.2%, compared to 71.8% with the baseline step-by-step prompt [10].
+Similarly, Self-Supervised Prompt Optimisation (SPO) demonstrates that optimised prompts, even without external references, consistently improve results across multiple datasets [11].
+These findings underline the power of prompt engineering: if we understood more clearly which features of a prompt drive performance, we could design instructions that yield more reliable outputs.
+Interpretability methods can provide the insights needed to guide such prompt design.
+Moreover, interpretability approaches must be model-agnostic.
+While Anthropic has recently showcased advanced interpretability research for Claude 3.5, using techniques such as circuit tracing and attribution graphs to visualise how the model plans, manipulates concepts, and sometimes even fabricates reasoning [12], such access to internal mechanisms is limited to others with privileged model access.
+For most practitioners and end users, leading LLMs like GPT and Claude are available only through APIs that restrict visibility into their parameters and architectures.
+As a result, practical interpretability techniques must be designed to explain model behaviour without relying on direct access to internal weights or structural details [6].
+To this end, we propose gSMILE, a generalised and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [13].
 gSMILE interprets model outputs by perturbing input prompts, measuring output shifts using Wasserstein distance, and fitting a local linear surrogate model.
-The result is a weighted heatmap that visualizes how individual words in a prompt contribute to the model’s response.
-We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including fidelity, consistency, stability, and accuracy.
+The result is a weighted heatmap that visualises how individual words in a prompt contribute to the model’s response.
+We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including ATT fidelity, ATT consistency, ATT stability, ATT faithfulness and ATT accuracy.
 Our findings demonstrate that gSMILE yields stable, interpretable, and reliable attributions, rendering it suitable for practical applications in real-world scenarios involving instruction-based generation.
@@
   • Token-level interpretability for LLMs: We present gSMILE, a model-agnostic method that quantifies the influence of each input token on the output of large language models using perturbation-based analysis and Wasserstein distance.
-  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behavior.
-  • Robust cross-model evaluation: We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, accuracy, fidelity, stability, and consistency that demonstrate its reliability and adaptability.
+  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behaviour.
+  • Robust cross-model evaluation:We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, including ATT accuracy, ATT fidelity, ATT stability, ATT faithfulness and ATT consistency, that demonstrate its reliability and adaptability.
+List of Abbreviations
+Abbreviation Definition
+LLM Large Language Model
+XAI Explainable Artificial Intelligence
+SMILE Statistical Model-agnostic Interpretability with Local Explanations
+gSMILE Generative SMILE (proposed method)
+WMD Word Mover’s Distance
+IWMD Inner Word Mover’s Distance (input-level distance)
+OWMD Outcome Word Mover’s Distance (output-level distance)
+WD Wasserstein Distance
+ATT Attribution
+TV Total Variation distance
+P Norm order in Wasserstein-$P$ metric
+$\pi^{(n)}$ Black-box text generation model
+$\Delta(x,\hat{x}_{j})$ Output-level semantic shift
+$w_{j}$ Similarity-based kernel weight
+$z_{j}$ Feature vector representation of perturbed prompt
 2 Literature review
@@
 2.1 Large Language Models
-Large language models (LLMs) have become increasingly prevalent due to significant advancements in deep learning [12], improved access to large-scale datasets, and the growing availability of powerful computing resources [13].
-These models are typically trained on massive corpora of text.
-They are capable of generating fluent, contextually appropriate responses across a wide variety of language-based tasks, including question answering, summarisation, translation, and content generation.
-Over time, several influential large language models (LLMs) have helped push the field forward, each introducing novel architectural innovations [4], training strategies [14], or alignment techniques [15] that have contributed to their growing capabilities and broader adoption.
-Among the most influential models, GPT (OpenAI) [1] stands out as a leading autoregressive transformer-based architecture capable of generating coherent and contextually rich text.
-It set a new standard for generative language models and introduced the broader public to powerful applications of AI through tools like ChatGPT.
-Expanding on this paradigm, Google’s T5 [14] reframed all NLP tasks as text-to-text problems, streamlining the transfer of learning across diverse tasks and contributing to more unified and efficient architectures.
-Palm (Google) [16], with its Pathways design, extended the multitask learning capability at scale, allowing a single model to generalize effectively across different NLP applications.
-In parallel, Anthropic’s Claude [17] introduced a new approach centered on safe and aligned AI behavior.
-Claude utilizes constitutional AI principles to ensure its outputs are consistent with predefined ethical standards, thereby promoting greater transparency and reducing the risk of harm.
-More recently, Google’s Gemma 2 [18] has focused on logical reasoning and structured data interpretation, offering a competitive option for research-driven and enterprise-level deployment.
-Meta’s LLaMA [19] series continues to draw attention as a scalable, open-source alternative, providing the research community with accessible tools that deliver high performance across a range of benchmarks.
-Community-driven efforts have also enriched the landscape.
-EleutherAI’s GPT-NeoX [20] was developed to provide an open-source alternative to proprietary GPT models, offering researchers greater flexibility in experimentation and deployment.
-AI21 Labs’ Jurassic-2 [21] targets commercial-scale applications with a focus on contextual awareness and enterprise integration.
-Meanwhile, Cohere’s Command R+ [22] integrates retrieval-augmented generation (RAG), enhancing factual consistency and minimizing hallucinations in the generative output — a key concern for the practical adoption of large language models (LLMs).
-Additionally, transformer-based models, such as BERT [23], have revolutionized NLP through bidirectional encoding, providing substantial improvements in tasks like question answering, named entity recognition, and sentiment analysis.
-Efficiency-focused models, such as Microsoft’s Phi-4, highlight a growing trend toward parameter-efficient design without sacrificing core performance, making large language models (LLMs) more accessible for low-resource environments.
-Mistral and Qwen 2.5 (by Alibaba) continue this direction by offering fine-tuning-friendly models with strong multilingual and domain adaptation capabilities.
-Several recent open-source contenders further enrich the ecosystem.
-DeepSeek-LLM [24] delivers performance comparable to proprietary models while remaining open and extensible.
-Mixtral and StableLM focus on task efficiency and targeted performance, ensuring their utility in production environments.
-DeepSeek-R1 has been noted for its emphasis on reasoning, reflecting a broader shift from general language generation to models optimized for critical thinking, chain-of-thought prompting, and interpretability.
-Cutting-edge advancements have introduced even more specialized architectures.
-Qwen2.5-Max utilizes large-scale Mixture-of-Experts (MoE) models to enhance throughput and scalability.
-OpenAI’s o3-mini emphasizes low-latency, cost-effective deployment while retaining high-quality output — ideal for mobile and edge computing.
-Perhaps most notably, DeepSeek-V3 represents a significant milestone by achieving GPT-4o-level performance in an open-source setting, suggesting that open-source models may soon rival or surpass closed models in both capability and adoption.
-The following models in Fig.
-2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications.
-These models continue to define the state of the art in machine-generated language, with a growing focus on transparency, efficiency, and alignment with human values.
-Category Models Description General-Purpose GPT-4 [25], GPT-4o [26], Claude 3 [27], Gemini 1.5 [28], Command R+ [29], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [32], Falcon 180B [33], DeepSeek-V3 [34], OpenChat [35] Customizable models for research applications Retrieval-Augmented Command R+ [29], GPT-4 Bing [25], Claude 3 RAG [27], Gemini Search [28] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [27], Gemini 1.5 [28], GPT-4o [26], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [25], Claude 3 Opus [27], Gemini 1.5 [28], Mixtral [32], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [27], Gemini [28], GPT-4 [25], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [32], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [26], Gemini 1.5 [28], Claude 3 [27], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [43], XLM-R [44], ByT5 [45] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [46], Galactica [47], Legal-BERT [48], StarCoder2 [49], Code LLaMA [50] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [51], Vicuna [52], ChatGLM [53], Baize [54], OpenChat [35] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [55], TinyLLaMA [56], DistilGPT2 [57], GPT-2 [58], LLaMA 2 7B [59] Designed for fast inference and deployment on edge devices
+Large language models (LLMs) have rapidly advanced due to progress in deep learning [14], the availability of massive datasets [15], and increasingly powerful hardware.
+Trained on large text corpora, they can generate fluent responses for tasks such as question answering, summarisation, translation, and content creation.
+Key innovations include new architectures [8], training strategies [16], and alignment methods [17].
+Several models have shaped the field.
+OpenAI’s GPT series [18] popularised autoregressive transformers through ChatGPT.
+Google’s T5 [16] reframed NLP as a text-to-text problem, while PaLM [19] scaled multitask learning.
+Anthropic’s Claude [12] introduced constitutional AI to emphasise safety, and Google’s Gemma [20] advanced reasoning-focused applications.
+Meta’s LLaMA models [3] provided a scalable open-source alternative.
+Community-driven projects also played a significant role.
+EleutherAI’s GPT-NeoX [21] offered an open substitute for proprietary GPT models.
+AI21’s Jurassic-2 [22] targeted enterprise integration, and Cohere’s Command R+ [23] combined generation with retrieval.
+Earlier, BERT [24] transformed bidirectional understanding.
+More recent lightweight models such as Phi-4, Mistral, and Alibaba’s Qwen 2.5 focus on efficiency, multilinguality, and adaptability.
+The latest generation continues this trend.
+DeepSeek-LLM [25], Mixtral [26], and StableLM prioritise open and efficient deployment, while DeepSeek-R1 highlights reasoning.
+Others push specialisation: Qwen2.5-Max uses Mixture-of-Experts for scalability, OpenAI’s o3-mini is designed for low-latency use, and DeepSeek-V3 has reached GPT-4-level performance in an open-source setting.
+As shown in Fig.
+2, today’s LLMs are not only powerful text generators but also increasingly efficient, transparent, and aligned with human values.
+They continue to shape the state of the art in both research and industry.
+Category Models Description General-Purpose GPT-4 [1], GPT-4o [27], Claude 3 [28], Claude 2.1 [29] Gemini 1.5 [35], Command R+ [23], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [26], Falcon 180B [32], DeepSeek-V3 [33], OpenChat [34] Customizable models for research applications Retrieval-Augmented Command R+ [23], GPT-4 Bing [1], Claude 3 RAG [28], Gemini Search [35] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [28], Gemini 1.5 [35], GPT-4o [27], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [1], Claude 3 Opus [28], Gemini 1.5 [35], Mixtral [26], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [28], Gemini [35], GPT-4 [1], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [26], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [27], Gemini 1.5 [35], Claude 3 [28], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [24], XLM-R [43], ByT5 [44] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [45], Galactica [46], Legal-BERT [47], StarCoder2 [48], Code LLaMA [49] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [50], Vicuna [51], ChatGLM [52], Baize [53], OpenChat [34], GPT-3.5 [54], LLaMA3.1 [55] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [56], TinyLLaMA [57], DistilGPT2 [58], GPT-2 [59], LLaMA 2 7B [3] Designed for fast inference and deployment on edge devices
 Figure 2: Extended LLM Categories, Models, and Descriptions
@@
 Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [60, 61].
-Explainability methods can be broadly categorized as intrinsic or post-hoc.
+Explainability methods can be broadly categorised as intrinsic or post-hoc.
 Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62].
@@
 Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives.
-Methods such as Q-LIME [67], S-LIME [68], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behavior, capturing non-linear relationships that the original method might miss.
-Furthermore, researchers have explored optimization strategies to enhance the efficiency and stability of LIME.
-Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
+Methods such as Q-LIME [68], S-LIME [69], Bay-LIME [70], and ALIME [71] use more advanced surrogate models to provide explanations that better reflect the underlying model’s behaviour, capturing non-linear relationships that the original method might miss.
+Furthermore, researchers have explored optimisation strategies to enhance the efficiency and stability of LIME.
+Techniques like OptiLIME [72] and G-LIME [73] work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
 Recent research has extended LIME to support specific data types and improve its interpretability.
 Adaptations such as PointNet LIME for 3D point cloud data [74], TS-MULE for time series data [75], Graph LIME for graph-structured data [76], Sound LIME for audio data [77], and B-LIME for ECG signal data [78] demonstrate how LIME can be tailored to diverse applications.
-In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimizing sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
+In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimising sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
 Other extensions include DSEG-LIME [79], which tailors LIME for segmentation models by generating perturbations at the object level rather than the pixel level, improving interpretability in image segmentation tasks.
@@
 US-LIME [82] employs uncertainty sampling to guide instance selection, increasing the reliability and faithfulness of local explanations.
-SS-LIME [78] uses self-supervised learning to identify important regions in the feature space and provide richer explanations.
-K-LIME [79] incorporates kernel-based methods into LIME’s regression step to model more complex local decision boundaries.
-Attention-LIME [80] leverages attention mechanisms to prioritize features during the explanation, aligning more closely with deep learning models.
-BMB-LIME [81] employs a boosted multi-branch approach to combine multiple local models, thereby enhancing consistency and robustness.
+S-LIME [83] uses self-supervised learning to identify important regions in the feature space and provide richer explanations.
 These methods aim to enhance interpretability across various domains, including segmentation, survival analysis, multimodal learning, and attention-based contexts.
-Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [82], LIME-SUP, LORE, ILIME, Aleph LIME, bLIMEy, NormLIME, DLIME, LIMEtree, MPS-LIME, Explain Explore, survLIME, survLIME-inf, CBR-LIME, GuidedLIME, audioLIME, MeLIME, C-LIME, gLIME, DIME, s-LIME, LIMESegment, UnRAvEL-LIME, Kernel-LIME, Specific-Input LIME, ExpLIMEable, GLIME, GMM-LIME, SEGAL, and Sig-LIME, reflecting the diverse and rapidly evolving landscape of post-hoc explainability [83].
+Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [84].
+Numerous other LIME-based and post-hoc explanation variants have been proposed, encompassing a wide range of methodological innovations.
+These have been comprehensively reviewed and categorised in [85].
 Figure 3: Taxonomy of LIME enhancement strategies: categorized into sampling techniques, surrogate model improvements, distance parameter adjustments, and optimization approaches.
@@
 [87] highlight the risks of opaque LLMs, including biases and misinformation, and call for increased transparency and accountability.
-Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [87, 88].
+Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [86].
 Amara et al.
 [89] recently introduced ConceptX, a concept-level attribution method highlighting semantically meaningful input phrases through a coalition-based Shapley framework, offering robust and human-aligned explanations.
-Mechanistic interpretability seeks to uncover how LLMs process information internally [90, 91, 92].
+Mechanistic interpretability seeks to uncover how LLMs process information internally [90, 28].
 Olah et al.
@@
 Recent advancements from the Transformer Circuits framework have further enriched mechanistic interpretability.
-Using causal scrubbing, Meng et al.
+Using causal scrubbing, Ameisen et al.
 [93] introduced Attribution Graphs to uncover emergent structures within neural networks.
@@
 [101] explore counterfactual explanations to improve human understanding.
-Recent works extend this perspective by enabling interactive explanations of black-box models through fine-tuned LLMs [105] and by benchmarking explanation quality in clinical domains [106].
+Recent works extend this perspective by enabling interactive explanations of black-box models [101] and by benchmarking explanation quality in clinical domains [102].
 These approaches improve the accessibility and transparency of LLM outputs in real-world decision-making.
+Recent research has explored cost-efficient, model-agnostic explanation methods for LLMs [103], which align well with the objectives of gSMILE; their approach leverages sampling and scoring strategies to generate faithful explanations without requiring access to model internals.
 Explainability enhances debugging, bias detection, regulatory compliance, and user trust in LLMs [60].
@@
 Recent work has demonstrated SMILE’s applicability beyond language, such as in instruction-based image editing [104], highlighting its modality-agnostic nature.
-New efforts address rare concept discovery via tailored sparse autoencoders [109] and propose hybrid architectures like concept bottleneck models [110] and symbolic program compositions [111] to enable more transparent reasoning.
-Future research should emphasise standardised benchmarks, hybrid mechanistic and human-centric methods, and integrating explainability into AI governance frameworks.
 3 Problem Definition
@@
 We begin by defining the key notation and problem setup, followed by a formal description of how we quantify the impact of input perturbations.
-We then describe the use of a surrogate model to approximate model behavior locally and provide theoretical justification based on the Wasserstein distance [63, 112, 113].
+We then describe the use of a surrogate model to approximate model behaviour locally and provide theoretical justification based on the Wasserstein distance [64, 105, 106].
 Lastly, we outline a structured methodology for implementing a Wasserstein-based surrogate framework tailored to large language models [62].
+Word Mover’s Distance (WMD) is the 1-Wasserstein (Earth Mover’s) distance between two documents represented as normalised bags-of-words in an embedding space, using the Euclidean distance between word embeddings as the ground cost [107].
+WMD provides a semantically meaningful measure of dissimilarity between documents by capturing the minimum amount of semantic “movement” needed to align their word distributions, making it well-suited for our interpretability framework.
 3.1 Notation and Problem Setting
 Setting the Stage.
-We aim to construct a local surrogate model that explains the behavior of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt.
+We aim to construct a local surrogate model that explains the behaviour of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt.
 The goal is to understand how specific parts of the input affect the model’s output distribution.
@@
 3.2 Input-Level Distance
-To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [115]:
-$\delta_{x_{j}}=\mathrm{WMD}(x,\hat{x}_{j}),$ (1)
+To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [107].
+We denote this input-level semantic distance as IWMD (Inner WMD):
+$\delta_{x_{j}}=\mathrm{IWMD}(x,\hat{x}_{j}),$ (1)
 Where $\delta_{x_{j}}\in\mathbb{R}_{\geq 0}$ denotes the semantic distance between the original and perturbed prompts.
 3.3 Output-Level Distribution Shift
-To assess the effect of input perturbation on the model’s behavior, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs:
-$\Delta(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
+To assess the effect of input perturbation on the model’s behaviour, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs.
+We refer to this as OWMD (Outcome WMD):
+$\mathrm{OWMD}(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
+$\Delta(x,\hat{x}_{j})=\mathrm{OWMD}(x,\hat{x}_{j}),$ (3)
 where $y\in\mathcal{Y}$, and $W(p,q)$ denotes the Wasserstein distance between two probability distributions $p$ and $q$.
@@
 Perturbations that are semantically closer to the original input are weighted more heavily. We use a Gaussian-based weighting scheme [64]:
-$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right),$ (3)
+$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma}\right)^{2}\right),$ (4)
 Where $w_{j}\in(0,1]$ denotes the relevance weight of $\hat{x}_{j}$, and $\sigma>0$ is a kernel width parameter that controls the rate of exponential decay.
@@
 $h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j},$ (5)
-where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimizing the weighted squared error.
+where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimising the weighted squared error.
 $\min_{\theta}\sum_{j=1}^{J}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2},$ (6)
-Where the loss ensures local fidelity between the surrogate and the observed output shifts.
+Where the loss ensures local ATT fidelity between the surrogate and the observed output shifts.
 3.6 Theoretical Justification (Lipschitz Smoothness)
 Shift Function Definition.
-We define a shift function that quantifies the model’s response to perturbation:
+We define a shift function that locally quantifies the model’s response to perturbation:
 $f(x_{a},\hat{x}_{a})=W\left(\pi^{(n)}(y\mid x_{a}),\pi^{(n)}(y\mid\hat{x}_{a})\right),$ (7)
-If $\pi^{(n)}$ is Lipschitz continuous concerning its input, then the function $f$ is also Lipschitz:
+If $\pi^{(n)}$ is Lipschitz continuous with respect to its input, then the function $f$ is also Lipschitz [108]:
 $\left|f(x_{a},\hat{x}_{a})-f(x_{b},\hat{x}_{b})\right|\leq L\times\left(\|x_{a}-x_{b}\|+\|\hat{x}_{a}-\hat{x}_{b}\|\right),\forall a,b\in\{1,\dots,N_{pert}\}$ (8)
 The notation $x_{a}$ and $\hat{x}_{a}$ represents a specific input and its corresponding perturbation.
-The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analyzing the Lipschitz continuity or stability across perturbations.
+The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analysing the Lipschitz continuity or stability across perturbations.
 $N_{pert}$ denotes the total number of perturbations.
-[…]
+where $L>0$ is the Lipschitz constant, and $\|\cdot\|$ denotes a norm over the input space (e.g., token embedding space).
 Relation to Total Variation.
@@
 $W(p,q)\leq\text{diam}(\mathcal{Y})\cdot\mathrm{TV}(p,q),$ (9)
-Where the total variation distance ($TV$) is given by:
+Where $\text{diam}(\mathcal{Y})=\sup_{y_{1},y_{2}\in\mathcal{Y}}d(y_{1},y_{2})$ denotes the diameter of the output space under the chosen ground metric $d$, i.e., the maximum possible distance between any two outputs.
+The total variation distance ($TV$) is given by:
 $\mathrm{TV}(p,q)=\frac{1}{2}\sum_{y\in\mathcal{Y}}|p(y)-q(y)|.$ (10)
-This bound highlights that small changes in distribution (in the TV sense) imply bounded Wasserstein distance.
+This bound highlights that small changes in distribution (in the TV sense) imply a bounded Wasserstein distance.
 3.7 Linear Surrogates Approximate Lipschitz Functions
@@
 where $\nabla f(x)^{\top}(x^{\prime}-x)$ denotes the dot product between the gradient of $f$ and the local displacement in input space.
-3.8 Putting It All Together
-The complete Wasserstein-based LIME surrogate procedure consists of:
-  1. 1.
-    Generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$.
-  2. 2.
-    Computing semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
-  3. 3.
-    Assigning similarity weights $w_{j}$ via Equation 3.
-  4. 4.
-    Extracting feature representations $z_{j}$ for each perturbation.
-  5. 5.
-    Fitting a local surrogate model $h_{\theta}$ by minimizing the loss (Equation 5).
-  6. 6.
-    Using $h_{\theta}$ to interpret the influence of each input feature on model behavior.
-This framework offers a principled and interpretable approach to analyzing large language models through localized perturbation and response analysis.
+The complete Wasserstein-based LIME surrogate procedure consists of generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$.
+The process then computes semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 3).
+Following this, similarity weights $w_{j}$ are assigned via Equation 4.
+The procedure continues by extracting feature representations $z_{j}$ for each perturbation and fitting a local surrogate model $h_{\theta}$ by minimising the loss (Equation 6).
+Finally, the fitted model $h_{\theta}$ is used to interpret the influence of each input feature on model behaviour.
+This framework offers a principled and interpretable approach to analysing large language models through localised perturbation and response analysis.
 4 Proposed Method
-The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analyzing how specific textual inputs influence the generated outputs.
+The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analysing how specific textual inputs influence the generated outputs.
 Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behavior [109, 62].
@@
 These perturbations reveal how different input elements contribute to the final prediction.
-Figure 4: SMILE flowchart for explaining image classification [11]
 SMILE records the model’s predictions for each perturbed input.
@@
 The results are used to train a weighted linear regression model that identifies the most influential input components.
-Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig.
+Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as illustrated in Fig.
 5.
-Instead of perturbing the generated image, as in SMILE, we perturb the input prompt by including or excluding certain words.
-We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between each generated output and the original output [115].
-This distance quantifies the difference in the production, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
-These similarity scores are then used as the outcome variable in a weighted linear regression model, with weights derived from the text distances and perturbations [116].
-This regression model helps determine how each word affects the generated text.
-The coefficients derived from this model quantify the influence of each word, which we use to create a visual heatmap highlighting the most influential words in the text prompt.
-This heatmap provides an intuitive visual representation of the elements of the prompt that contribute the most to the generated output.
+Instead of perturbing the generated image, as in the original SMILE framework, we perturb the input prompt.
+Each perturbed prompt $\hat{x}_{j}$ produces a corresponding text output, from which we compute the outcome-level semantic shift $\Delta(x,\hat{x}_{j})$ using the IWMD as defined  1.
+This allows us to identify which words in the input prompt most influence the model’s output.
+The set of output shifts $\{\Delta(x,\hat{x}_{j})\}_{j=1}^{J}$ is then used as the outcome variable in a weighted linear regression model, with weights $w_{j}$ computed from the input-level semantic distances $\delta_{x_{j}}$ via a Gaussian kernel (Eq.
+4).
+This regression estimates how each input feature (token) influences the generated output within the local neighbourhood of $x$.
+The resulting coefficients $\theta$ quantify the relative importance of each token, which we visualise as a heatmap highlighting the most influential words in the prompt.
+Figure 4: SMILE flowchart for explaining image classification [13]
 Using this enhanced three-step process, which builds on SMILE’s methodology [13], we aim to help users understand how particular elements of text prompts impact text generation.
-This technique enhances user control and predictability in text generation models by providing insight into which aspects of textual input, such as stylistic instructions or descriptive terms, have the most significant impact on the final output.
-The first step utilizes a text generation model that is trained on input text.
-The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output.
-The text generation model is expected to produce responses similar to the original response, varying only based on the critical components of the input texts [117].
-Different text permutations are utilized from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into separate words.
-The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [118].
-While the text generation model faces challenges in tasks such as counting objects and spatial reasoning, these issues can be mitigated by carefully selecting perturbation texts and restricting the input domain for perturbations [119].
-The text generation models enable the proposed method to create responses that differ from the original text solely in terms of keywords related to the perturbation texts.
-We use the Wasserstein distance to quantify these differences, comparing the distributions in both text embeddings [120, 112].
-This metric is employed because it provides a more comprehensive measure of the differences between text embeddings compared to other ECDF-based distances, such as Kolmogorov–Smirnov or Cramér–von Mises.
-Although these methods focus on specific aspects, the Wasserstein distance captures both broad and subtle changes in the data [120, 113].
-Finally, by applying linear regression, the space of input prompts is linked to the output space of the text generation model.
-As a result, the model’s behaviour can be interpreted by examining the regression coefficients [121, 116].
+This approach improves user control and predictability by showing which textual components, such as stylistic instructions or descriptive terms, most influence the final output.
+In the first step, we employ a black-box text generation model $\pi^{(n)}$ that receives an original prompt $x$ and produces an associated output $\pi^{(n)}(y\mid x)$.
+We then construct a set of perturbed prompts $\{\hat{x}_{j}\}_{j=1}^{J}$ by applying small, semantically meaningful edits to $x$ (e.g., removing or replacing specific tokens).
+For each $\hat{x}_{j}$, the model generates $\pi^{(n)}(y\mid\hat{x}_{j})$.
+Ideally, these outputs remain similar to the original output, with variation arising primarily from the altered components of the prompt [110].
+We quantify the distributional shift between the original and perturbed outputs using the OWMD as defined 2 $\Delta(x,\hat{x}_{j})$, as defined in Eq.
+3, computed in the output embedding space [111, 105].
+This metric is chosen for its ability to capture both broad and subtle differences compared to ECDF-based measures [106].
+Finally, we fit a weighted linear surrogate model $h_{\theta}$ (Eq.
+5) by minimising the loss in Eq.
+6, linking the feature representation $z_{j}$ of each perturbed prompt to its corresponding $\Delta(x,\hat{x}_{j})$.
+The learned coefficients $\theta$ provide an interpretable mapping between input prompt components and changes in the model’s generated output [112, 113].
+Figure 5: gSMILE flowchart for Large Language Models
 4.1 Text Generating and Perturbed Prompts
-To begin, we generate variations of the original text prompt, referred to as “perturbed prompts,” to analyze how subtle changes in the input affect the output of a text generation model [122, 123].
-The original prompt $p_{\text{org}}$ is first tokenized into individual words.
-From these tokens, multiple perturbed prompts $\{p_{\text{pert}}^{i}\}_{i=1}^{N}$ are constructed by selectively including or excluding specific words.
-Each perturbed prompt is then passed to the text generation model, which produces a corresponding output.
-The output generated from the original prompt, $t_{\text{org}}^{\prime}$, serves as the reference (or baseline) for comparison.
-For the original text prompt $p_{\text{org}}$, the text generation model $\phi_{\text{gen}}$ produces the output:
-$t_{\text{org}}^{\prime}=\phi_{\text{gen}}(p_{\text{org}})$ (11)
-Similarly, for each perturbed prompt $p_{\text{pert}}^{i}\in S_{\text{pert}}$, the model generates:
-$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in\{1,\dots,N\}$ (12)
-Here, $\phi_{\text{gen}}$ is the text generation function, and $t_{\text{pert}}^{\prime i}$ denotes the generated output for the $i$-th perturbation.
-These outputs are used to measure how individual word-level changes in the prompt influence the model’s behavior [63, 118].
+We begin by generating variations of the original text prompt $x$, referred to as “perturbed prompts” $\{\hat{x}_{j}\}_{j=1}^{J}$, to analyse how subtle changes in the input affect the output of a text generation model [114].
+The original prompt $x$ is first tokenised into individual tokens.
+From these tokens, multiple perturbed prompts are constructed by selectively including, excluding, or substituting specific words or phrases.
+Each perturbed prompt $\hat{x}_{j}$ is then passed to the text generation model $\pi^{(n)}$, which produces a corresponding output.
+The output generated from the original prompt, $\pi^{(n)}(y\mid x)$, serves as the reference (or baseline) for comparison.
+Formally, the outputs for the original and perturbed prompts follow the notation in Section 3:
+the model generates $y_{\text{org}}\sim\pi^{(n)}(y\mid x)$ for the original prompt, and $y_{\text{pert},j}\sim\pi^{(n)}(y\mid\hat{x}_{j})$ for each perturbation $j\in\{1,\dots,J\}$.
+Here, $\pi^{(n)}$ denotes the black-box text generation model, $y_{\text{org}}$ is the output for the original prompt, and $y_{\text{pert},j}$ denotes the output for the $j$-th perturbation.
+These outputs are subsequently used in the computation of the output shift $\Delta(x,\hat{x}_{j})$ (Eq.
+3) to measure how individual token-level changes in the prompt influence the model’s behavior [64, 115].
 4.2 Creating the Interpretable Space
-To represent the output space of the text generation model in a one-dimensional similarity space, we compute the Wasserstein distance between each generated text from a perturbed prompt and the output generated from the original prompt [112, 113].
-Throughout this process, the input structure remains unchanged, and only the perturbed words differ across inputs.
-To capture the semantic similarity between generated outputs, we use Word Mover’s Distance (WMD) to embed the texts into a feature space [115].
-These embeddings enable the focus to remain on high-level semantic content, thereby mitigating lexical noise and enhancing the reliability of the computed distances.
-The resulting scalar distances are used as the dependent variable in the subsequent regression model.
-The Wasserstein distance between each perturbed output $t_{\text{pert}}^{\prime i}$ and the original output $t_{\text{org}}^{\prime}$ is computed as:
-$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{P}\right)^{\frac{1}{P}},\quad\forall i\in\{1,\dots,N\}$ (13)
-In this equation:
-  • $\delta_{\text{WMD}}(\cdot)$ denotes the semantic embedding of a generated text,
-  • $n$ is the dimensionality of the embedding space,
-  • $P$ specifies the norm order used in the Wasserstein metric,
-  • and $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance computed for the $i$-th perturbation.
-Using p-values is crucial for validating the significance of the observed Wasserstein distances (WD) between texts generated under different prompt perturbations.
-By calculating a p-value, we test whether the observed WD could have occurred by chance, determining if the perturbation-induced changes are meaningful [124].
-To compute this p-value, we use the Bootstrap algorithm [125, 126], which estimates the probability of observing a WD as large as the observed one.
-This involves generating distributions of WDs through repeated sampling and comparing each sampled WD with the observed value.
-If a small proportion of sampled WDs exceed the observed WD, the resulting p-value confirms that the differences are statistically significant, supporting the validity of the perturbation effects.
-The Wasserstein-$P$ metric enables adjustment of the sensitivity of the Wasserstein Distance (WD) to different types of text variations by tuning the norm order $P$.
-Smaller values of $P$ emphasize fine-grained, local differences between distributions, while larger values capture broader, more global shifts [113].
-By experimenting with various values of $P$, we can select the one that best captures the semantic differences introduced by prompt perturbations, thereby improving both the reliability and interpretability of the analysis [113].
-To assess the statistical significance of the computed Wasserstein distances, we apply hypothesis testing using p-values.
-Distances associated with high p-values, indicating non-significant variation, are excluded from further consideration [124].
-As detailed in Algorithm 1, the Bootstrap algorithm [127, 108] is used to estimate these p-values for each observed WD.
-For example, let X and Y denote two sets of text-based embeddings.
-The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y.
-In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed.
-If boostWD exceeds the original observed WD, a counter (bigger) is incremented.
-After all iterations, the p-value (pVal) is estimated as the proportion of iterations where boostWD exceeds the observed WD, as outlined in Algorithm 1.
-Input: Two sets of text-based distances: $X$ and $Y$
-Output: Observed Wasserstein Distance (WD) and corresponding p-value (pVal)
+To represent the output space of the text generation model in a one-dimensional similarity space, we compute the OWMD distance between the output for each perturbed prompt $\hat{x}_{j}$ and the output for the original prompt $x$ (as defined in Eq.
+3[105, 106].
+Throughout this process, the input structure remains unchanged, and only the perturbed tokens differ across inputs.
+To capture semantic similarity between generated outputs, we employ the Outcome Word Mover’s Distance (OWMD) to measure distances between texts based on their representations in a semantic embedding space [107].
+Operating in this embedding space focuses the comparison on high-level semantic content, thereby mitigating lexical noise and improving the reliability of the computed distances.
+The resulting scalar values $\Delta(x,\hat{x}_{j})$ are then used as the dependent variable in the weighted regression model.
+Here, $\phi(\cdot)$ denotes the output embedding function, $n$ is the dimensionality of the embedding space, and $P$ specifies the norm order used in the Wasserstein metric.
+For each perturbation $j$, $\Delta(x,\hat{x}_{j})$ measures the distributional shift in the generated output relative to the baseline.
+After computing $\Delta(x,\hat{x}_{j})$ for each perturbation, we assess their statistical significance using p-values [116].
+This ensures that only meaningful output shifts are considered in the regression model.
+A p-value quantifies the probability that an observed Outcome Word Mover’s Distance (OWMD) between outputs could have occurred purely by random chance, under the null hypothesis that the perturbation has no real effect.
+To control the sensitivity of OWMD to different types of changes, we use the Wasserstein-$P$ metric [106].
+The norm order $P$ determines the scale of differences captured: smaller $P$ values emphasise fine-grained, local variations, whereas larger $P$ values capture broader, global shifts.
+By experimenting with different $P$ values, we select the one that best captures the semantic effects of prompt perturbations, thereby improving both the robustness and interpretability of the analysis.
+We compute p-values for each observed OWMD using a Bootstrap resampling procedure [117, 104], as detailed in Algorithm 1.
+This non-parametric approach repeatedly resamples the combined set of embeddings from both the original and perturbed outputs to construct an empirical distribution of OWMD values under the null hypothesis.
+The p-value is then estimated as the proportion of bootstrap distances greater than the observed OWMD.
+Perturbations with high p-values, indicating non-significant differences, are excluded from further interpretation.
+Input: Two sets of text embeddings: $X$ and $Y$
+Output: Observed Outcome Wasserstein Distance (OWMD) and p-value (pVal)
 1
 […]
-[…]
-[…]
-[…]
-6 $n\leftarrow\texttt{LX}+\texttt{LY}$;
+3 $\texttt{OWMD}\leftarrow\texttt{Wasserstein\_Dist}(X,Y)$;
+4 $\texttt{XY}\leftarrow X\cup Y$;
+5 $L_{X}\leftarrow|X|$, $L_{Y}\leftarrow|Y|$;
 6 $\texttt{bigger}\leftarrow 0$;
-8
-9for $i\leftarrow 1$ to MaxItr do
-10       Draw a random sample $e$ of size LX from $XY$;
-11       Draw a random sample $f$ of size LY from $XY$;
-[…]
-[…]
+7
+8for $i\leftarrow 1$ to MaxItr do
+9   Sample $e$ of size $L_{X}$ from XY without replacement;
+10   Sample $f$ of size $L_{Y}$ from XY without replacement;
+11 $\texttt{bootOWMD}\leftarrow\texttt{Wasserstein\_Dist}(e,f)$;
+12 if $\texttt{bootOWMD}>\texttt{OWMD}$ then
 13    $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
-15
+14
 15   end if
-17
+16
 17 end for
 […]
-[…]
-Algorithm 1 Bootstrap Algorithm for Wasserstein Distance (WD) P-Value Calculation
-Figure 5: gSMILE flowchart for Large Language Models
-When generating perturbation texts from the original prompt, the key terms in the prompt are essential for maintaining semantic similarity between perturbed prompts and the source.
-To account for this, we define a similarity metric between each perturbed prompt and the original, which is then used as a sample weight in the interpretable model.
-This similarity is quantified using Word Mover’s Distance (WMD), which calculates the minimum cost to transport the embedded words of one document to match another [115].
-The WMD between the original prompt $p_{\text{org}}$ and a perturbed prompt $p_{\text{pert}}^{i}$ is defined as:
-$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\cdot d(w_{k},w_{l})$ (14)
-subject to $\sum_{l=1}^{m}T_{kl}=d_{k},\quad\forall k=1,\dots,n$
-$\sum_{k=1}^{n}T_{kl}=d^{\prime}_{l},\quad\forall l=1,\dots,m$
-$T_{kl}\geq 0$
-Here:
-  • $w_{k}$ and $w_{l}$ denote the $k$-th and $l$-th words from the original and perturbed prompts respectively,
-  • $d(w_{k},w_{l})$ is the distance (typically cosine or Euclidean) between the embeddings of $w_{k}$ and $w_{l}$,
-  • $T_{kl}$ is the flow variable indicating how much of $w_{k}$’s weight is transported to $w_{l}$,
-  • $d_{k}$ and $d^{\prime}_{l}$ represent the normalized bag-of-words weights for the original and perturbed texts.
-To use this distance as a weighting factor in the regression model, we normalize it using a Gaussian kernel:
-$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})=\exp\left(-\left(\frac{\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma}\right)^{2}\right),\quad\forall i\in\{1,\dots,N\}$ (15)
-Here, $\sigma$ is a kernel width parameter that controls sensitivity to dissimilar inputs.
-The resulting similarity score $\pi_{i}$ falls in the range $[0,1]$, assigning higher weights to perturbed prompts that are semantically closer to the original.
+19 return $\texttt{OWMD},\texttt{pVal}$;
+Algorithm 1 Bootstrap-based P-Value Calculation for Outcome Wasserstein Distance (OWMD)
+When generating perturbed prompts from the original prompt $x$, preserving key terms is essential for maintaining semantic similarity between each perturbed prompt $\hat{x}_{j}$ and the source.
+To capture this similarity, we compute the input-level semantic distance $\delta_{x_{j}}$ using the Inner Wasserstein Distance over tokens (IWMD) [107], as introduced in Eq.
+1.
+This value is later transformed into a sample weight $w_{j}$ for the interpretable surrogate model (Eq.
+4).
+To incorporate input similarity into the regression model, each IWMD value is converted into a similarity weight via a Gaussian kernel, following Eq. 4:
+$w_{j}(x,\hat{x}_{j})=\exp\left[-\left(\frac{\mathrm{IWMD}(x,\hat{x}_{j})}{\sigma}\right)^{2}\right],\quad\forall j\in\{1,\dots,J\}.$ (12)
+Here, $\sigma>0$ is the kernel width parameter that controls the rate of decay in weight as semantic dissimilarity increases.
+The resulting weights $w_{j}\in(0,1]$ assign greater importance to perturbed prompts that are more semantically similar to the original prompt $x$, and down-weight those that differ substantially.
 4.3 Developing the Interpretable Surrogate Model
-To interpret the influence of individual words on text generation, we map the output space of the generation model to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [128].
-This enables the training of a local surrogate model that relates changes in input structure to variations in output behavior.
-In this weighted linear regression framework, the perturbation vectors (i.e., binary or one-hot encodings of included or excluded tokens) serve as independent variables.
-The response variable is the Wasserstein distance between the output of a perturbed prompt and that of the original prompt.
-Each perturbation is assigned a weight based on its semantic similarity to the original prompt [114, 129].
-The loss function minimized during training is a weighted squared error:
-$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\cdot\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$ (16)
-Here:
-  • $\phi_{\text{gen}}$ is the text generation model,
-  • $\delta_{\text{WMD}}$ denotes the function used to embed generated texts,
-  • $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance between the generated outputs,
-  • $\pi_{i}$ is the similarity-based kernel weight for the $i$-th perturbation,
-  • and $\hat{f}_{t}(p_{\text{pert}}^{i})$ is the prediction from the linear surrogate model.
-The surrogate model $\hat{f}_{t}$ is trained on the one-hot vector representation of the perturbation prompts $p_{\text{pert}}^{i,01}$, using the computed Wasserstein distances as targets [114].
+To interpret the influence of individual tokens on the model’s output, we map the output space of the black-box text generation model $\pi^{(n)}$ to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [118].
+This mapping enables the training of a local surrogate model that relates structural changes in the input prompt to variations in output behaviour.
+In this weighted linear regression framework, the independent variables are binary (or one-hot) perturbation vectors indicating which tokens from the original prompt $x$ are retained or altered in $\hat{x}_{j}$.
+The dependent variable is the output-level distribution shift $\Delta(x,\hat{x}_{j})$, defined in Eq.
+3 as the Wasserstein distance between the output distributions for $x$ and $\hat{x}_{j}$.
+Each perturbation is assigned a similarity-based weight $w_{j}$ from Eq.
+4, computed via the Gaussian kernel over the input-level distance $\delta_{x_{j}}$.
+The surrogate model parameters are estimated by minimising the weighted squared error:
+$\mathcal{L}(\pi^{(n)},\delta_{x_{j}},x)=\frac{1}{J}\sum_{j=1}^{J}w_{j}(x,\hat{x}_{j})\cdot\left(\Delta(x,\hat{x}_{j})-h_{\theta}(z_{j})\right)^{2}$ (13)
+The fitted coefficients $\theta$ quantify the contribution of each token to the observed output shift, enabling interpretable attribution.
+We evaluate the proposed attribution method using several state-of-the-art large language models: OpenAI’s gpt-3.5-turbo-instruct [54], Anthropic’s Claude 2.1 [29], and Meta’s LLaMA 3.1 Instruct Turbo (70B) [55].
 4.4 Evaluation Metrics
 In our investigation, we adopt a suite of evaluation metrics inspired by foundational work provided by Google [119], emphasising the multifaceted nature of assessing explainable models.
-This work highlights the significance of metrics such as accuracy, stability, fidelity and consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models.
+This work highlights the significance of metrics such as ATT accuracy, ATT stability, ATT fidelity, ATT faithfulness, and ATT consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models, as you can see in Fig.
+6.
 The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [119].
-Figure 6:
-4.4.1 Accuracy
-Accuracy measures how well the model’s output aligns with the expected results (ground truth).
+4.4.1 Attribution Accuracy
+Att Accuracy measures how well the model’s output aligns with the expected results (ground truth).
 Specifically, we compare the model’s attribution scores (attention to text elements) against ground truth labels that identify the most relevant text elements for specific generated text segments [120].
@@
 An accurate model should correctly assign higher attention to the text elements that most influence the generated text output, improving the interpretability and reliability of the model [120].
-4.4.2 Stability
-Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [134].
+Figure 6: Given a prompt (e.g., “What is the meaning of life?”), We compare ground truth attributions with model-generated attributions for input tokens.
+Attribution quality is evaluated using four metrics: ATT Accuracy quantifies alignment of attributions with ground-truth (via AttAUC scores).
+ATT faithfulness measures whether attributions reflect the importance of tokens for the model’s response (via auROC).
+ATT consistency assesses whether results remain stable across different model runs or hyperparameters.
+ATT stability examines the sensitivity of attributions to small perturbations in the input prompt.
+4.4.2 Attribution Faithfulness
+ATT faithfulness refers to the extent to which an explanation accurately reflects a model’s reasoning process.
+A faithful attribution assigns high importance to input tokens that genuinely drive the model’s output, rather than highlighting spurious correlations or irrelevant features.
+This property is widely recognised as essential for trustworthy interpretability in natural language processing and other domains [123, 124].
+In our study, we evaluate ATT faithfulness through correlation analysis between attribution-based metrics (ATT ACC, ATT F1, ATT AUROC) and externally reported benchmark accuracies of the models being assessed, including OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1 [125, 126, 127].
+Formally, we compute the Pearson correlation coefficient:
+$r=\frac{\sum_{i=1}^{n}\bigl{(}x_{i}-\bar{x}\bigr{)}\bigl{(}y_{i}-\bar{y}\bigr{)}}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\,\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}},$ (14)
+where $x_{i}$ are published benchmark ATT Accuracy values, $y_{i}$ are the corresponding attribution scores, and $\bar{x},\bar{y}$ denote their means.
+A strong positive $r$ indicates that attribution metrics faithfully track external measures of model competence.
+Thus, ATT faithfulness in our framework is treated as an externally validated property: explanations are considered reliable to the extent that attribution scores correlate with established benchmarks, ensuring interpretability methods provide meaningful insights into large language models, particularly in high-stakes applications [128].
+4.4.3 Attribution Stability
+ATT stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [129].
 In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations.
-Stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61].
+ATT stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61].
 It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [130].
@@
 6, assigns similar importance to the words make and rainy in both prompts, even though the token  was added in one case.
-This consistency is critical for ensuring the model behaves predictably across minor input variations [136].
-Without such stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [63].
-We use the Jaccard index to quantify stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 17:
+This ATT consistency is critical for ensuring the model behaves predictably across minor input variations [131].
+Without such ATT stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [64].
+We use the Jaccard index to quantify ATT stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 15:
 $\text{Jaccard}(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (15)
-This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater stability and similarity in explanations [134].
+This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater ATT stability and similarity in explanations [129].
 A higher Jaccard index confirms the model’s ability to yield stable outputs, thus providing consistent and reliable interpretations [8].
-By ensuring stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [61].
-4.4.3 Consistency
-In evaluating the consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times.
-Consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [63].
+By ensuring ATT stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [61].
+4.4.4 Attribution consistency
+In evaluating the ATT consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times.
+ATT consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [64].
 As shown in Fig.
 6, when the input phrase “what is the meaning of life?” was fed into the model multiple times, the output remained stable across iterations.
-This consistency is crucial for applications where repeatable and reliable results are required.
+This ATT consistency is crucial for applications where repeatable and reliable results are required.
 In our case, the model consistently assigns higher weights to the words “meaning” and “life”, demonstrating that it understands the importance of these words in generating the desired output [8].
 Such behaviour underscores the model’s ability to focus on the semantic core of the input, a trait essential for maintaining coherence in practical use cases [130].
-In practical applications—such as automated content generation, interactive systems, or scenario simulations—users depend on models to produce results that do not fluctuate unexpectedly.
+In practical applications, such as automated content generation, interactive systems, or scenario simulations, users depend on models to produce results that do not fluctuate unexpectedly.
 Any inconsistencies could undermine user trust and compromise the system’s overall effectiveness.
 Our model enhances user confidence by demonstrating consistent outputs and establishes a foundation for further optimisation and integration into more complex systems.
-This ability to maintain consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [61].
-4.4.4 Fidelity
-Fidelity is a metric used to evaluate how closely an explainable model replicates the behavior of a black-box model.
-Specifically, the predictions of the black-box model denoted as $f(X_{i})$, are compared to those of the explainable model, $g(X_{i})$, for each input perturbation $i$.
+This ability to maintain ATT consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [61].
+4.4.5 Attribution Fidelity
+ATT fidelity is a metric used to evaluate how closely an explainable model replicates the behaviour of a black-box model.
+Specifically, the predictions of the black-box model, denoted as $f(X_{j})$, are compared to those of the explainable model, $g(X_{j})$, for each input perturbation $j$.
 A variety of accuracy-based and error-based metrics are employed to assess this alignment, thereby quantifying how well the explainable model approximates the responses of the black-box model.
@@
 It is defined as:
-$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f})^{2}}$ (18)
-where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{i})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
+$R^{2}=1-\frac{\sum_{j=1}^{N_{p}}(f(X_{j})-g(X_{j}))^{2}}{\sum_{j=1}^{N_{p}}(f(X_{j})-\bar{f})^{2}}$ (16)
+where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{j})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
 In cases where data samples have different levels of importance, a weighted variant is used. The weighted coefficient of determination $R_{w}^{2}$ is given by [131, 133]:
-$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}_{w})^{2}}$ (19)
+$R_{w}^{2}=1-\frac{\sum_{j=1}^{N_{p}}(f(X_{j})-g(X_{j}))^{2}}{\sum_{j=1}^{N_{p}}(f(X_{j})-\bar{f}_{w})^{2}}$ (17)
 Here, $\bar{f}_{w}$ is the weighted mean of the black-box predictions.
@@
 The classical unweighted forms Mean Squared Error (MSE) and Mean Absolute Error (MAE) are:
-$\text{MSE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (21)
-$\text{MAE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (22)
-For situations requiring sample-level importance, the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE) are defined as [138]:
-$\text{WMSE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}w_{i}}$ (23)
-$\text{WMAE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot|f(X_{i})-g(X_{i})|}{\sum_{i=1}^{N_{p}}w_{i}}$ (24)
-Here, $w_{i}$ denotes the sample weight associated with the $i$-th perturbation.
+$\text{WMSE}=\frac{\sum_{j=1}^{N_{p}}w_{j}\cdot(f(X_{j})-g(X_{j}))^{2}}{\sum_{j=1}^{N_{p}}w_{j}}$ (19)
+$\text{WMAE}=\frac{\sum_{j=1}^{N_{p}}w_{j}\cdot|f(X_{j})-g(X_{j})|}{\sum_{j=1}^{N_{p}}w_{j}}$ (20)
+Here, $w_{j}$ denotes the sample weight associated with the $j$-th perturbation.
 Moreover, the mean $L_{1}$ and mean $L_{2}$ norms are also applied to evaluate absolute and squared differences respectively [134, 131]:
-$L_{1}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (25)
-$L_{2}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (26)
-These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$—together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
-We compute fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model.
-This includes analyzing how different input perturbations affect fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behavior.
+$L_{1}=\frac{1}{N_{p}}\sum_{j=1}^{N_{p}}|f(X_{j})-g(X_{j})|$ (21)
+$L_{2}=\frac{1}{N_{p}}\sum_{j=1}^{N_{p}}(f(X_{j})-g(X_{j}))^{2}$ (22)
+These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$, together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
+We compute ATT fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model.
+This includes analysing how different input perturbations affect ATT fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behaviour.
+The overall workflow for computing ATT fidelity is shown in Figure 7.
+As depicted, input prompts are perturbed, and both the Global Model and the Local Surrogate Model produce similarity signals that are subsequently compared to evaluate ATT fidelity.
+Figure 7: ATT fidelity evaluation framework.
+The input prompt is perturbed, and both a Global Model (e.g., ChatGPT) and a Local Surrogate Model (SMILE) generate similarity signals.
+Their comparison quantifies how closely the surrogate approximates the black-box model’s behaviour.
 5 Experimental Results
-This section evaluates the proposed method’s ability to enhance explainability in instruction-based text generation.
-We analyse how the structure of textual prompts affects interpretability, how variations in input text influence the clarity of explanations, and how the proposed solution performs across diverse scenarios.
-Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI GPT [1], LLaMA [19], and Claude-AI [140] to ensure a comprehensive assessment.
+This section evaluates the proposed method’s ability to improve explainability in instruction-based text generation, focusing on both performance and its implications for model transparency and usability.
+We analyse how prompt structure influences interpretability, how phrasing variations affect the clarity of explanations, and how the method performs across diverse scenarios.
+For researchers, the analysis highlights linguistic factors that most affect interpretability; for practitioners, it offers actionable insights for prompt engineering and model debugging.
+Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, ensuring a robust and generalisable evaluation across architectures.
 5.1 Qualitative Results
-The qualitative evaluation demonstrates the proposed framework’s interpretability using scenario-based testing and visualisations.
-We designed diverse testing scenarios by varying textual prompts, including changes in sentence complexity, rephrasing, and the inclusion of domain-specific terminology.
-These scenarios examine how the model interprets and generates outputs based on nuanced variations in textual inputs.
+We assess interpretability through controlled variations of input prompts, including changes in sentence complexity, rephrasing, and the use of domain-specific terminology.
+This approach allows us to examine whether attribution patterns remain stable and meaningful under different linguistic conditions, a key requirement for trustworthy explainability tools.
+Observing these changes provides practical guidance for designing effective prompts and identifying potential reliability issues before deployment.
 5.1.1 Input Prompt Explainability
-The proposed framework applies scenario-based testing across multiple prompt variations to evaluate the interpretability of the model’s behaviour regarding different inputs.
-These variations include rephrased questions, descriptive modifiers, and changes in word choice.
-The system generates a text heatmap highlighting individual words’ contribution to the model’s output for each input, with colours indicating the importance scores.
-Figure 7 shows how varying the input prompt’s phrasing or focus leads to different word importance patterns, demonstrating the model’s sensitivity to linguistic changes.
+The framework applies multiple prompt variations, such as rewording questions, adding descriptive modifiers, or altering word choice.
+It generates heatmaps showing the contribution of each token to the model’s output, with colour intensity indicating importance scores.
+Figure 8 illustrates how small changes in prompt phrasing or focus can lead to noticeably different attribution patterns when using OpenAI’s gpt-3.5-turbo-instruct.
+These differences are not merely cosmetic: they reveal how the model’s attention can shift even for semantically similar inputs.
+Understanding such sensitivity helps practitioners design prompts that minimise unwanted variation, while also highlighting cases where model behaviour may be unstable, valuable both for prompt-engineering strategies and for reliability assessments in real-world applications.
 Figure 8: Text heatmaps showing word-level importance for various input prompts. The colour intensity represents the influence of each word on the model’s output.
 5.1.2 Reasoning Path and Token Attribution
-Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response.
-Figure 8 presents an example where the gSMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram.
-This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
-Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models.
-As shown in Figure 8, gSMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [141].
-Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the gSMILE framework.
-The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates gSMILE’s token-level importance heatmap for the same prompt.
+Beyond token-level heatmaps, the proposed framework generates structured reasoning paths that show how the model connects input elements to form its response.
+As illustrated in Figure 9, gSMILE not only assigns attribution scores to individual tokens but also produces a reasoning flow diagram, offering a unified view of both local word-level contributions and the broader logical structure behind the output.
+This dual-layer explanation bridges the gap between fine-grained token attribution and higher-level reasoning, enabling users to trace how specific inputs influence each decision step.
+The approach aligns with advanced methods used by leading AI research organisations, such as Anthropic’s Attribution Graphs, while extending flexibility across diverse tasks and models.
+In the example shown, gSMILE’s interpretability matches or surpasses industry standards, providing a richer and more adaptable explainability toolkit for practical deployment [12].
+Figure 9: Comparison between attribution visualisations by Anthropic’s Attribution Graphs (top) and the gSMILE framework (bottom) using OpenAI’s gpt-3.5-turbo-instruct.
+The former shows the reasoning path, while the latter presents token-level importance for the same prompt.
 [12]
 5.1.3 Word-Level Attribution Analysis on MMLU Using gSMILE
-To evaluate the interpretability of ChatGPT using the gSMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [142].
-This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities.
-For this study, we accessed a pre-formatted version of the dataset available on Kaggle [143] to facilitate prompt construction and experimentation.
-Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision.
-Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter.
-These visualisations demonstrate gSMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
-Figure 9: Attribution heatmaps generated by gSMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
+We assessed the interpretability of OpenAI’s gpt-3.5-turbo-instruct with the gSMILE framework on the Massive Multitask Language Understanding (MMLU) benchmark [135], which contains four-choice questions across a broad range of academic and professional domains.
+A pre-formatted version from Kaggle [136] was used to streamline prompt creation and experimentation.
+Figure 10 presents attribution heatmaps that highlight which tokens most strongly influenced the model’s choice.
+Darker red tones correspond to higher importance, while lighter shades indicate weaker influence.
+By surfacing these token-level attributions, gSMILE exposes the model’s internal focus when answering complex, domain-specific questions.
+Such analysis is valuable for diagnosing whether the model bases its answers on semantically relevant cues or potentially spurious correlations.
+For example, consistent highlighting of key domain terms across questions suggests stable reasoning.
+In contrast, high attribution to irrelevant tokens may indicate susceptibility to misleading prompt elements, critical information for practitioners deploying LLMs in knowledge-intensive tasks.
+Figure 10: Attribution heatmaps generated by gSMILE using OpenAI’s gpt-3.5-turbo-instruct on selected MMLU samples. Highlighted tokens are those with the greatest influence on the model’s prediction.
 5.1.4 Optimised Instructions from Google Research
-To further explore how instruction phrasing influences interpretability, we applied gSMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [144].
-These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
-We first applied gSMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem.
-” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
-We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, gSMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
+To examine how instruction phrasing affects interpretability, we applied gSMILE to two high-performing prompts from Google Research’s OPRO framework [137], which were optimised for zero-shot mathematical reasoning in LLMs.
+For GPT-3.5, we analysed the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem.” using 1024 perturbations.
+As shown in Figure 11, the tokens “arithmetic”, “approach”, and “solution” received the highest importance scores, indicating their role in directing the model toward structured problem-solving.
+For GPT-4, we analysed the prompt: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” Figure 12 shows that “numerical”, “command”, and “answer” were most influential, suggesting that technical and goal-oriented terms firmly steer reasoning in high-performing instructions.
+This Comparison highlights how effective prompt design amplifies semantically critical terms, improving reasoning reliability, a valuable insight for crafting robust instructions across different LLM architectures.
 Figure 11: Word-level importance in a high-performing instruction from Google Research, identified using gSMILE with GPT-3.5.
-Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using gSMILE with GPT-4.
+Figure 12: Word-level importance in an instruction optimised by Google Research, analysed using gSMILE with GPT-4.
 5.2 Quantitative Results
-In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to assess our explainability method across different text generation models and different scenarios.
-5.2.1 Accuracy
-To evaluate the performance of instruction-based text generation models, we tested ten different prompts and their variations through 64 perturbations.
-Heatmaps were generated to display the weights of each keyword in the prompts across various models, including OpenAI GPT, LLAMA, and Claude-AI.
-A ground truth was defined, assigning weights of 1 to critical keywords in the prompts and 0 to less relevant or auxiliary words.
-These ground truth values were compared with the extracted heatmaps to assess the interpretability and accuracy of each model.
-To quantify accuracy, we employed multiple metrics, including Attention Accuracy (ATT ACC), F1-Score for Attention (ATT F1), and Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC).
-The results were averaged across the ten different prompts, and each model’s performance is summarised in Table 1, showcasing their effectiveness based on these metrics.
-Among the evaluated models, Claude-AI demonstrated the strongest performance, achieving the highest scores in all three metrics: 0.82 in ATT ACC, 0.67 in ATT F1, and 0.88 in ATT AUROC.
-These results indicate Claude-AI’s superior ability to consistently focus on essential input tokens across varying prompts.
-In contrast, LLAMA achieved the second-highest accuracy but showed a notably lower F1-score (0.40), suggesting potential inconsistency in identifying key tokens.
-OpenAI GPT yielded balanced results but lagged slightly behind Claude-AI in all metrics, highlighting room for improvement in its attention fidelity.
-Ll, Claude-AI provides more reliable and semantically focused attention compared to its counterparts.
-Table 1: Performance metrics (ATT ACC, ATT F1, and ATT AUROC) for various
-models, evaluated under two conditions: different prompts and different images.
-Model name ATT ACC ATT F1 ATT AUROC
-OpenAI-GPT 0.70 0.59 0.84
-LLaMA 0.76 0.40 0.84
-Claude-AI 0.82 0.67 0.88
-5.2.2 Stability
-To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap.
-For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word.
-To evaluate stability, we added only the token  at the end of the text, extracted the coefficients again, and computed the Jaccard index between the coefficients before and after the perturbation.
-The averaged Jaccard indices across all ten prompts for each model are reported in Table 2.
-Higher scores indicate greater stability under minor input changes.
-Table 2: Stability scores across models based on average Jaccard index over ten prompts
+We quantitatively evaluate the proposed explainability method using multiple metrics, ATT Accuracy, ATT Stability, ATT Consistency, ATT Fidelity, and computational complexity, across different text generation models and scenarios.
+These metrics provide a multi-faceted assessment, capturing not only the accuracy of attributions but also their robustness, reproducibility, and efficiency in practical applications.
+5.2.1 Attribution Accuracy
+We assessed attribution accuracy by testing ten distinct prompts and their variations, each subjected to 64 perturbations.
+For each prompt, we generated heatmaps showing the weight assigned to every keyword across three models: OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1.
+A ground-truth attribution map was defined by assigning a weight of 1 to critical keywords and 0 to auxiliary or less relevant tokens.
+Model-generated heatmaps were then compared to this ground truth.
+Accuracy was quantified using three complementary metrics: Attention Accuracy (ATT ACC), Attention F1-Score (ATT F1), and the Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC).
+Scores were averaged over all prompts, and the results are reported in Table 1.
+Anthropic’s Claude 2.1 achieved the highest scores across all metrics, 0.82 ATT ACC, 0.67 ATT F1, and 0.88 ATT AUROC, indicating a strong and consistent focus on the most relevant tokens.
+Meta’s LLaMA 3.1 Instruct Turbo (70B) ranked second in accuracy but had a notably lower ATT F1 (0.40), suggesting occasional inconsistency in identifying all key tokens within a prompt.
+OpenAI’s gpt-3.5-turbo-instruct showed balanced but slightly lower scores compared to Claude 2.1.
+These findings highlight that while all models can identify necessary tokens to some extent, Claude 2.1 demonstrates a more semantically precise and reliable attribution pattern, an essential property for trust in explainability-driven applications.
+Table 1: Performance metrics (ATT ACC, ATT F1, ATT AUROC, and published accuracy) for various models, evaluated under two conditions: different prompts and different images.
+Model name ACC ATT ACC ATT F1 ATT AUROC
+OpenAI’s gpt-3.5-turbo-instruct 0.70 0.70 0.59 0.84
+Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.69 0.76 0.40 0.84
+Anthropic’s Claude 2.1 0.79 0.82 0.67 0.88
+5.2.2 Attribution Faithfulness
+We assessed attribution faithfulness across the three evaluated large language models, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, by measuring how well attribution scores align with the models’ externally reported factual accuracies.
+Table 1 presents both the published accuracies and our attribution metrics (ATT ACC, ATT F1, ATT AUROC).
+The analysis reveals a clear positive relationship: models with higher factual accuracy tend to exhibit more faithful attributions.
+For instance, Claude 2.1, with a reported factual accuracy of 78.5%, achieves the highest ATT AUROC (0.88), reflecting strong alignment between its most essential tokens and its correct predictions.
+In Comparison, OpenAI’s gpt-3.5-turbo-instruct (70% accuracy) and LLaMA 3.1 (68.8% accuracy) both yield a slightly lower ATT AUROC of 0.84.
+We quantified this trend by computing the Pearson correlation between published accuracy and each attribution metric (Table 2).
+ATT AUROC showed the most substantial alignment ($r\approx 0.994$), followed by ATT ACC ($r\approx 0.804$) and ATT F1 ($r\approx 0.802$).
+These results indicate that attribution faithfulness can serve as a practical proxy for overall model reliability.
+When a model performs better on external benchmarks, it is also more likely to assign importance to semantically relevant input tokens.
+This makes ATT AUROC, in particular, a valuable diagnostic for selecting models in high-stakes interpretability scenarios.
+Table 2: Pearson correlation coefficients ($r$) between published accuracy and attribution metrics.
+Metric Pearson $r$
+ATT ACC 0.804
+ATT F1 0.802
+ATT AUROC 0.994
+5.2.3 Attribution Stability
+We measure attribution stability using the Jaccard index, which quantifies the overlap between sets of necessary tokens before and after a small, controlled input perturbation.
+For each of the three evaluated models, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, we selected ten prompts and recorded the token-level attribution scores.
+We then introduced a minimal change by appending the token  to the end of each prompt, recalculated the attributions, and computed the Jaccard index between the pre- and post-perturbation sets of essential tokens.
+Table 3 reports the average stability scores across all prompts.
+Higher values indicate greater resilience of attribution patterns to irrelevant input changes.
+OpenAI’s gpt-3.5-turbo-instruct achieved the highest stability score (0.62), suggesting that its attention to key tokens is less likely to shift under trivial modifications.
+In contrast, both LLaMA 3.1 (0.45) and Claude 2.1 (0.44) exhibited lower stability, indicating greater sensitivity to small, semantically insignificant changes.
+From an interpretability standpoint, higher ATT stability is desirable in real-world deployments: it means that explanations are less likely to fluctuate unpredictably when prompts are rephrased or slightly altered, improving trust and consistency in model behaviour.
+Table 3: ATT stability scores across models based on average Jaccard index over ten prompts.
 Model name Jaccard Index
-OpenAI-GPT 0.62
-LLaMA 0.45
-Claude-AI 0.44
-5.2.3 Consistency
-To compute consistency, we ran the same code with 64 perturbations for each prompt over 10 iterations, ensuring that the outputs remained consistent across these repetitions for text generation models such as OpenAI GPT, LLAMA, and Claude-AI, as shown in Table 3.
-We computed variance and standard deviation metrics for each word coefficient.
-This step is essential to confirm that the model’s predictions are not random or heavily dependent on initialisation factors but are instead deterministic and robust.
-Among all models, OpenAI-GPT exhibited the highest consistency, achieving a variance of 0.0000 and a standard deviation of 0.0046.
-This indicates that its predictions are the most stable across repeated runs, confirming its robustness and reliability in generating interpretable outputs.
-Table 3: Consistency metrics for different models on the prompt
-“What is the meaning of life?”
+OpenAI’s gpt-3.5-turbo-instruct 0.62
+Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.45
+Anthropic’s Claude 2.1 0.44
+5.2.4 Attribution Consistency
+Attribution consistency measures how reproducible token-level importance scores are when the same analysis is repeated under identical conditions.
+For each model, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, we ran 64 perturbations for a given prompt across 10 independent iterations.
+For each token, we computed the variance and standard deviation of its attribution score across runs.
+Low variance indicates that the model’s attribution patterns are stable and not influenced by randomness in the perturbation process or internal model initialisation.
+Table 4 summarises the results using the example prompt: “What is the meaning of life?” OpenAI’s gpt-3.5-turbo-instruct achieved near-perfect reproducibility, with a variance of 0.0000 and a standard deviation of only 0.0046.
+In contrast, LLaMA 3.1 and Claude 2.1 displayed higher variability, with LLaMA showing the least consistency overall.
+From an interpretability perspective, high ATT consistency is crucial: if attribution maps fluctuate unpredictably between identical runs, users cannot reliably trust the explanations.
+The results here suggest that, at least for this prompt, GPT-3.5 provides the most deterministic and reliable attributions.
+Table 4: ATT consistency metrics for different models on the prompt “What is the meaning of life?”
 Model name Variance Standard Deviation
-OpenAI-GPT 0.0000 0.0046
-LLaMA 0.0048 0.0663
-Claude-AI 0.0011 0.0331
-5.2.4 Fidelity for Different Instruct image editing diffusion based Models (IED)
-[…]
-Table 4: Fidelity metrics for different models on the prompt
-“What is the meaning of life?”
-[…]
-OpenAI-GPT 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
-LLaMA 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
-Claude-AI 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
-From the results, we observe that Claude-AI achieves the lowest WMSE, WMAE, and mean-L1, indicating superior performance in minimizing error.
-On the other hand, OpenAI-GPT performs best in terms of $R^{2}_{\omega}$, mean-L2, and adjusted $R^{2}_{\hat{\omega}}$, demonstrating stronger alignment and generalization of the surrogate model.
-LLaMA ranks in the middle across most metrics but does not lead in any.
-This suggests that different models excel in different fidelity dimensions, highlighting the importance of multi-metric evaluation when interpreting black-box model behavior.
-5.2.5 Fidelity Across Different number of Text Perturbations
-Tables 56, and 7 present the results of fidelity computations for varying numbers of perturbations across the OpenAI GPT, LLAMA, and Claude-AI models, respectively.
-These results are derived using a weighted linear regression as the surrogate model and the Wasserstein distance for computing distances.
-The evaluation metrics include weighted mean squared error (WMSE), weighted mean absolute error (WMAE), coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$), and average loss metrics such as mean-L1 and mean-L2.
-Table 5: Performance metrics for different numbers of perturbations in OpenAI GPT
+OpenAI’s gpt-3.5-turbo-instruct 0.0000 0.0046
+Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.0048 0.0663
+Anthropic’s Claude 2.1 0.0011 0.0331
+5.2.5 ATT Fidelity for Different Instruction-Based Models
+ATT fidelity measures how well a surrogate model can approximate the attribution patterns of a black-box model.
+For the prompt “What is the meaning of life?”, we evaluated OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1 using 32 perturbations.
+Fidelity was computed with a weighted linear regression surrogate and Wasserstein distance, assessing WMSE, weighted $R^{2}$ ($R^{2}_{\omega}$), WMAE, mean $L_{1}$, mean $L_{2}$, and weighted adjusted $R^{2}$ ($R^{2}_{\hat{\omega}}$).
+Table 5: ATT fidelity metrics for different models on the prompt “What is the meaning of life?”
+Model name WMSE $R^{2}_{\omega}$ WMAE $L_{1} mean$-$L_{2}$ $R^{2}_{\hat{\omega}}$
+OpenAI’s gpt-3.5-turbo-instruct 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
+Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
+Anthropic’s Claude 2.1 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
+The results reveal a trade-off between error minimisation and alignment quality.
+Anthropic’s Claude 2.1 achieves the lowest WMSE, WMAE, and mean-$L_{1}$, suggesting it produces the most faithful low-error surrogate approximations.
+Conversely, OpenAI’s gpt-3.5-turbo-instruct leads in $R^{2}_{\omega}$, mean-$L_{2}$, and adjusted $R^{2}_{\hat{\omega}}$, indicating stronger overall correlation and generalisation of the surrogate model to the original attribution patterns.
+LLaMA 3.1 shows mid-range performance across most metrics, without topping any category.
+From a practical perspective, these findings underscore that fidelity is multi-dimensional: a model optimising for minimal error may not consistently achieve the highest correlation with its surrogate, and vice versa.
+Thus, a multi-metric approach is essential when selecting models for high-stakes interpretability tasks.
+5.2.6 Attribution Fidelity Across Different Numbers of Text Perturbations
+Tables 68 report ATT fidelity for OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1 using varying numbers of perturbations.
+Fidelity was measured with a weighted linear regression surrogate and Wasserstein distance, using WMSE, WMAE, $R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$, mean-$L_{1}$, and mean-$L_{2}$.
+For gpt-3.5-turbo-instruct and LLaMA 3.1, larger perturbation counts (especially 256) generally reduced error metrics (WMSE, WMAE, $L_{1}, mean$-$L_{2}$), suggesting a finer-grained local approximation of model attributions.
+In contrast, Claude 2.1 achieved its lowest WMSE and WMAE at only 64 perturbations, while still benefiting from higher counts in some loss metrics.
+Interestingly, all three models reached their highest $R^{2}_{\omega}$ at 32 perturbations, indicating that correlation-based alignment peaks early, while additional perturbations may improve local fit but dilute global attribution patterns.
+From a practical standpoint, these results suggest a trade-off: more perturbations can yield more precise local surrogate fits but may reduce overall attribution alignment.
+For interpretability workflows, an intermediate number of perturbations (e.g., 64–128) may offer a balanced choice, depending on whether correlation or error minimisation is the primary goal.
+Table 6: Performance metrics for different numbers of perturbations in OpenAI’s gpt-3.5-turbo-instruct
 […]
@@
 256 0.0116 0.4276 0.0675 0.0872 0.0163 0.4138
-Table 6: Performance metrics for different numbers of perturbations in LLaMA
+Table 7: Performance metrics for different numbers of perturbations in Meta’s LLaMA 3.1 Instruct Turbo (70B)
 […]
@@
 256 0.0058 0.2387 0.0605 0.0667 0.0118 0.2203
-Table 7: Performance metrics for different numbers of perturbations in Claude-AI
+Table 8: Performance metrics for different numbers of perturbations in Anthropic’s Claude 2.1
 […]
@@
 256 0.0191 0.2670 0.0902 0.0873 0.0163 0.2494
-For OpenAI GPT and LLaMA, increasing the number of perturbations generally improves fidelity across most metrics, with 256 perturbations achieving the best WMSE, WMAE, and loss values.
-Interestingly, for Claude-AI, 64 perturbations yield the lowest WMSE and WMAE, while 256 achieves the best loss (mean-L2 and mean-L1).
-However, the highest $R^{2}$ scores for all three models occur at 32 perturbations, suggesting that beyond a certain point, more perturbations may dilute interpretability despite improving local fit.
-Therefore, there is a trade-off between the number of perturbations and the reliability of the surrogate model, depending on the metric of interest.
-5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
-Fidelity is also computed by comparing various distance metrics for text generation models such as OpenAI GPT using text-to-text comparisons.
-The surrogate models used are Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME), as shown in Table 8.
-We explore different combinations of distance metrics, including Cosine similarity and Wasserstein Distance (WD).
-Fidelity is measured using WMSE, WMAE, mean-$L_{1}$, mean-$L_{2}$, and coefficient of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$).
-Table 8: Fidelity results for different distance measures with 30 perturbations
-WLR Fidelity Metrics
+5.2.7 Attribution Fidelity for Different Distance Metrics and Surrogate Models
+We compared ATT fidelity across combinations of distance metrics and surrogate models for OpenAI’s gpt-3.5-turbo-instruct, using 30 perturbations.
+Surrogates included Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME) [70], and distance metrics included Cosine similarity, Wasserstein Distance (WMD), and their combination (WMD+C).
+Fidelity was evaluated using WMSE, WMAE, $L_{1}, mean$-$L_{2}$, and coefficients of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$), as shown in Table 9.
+Table 9: ATT fidelity results for different distance measures with 30 perturbations.
+WLR ATT Fidelity Metrics
 […]
 Cosine Cosine 0.0172 0.3151 0.0659 0.1277 0.0412 0.1508
-Cosine WD 0.0216 0.4197 0.0899 0.1332 0.0385 0.2805
-WD WD 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
-WD Cosine 0.0048 0.4026 0.0329 0.0871 0.0296 0.2593
-WD+C WD+C 0.0349 0.5349 0.1589 0.3050 0.1468 0.4233
+Cosine WMD 0.0216 0.4197 0.0899 0.1332 0.0385 0.2805
+WMD WMD 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
+WMD Cosine 0.0048 0.4026 0.0329 0.0871 0.0296 0.2593
+WMD+C WMD+C 0.0349 0.5349 0.1589 0.3050 0.1468 0.4233
 BayLIME
 Cosine Cosine 0.0200 0.2285 0.0499 0.1067 0.0464 0.0434
-Cosine WD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
-WD WD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
-WD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
-WD+C WD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
-Our Proposed Method: gSMILE.
-Based on the results in Table 8, the combination where Wasserstein Distance (WD) is used for both input similarity and output shift consistently yields the highest values of the weighted coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$).
-Specifically, underweighted linear regression, the WD–WD configuration achieves $R^{2}_{\omega}=\textbf{0.7104}$ and $R^{2}_{\hat{\omega}}=\textbf{0.6409}$, which significantly outperform other configurations.
-Motivated by these findings, we introduce our method, called gSMILE (generalized SMILE), which utilizes the Wasserstein distance for both input and output comparisons.
-This choice enhances the surrogate model’s fidelity and stability in capturing local shifts in model behavior.
-We use gSMILE throughout subsequent sections as the foundation for explainability evaluations on large language models (LLMs).
-These include assessments of fidelity, consistency, and stability, ensuring the proposed approach is both theoretically grounded and empirically adequate.
+Cosine WMD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
+WMD WMD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
+WMD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
+WMD+C WMD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
+Key Findings and the gSMILE Approach.
+Across both surrogate models, using Wasserstein Distance for both input similarity and output shift consistently produced the most substantial alignment between the surrogate and the target model.
+In WLR, the WMD–WMD configuration achieved the highest $R^{2}_{\omega}$ (0.7104) and $R^{2}_{\hat{\omega}}$ (0.6409), substantially outperforming cosine-based approaches.
+This result motivated our proposed method, gSMILE (generalised SMILE), which standardises on WMD–WMD for attribution fidelity.
+The Wasserstein Distance captures fine-grained distributional differences in token attributions, making it particularly effective for text-to-text comparisons.
+By adopting WMD–WMD, gSMILE improves not only ATT fidelity but also stability across perturbations, providing a consistent basis for all subsequent evaluations (fidelity, consistency, stability) in this work.
+This design choice balances theoretical robustness with empirical performance, making it well-suited for LLM interpretability workflows.
 5.2.8 Computation Complexity
-We evaluate the computational efficiency of our proposed method, gSMILE, using three instruction-based text generation models: OpenAI GPT, LLaMA, and Claude-AI.
-Although these models differ in scale and architecture, the objective here is not to compare the models themselves but rather to assess the performance of different explainability methods applied to each model under identical hardware and runtime conditions.
-All evaluations are conducted with 60 perturbations per input prompt.
-We compare three explainability approaches—LIME, gSMILE, and Bay-LIME—based on their execution time.
-As reported in Table 9, gSMILE consistently requires the highest runtime across all model frameworks.
-This is mainly due to its reliance on the Wasserstein distance to compute similarity between output distributions, which involves solving high-dimensional optimization problems with time complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [145].
-By contrast, LIME and Bay-LIME utilize cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
-Despite the higher computational cost, gSMILE delivers significantly better fidelity, stability, and robustness to adversarial perturbations—essential properties for trustworthy explanations.
-These advantages are supported by results in Tables 8 and 2 and align with findings from prior work on SMILE [11], which demonstrated superior robustness against adversarial input variations.
-These findings suggest that gSMILE’s computational overhead is justified by the quality of explanations it provides, making it highly suitable for applications where interpretability, consistency, and resistance to manipulation are paramount.
-Table 9: Execution times (in seconds) for each framework and explainability method with 60 perturbations
+We assessed the computational efficiency of gSMILE using three instruction-based text generation models: OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1.
+The goal here is not to compare these models, but to evaluate how different explainability methods perform under identical hardware and runtime conditions, with 60 perturbations per input prompt.
+We compared three approaches, LIME, gSMILE, and Bay-LIME, based on execution time.
+As shown in Table 10, gSMILE consistently requires the most runtime.
+This is primarily due to its use of the Wasserstein distance for comparing output distributions, which involves solving high-dimensional optimisation problems with complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [138].
+By contrast, LIME and Bay-LIME rely on cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
+Although gSMILE incurs a higher computational cost, it delivers substantially better ATT fidelity, ATT stability, and robustness to adversarial perturbations, critical for producing trustworthy explanations.
+These benefits are supported by the results in Tables 9 and 3 and are consistent with prior findings on SMILE [13], which showed improved robustness against adversarial input variations.
+In applications where interpretability, stability, and resistance to manipulation are priorities, gSMILE’s additional runtime is a justified trade-off for higher-quality, more reliable explanations.
+Table 10: Execution times (in seconds) for each framework and explainability method with 60 perturbations.
 Framework LIME gSMILE Bay-LIME
-OpenAI-GPT 156.82 170.70 161.43
-LLaMA 113.00 118.99 116.40
-Claude-AI 355.09 372.95 349.88
+OpenAI’s gpt-3.5-turbo-instruct 156.82 170.70 161.43
+Meta’s LLaMA 3.1 Instruct Turbo (70B) 113.00 118.99 116.40
+Anthropic’s Claude 2.1 355.09 372.95 349.88
 5.2.9 Token-Level Attention for Different Sentence Structures
-To investigate how sentence structure affects token-level attention in large language models (LLMs), we analyzed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts.
+To investigate how sentence structure affects token-level attention in large language models (LLMs), we analysed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts.
 These prompts ranged from questions (e.g., “What is the meaning of life?”), commands (e.g., “Please give me the meaning of life”), to declarative formulations (e.g., “You must explain the meaning of life”).
 For each sentence, we extracted the attention values directly from the final layer of the LLM.
-We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualized using a box plot (see Figure 12), where each point represents a unique sentence.
+We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualised using a box plot (see Figure 13), where each point represents a unique sentence.
 Full-sentence prompts are viewable through hover text, allowing qualitative comparison of sentence structure with attention distribution.
+Figure 13: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures.
+Each point represents the attention score assigned to the token in a specific sentence instance.
 The analysis reveals noticeable variability in how these tokens are weighted depending on syntactic form.
@@
 These findings underscore the sensitivity of LLMs to surface-level structure in shaping token-level interpretability, offering insight into both model behaviour and prompt engineering strategies.
-Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures.
-Each point represents the attention score assigned to the token in a specific sentence instance.
 6 Limitations
@@
 Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
-Whereas most existing interpretability methods highlight important tokens in the model’s output [63, 64], gSMILE shifts attention to the input space.
+Whereas most existing interpretability methods highlight essential tokens in the model’s output [64, 65], gSMILE shifts attention to the input space.
 By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
-Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma.
-The resulting heatmaps and attribution scores provide an intuitive understanding of model behavior, empowering users to control and refine prompt design better.
+Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, Meta’s LLaMA 3.1 Instruct Turbo (70B), Claude.ai, and Gemma.
+The resulting heatmaps and attribution scores provide an intuitive understanding of model behaviour, empowering users to control and refine prompt design better.
 gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.

### v5

@@
 We present gSMILE (generative SMILE), a model-agnostic, perturbation-based framework for token-level interpretability in LLMs.
-Extending the SMILE methodology, gSMILE uses controlled prompt perturbations, Wasserstein distance metrics, and weighted linear surrogates to identify input tokens with the most significant impact on the output.
+Extending the SMILE methodology, gSMILE utilises controlled prompt perturbations, Wasserstein distance metrics, and weighted linear surrogates to identify input tokens that have the most significant impact on the output.
 This process enables the generation of intuitive heatmaps that visually highlight influential tokens and reasoning paths.
-We evaluate gSMILE across leading LLMs (OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo, and Anthropic’s Claude 2.1) using attribution fidelity, attribution consistency, attribution stability, attribution faithfulness, and attribution accuracy as metrics.
+We evaluate gSMILE across leading LLMs (OpenAI’s gpt-3.5-turbo-instruct, LLaMA 3.1 Instruct Turbo, and Anthropic’s Claude 2.1) using attribution metrics such as fidelity, consistency, stability, faithfulness, and accuracy.
 Results show that gSMILE delivers reliable human-aligned attributions, with Claude 2.1 excelling in attention fidelity and GPT-3.5 achieving the highest output consistency.
-These findings demonstrate gSMILE’s ability to balance model performance and interpretability, enabling more transparent and trustworthy AI systems.
+These findings demonstrate gSMILE’s ability to strike a balance between model performance and interpretability, enabling more transparent and trustworthy AI systems.
 Figure 1: Overview of gSMILE interpretability process: An input prompt is processed by a transformer-based LLM, and gSMILE generates a token-level heatmap highlighting influential words contributing to the output.
 1 Introduction
-Large Language Models (LLMs) such as GPT-4 [1], Claude [2], and LLaMA [3] have rapidly become central to artificial intelligence, powering applications in healthcare, education, law, and scientific research [4, 5, 6, 7].
-Built on the Transformer architecture [8] and trained on massive corpora, they generate fluent, context-aware responses for tasks ranging from dialogue and summarisation to reasoning and information extraction [6].
+Large Language Models (LLMs) such as GPT‑4 [1], Claude [2], and LLaMA [3] have rapidly become central to artificial intelligence, powering applications in healthcare, education, law, and scientific research [4, 5, 6, 7].
+Built on the Transformer architecture [8] and trained on massive corpora, they generate fluent, context‑aware responses for tasks ranging from dialogue and summarisation to reasoning and information extraction [6].
 Today, LLMs are embedded in widely used technologies, assisting clinicians in healthcare [4], supporting personalised education [5], streamlining customer service [6], and aiding legal research [7].
@@
 Interpretability methods can provide the insights needed to guide such prompt design.
-Moreover, interpretability approaches must be model-agnostic.
+Moreover, interpretability approaches must be model‑agnostic.
 While Anthropic has recently showcased advanced interpretability research for Claude 3.5, using techniques such as circuit tracing and attribution graphs to visualise how the model plans, manipulates concepts, and sometimes even fabricates reasoning [12], such access to internal mechanisms is limited to others with privileged model access.
@@
 The resulting coefficients $\theta$ quantify the relative importance of each token, which we visualise as a heatmap highlighting the most influential words in the prompt.
-Figure 4: SMILE flowchart for explaining image classification [13]
 Using this enhanced three-step process, which builds on SMILE’s methodology [13], we aim to help users understand how particular elements of text prompts impact text generation.
@@
 Ideally, these outputs remain similar to the original output, with variation arising primarily from the altered components of the prompt [110].
+Figure 4: SMILE flowchart for explaining image classification [13]
 We quantify the distributional shift between the original and perturbed outputs using the OWMD as defined 2 $\Delta(x,\hat{x}_{j})$, as defined in Eq.
@@
 The learned coefficients $\theta$ provide an interpretable mapping between input prompt components and changes in the model’s generated output [112, 113].
-Figure 5: gSMILE flowchart for Large Language Models
 4.1 Text Generating and Perturbed Prompts
@@
 3) to measure how individual token-level changes in the prompt influence the model’s behavior [64, 115].
+Figure 5: gSMILE flowchart for Large Language Models
 4.2 Creating the Interpretable Space
@@
 The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [119].
+Figure 6: Given a prompt (e.g., “What is the meaning of life?”), We compare ground truth attributions with model-generated attributions for input tokens.
+Attribution quality is evaluated using four metrics: ATT Accuracy quantifies alignment of attributions with ground-truth (via AttAUC scores).
+ATT faithfulness measures whether attributions reflect the importance of tokens for the model’s response (via auROC).
+ATT consistency assesses whether results remain stable across different model runs or hyperparameters.
+ATT stability examines the sensitivity of attributions to small perturbations in the input prompt.
 4.4.1 Attribution Accuracy
@@
 An accurate model should correctly assign higher attention to the text elements that most influence the generated text output, improving the interpretability and reliability of the model [120].
-Figure 6: Given a prompt (e.g., “What is the meaning of life?”), We compare ground truth attributions with model-generated attributions for input tokens.
-Attribution quality is evaluated using four metrics: ATT Accuracy quantifies alignment of attributions with ground-truth (via AttAUC scores).
-ATT faithfulness measures whether attributions reflect the importance of tokens for the model’s response (via auROC).
-ATT consistency assesses whether results remain stable across different model runs or hyperparameters.
-ATT stability examines the sensitivity of attributions to small perturbations in the input prompt.
 4.4.2 Attribution Faithfulness
@@
 Figure 8: Text heatmaps showing word-level importance for various input prompts. The colour intensity represents the influence of each word on the model’s output.
+5.1.2 Gender Fairness via Token-Level Attributions
+We further examined whether gendered phrasing affects attribution patterns by applying gSMILE to three closely related prompts: “What is the meaning of life?”, “What is the meaning of life from men’s perspective?”, and “What is the meaning of life from women’s perspective?”.
+In all three models, the neutral formulation consistently highlighted the tokens meaning and life, underscoring their central role in shaping the response.
+Once a gender qualifier was introduced, however, the attribution heatmaps shifted, revealing model-specific sensitivities.
+As shown in Fig.
+9, GPT-3.5 tended to assign greater weight to the token women than men, suggesting an asymmetric influence of gender terms.
+Claude 3.5 distributed its attention more evenly, balancing the contribution of both gendered tokens while still raising their importance relative to the baseline.
+By contrast, LLaMA 3.1 exhibited higher variability, with attribution sometimes amplifying one gender reference more strongly than the other.
+Figure 9: Comparison of gSMILE token-level attributions for GPT-3.5, LLaMA 3.1, and Claude 3.5 under three prompts: neutral, “men perspective”, and “women perspective”.
+Colour intensity encodes the relative influence of each token on the output distribution.
+These findings illustrate how gSMILE exposes fairness-related sensitivities that remain hidden if only the generated text is inspected.
+A gender term that absorbs disproportionate attribution suggests an underlying association that could bias the model’s reasoning.
+From a practical standpoint, this analysis highlights that avoiding gendered modifiers in prompts helps maintain attention on the semantic core of the task, thereby reducing the risk of unintended attribution drift.
 5.1.3 Reasoning Path and Token Attribution
@@
 A pre-formatted version from Kaggle [136] was used to streamline prompt creation and experimentation.
+Figure 11: Attribution heatmaps generated by gSMILE using OpenAI’s gpt-3.5-turbo-instruct on selected MMLU samples. Highlighted tokens are those with the greatest influence on the model’s prediction.
 Figure 11 presents attribution heatmaps that highlight which tokens most strongly influenced the model’s choice.
@@
 In contrast, high attribution to irrelevant tokens may indicate susceptibility to misleading prompt elements, critical information for practitioners deploying LLMs in knowledge-intensive tasks.
-Figure 10: Attribution heatmaps generated by gSMILE using OpenAI’s gpt-3.5-turbo-instruct on selected MMLU samples. Highlighted tokens are those with the greatest influence on the model’s prediction.
 5.1.5 Optimised Instructions from Google Research
@@
 4) Some models, particularly research-stage or open-source LLMs, do not provide API access and require manual local deployment.
+Taken together, these limitations highlight the trade-off between generality and practicality in perturbation-based interpretability.
+They also point toward future work in designing adaptive, cost-efficient perturbation strategies and incorporating causal verification mechanisms to strengthen the robustness of explanations.
 7 Conclusion
-Interpretability remains a cornerstone of responsible AI, particularly in the context of large language models (LLMs) and instruction-based text generation.
-As AI systems become increasingly embedded in domains such as business, education, healthcare, and public services, it becomes essential for users to understand how these models generate outputs from their prompts.
-Without such insight, trust in AI systems diminishes, limiting adoption in real-world workflows.
-This underscores the need for transparent and interpretable solutions to demystify the behaviour of these black-box models.
-Interpretability is not only a research priority but is likely to become a legal and operational requirement in alignment with emerging AI regulations and ethical standards.
-Organisations that fail to address interpretability may encounter regulatory hurdles and reduced user acceptance.
-This work introduced a novel interpretability framework tailored for instruction-based text generation models.
-Leveraging gSMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output.
-Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
+This work introduced gSMILE, a generative, model-agnostic framework for explaining LLM behaviour through token-level attribution and visual reasoning aids.
 Whereas most existing interpretability methods highlight essential tokens in the model’s output [64, 65], gSMILE shifts attention to the input space.
-By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
-Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, Meta’s LLaMA 3.1 Instruct Turbo (70B), Claude.ai, and Gemma.
-The resulting heatmaps and attribution scores provide an intuitive understanding of model behaviour, empowering users to control and refine prompt design better.
-gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.
-This work lays the groundwork for more transparent and trustworthy natural language systems.
-Future directions include extending gSMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
-8 Future Works
+By integrating perturbation analysis with Wasserstein-based distance measures and weighted linear surrogates, gSMILE offers a principled approach to local interpretability without requiring access to internal model parameters.
+Our evaluation across three instruction-tuned LLMs revealed clear performance contrasts:
+  • Claude 2.1 achieved the highest attribution fidelity (AttAUROC = 0.88) and lowest error measures, indicating more semantically focused attention.
+  • GPT-3.5-turbo-instruct exhibited exceptional consistency (lowest variance and standard deviation across repeated runs), making it highly predictable in explanation output.
+  • LLaMA 3.1 performed competitively but showed greater variability in token attribution, suggesting sensitivity to prompt phrasing.
+Overall, the resulting heatmaps and attribution scores provide an intuitive understanding of model behaviour, empowering users to control and refine prompt design better.
+gSMILE was evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including attribution metrics such as fidelity, consistency, stability, faithfulness, and accuracy.
+We systematically assessed how well each method captured token-level influence in LLMs.
+The results show that gSMILE consistently achieved higher fidelity in approximating black-box behaviour, while also producing more stable and repeatable attributions under small prompt perturbations.
+Compared to LIME and Bay-LIME, gSMILE demonstrated improved robustness and alignment with human-interpretable patterns, underscoring its suitability for practical use in settings where reliable interpretability is critical.
+These findings justify gSMILE’s contribution in two ways.
+First, it demonstrates that explainability metrics can meaningfully differentiate between LLMs beyond standard benchmark accuracy, providing new insight into their operational reliability.
+Second, it shows that a purely black-box, post-hoc approach can produce explanations with stability and fidelity comparable to privileged-access interpretability tools.
+At the same time, the limitations identified (prompt sensitivity, hallucination effects, computational overhead, and deployment constraints) highlight that the path toward fully practical interpretability remains open.
+Addressing these will be central to future research.
+Looking ahead, we see three promising directions:
   1. 1.
-    Extension to Mathematical and Equation-Based Text Generation: Future work will explore the application of the proposed interpretability framework to models focused on symbolic reasoning and math-related tasks, such as equation solving, theorem generation, and formal derivation.
-Instruction-based mathematical models like Minerva and MathGLM require high precision in prompt design, and understanding how prompt components affect symbolic output will be essential in educational and scientific settings.
+    Adaptive perturbation strategies: Future work can move beyond random sampling by using targeted perturbations that focus only on influential tokens or phrases.
+This reduces computational overhead while preserving fidelity.
+Inspired by approaches such as IBM’s CELL framework, perturbations could involve changing specific words rather than only removing or keeping them.
+This makes it possible to highlight the importance of selected words in longer sentences efficiently.
   2. 2.
-    Application to Paraphrase Generation Models: The methodology can be extended to paraphrasing models, where semantic preservation and variation are central.
-Analysing how individual words influence paraphrased outputs’ style, tone, or fidelity can support better user control in writing assistance, content rephrasing, and translation tools.
+    Causal verification methods: To mitigate hallucination-driven or spurious attributions, causal analysis can be integrated (e.g., perturbing tokens while holding other context fixed) to confirm whether an attribution genuinely drives the model’s behaviour.
+This would strengthen robustness in high-stakes applications.
   3. 3.
-    Adaptation to Multi-Turn and Context-Aware Generation: Another direction is to adapt the framework for multi-turn and context-aware systems, including dialogue agents.
-This would involve tracking how prompt changes impact responses’ consistency, coherence, and relevance across conversational sequences.
-  4. 4.
-    Cross-Task Generalisation: While this work focuses on instruction-based text generation, the interpretability strategy can be generalised to broader NLP tasks, including summarisation, classification, and question answering.
-Evaluating how different model families respond to prompt perturbations can uncover model-specific behaviours and improve transparency.
-  5. 5.
-    Benchmarking Interpretability and Prompt Sensitivity: We plan to contribute to standardised evaluation protocols for prompt sensitivity, attribution robustness, and local fidelity.
-These benchmarks will support consistent assessment across models and help identify best practices for interpretable generation.
-  6. 6.
-    Optimisation of Computational Efficiency: To address the computational demands of gSMILE, future work will investigate dimensionality reduction techniques and approximation strategies.
-These approaches aim to reduce execution time while preserving the integrity and interpretability of the model’s outputs.
+    Extension to multimodal and multilingual models: gSMILE currently focuses on English text.
+Extending the framework to multilingual LLMs and multimodal systems (text–image or text–speech) would test whether stability and fidelity hold across diverse input spaces, improving generalisability for real-world deployment.
+By uniting rigorous evaluation with interpretable visualisations, gSMILE makes LLMs more transparent without privileged access to their internals, thereby providing both researchers and practitioners with a practical tool for building more trustworthy AI systems.
 Data Availability
@@
 For the gSMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy.
-For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.
+For the proposed algorithm introduced in this paper, see Sara068/LLM_SMILE.